### v1

# Vizing’s Theorem in Near-Linear Time

Abstract
Vizing’s theorem states that any $n$-vertex $m$-edge graph of maximum degree $\Delta$ can be edge colored using at most $\Delta+1$ different colors [Vizing, 1964]. Vizing’s original proof is algorithmic and shows that such an edge coloring can be found in $O(mn)$ time. This was subsequently improved to $\tilde{O}(m\sqrt{n})$ time, independently by [Arjomandi, 1982] and by [Gabow et al., 1985]. [^1] Throughout, we use $\tilde{O}(f):=O(f\operatorname*{polylog}{(n)})$ to suppress log-factors in the number of vertices of the graph.
Very recently, independently and concurrently, using randomization, this runtime bound was further improved to $\tilde{O}(n^{2})$ by [Assadi, 2024] and $\tilde{O}(mn^{1/3})$ by [Bhattacharya, Carmon, Costa, Solomon and Zhang, 2024] (and subsequently to $\tilde{O}(mn^{1/4})$ time by [Bhattacharya, Costa, Solomon and Zhang, 2024]).
We present an algorithm that computes a $(\Delta+1)$-edge coloring in $\tilde{O}(m)$ time—in fact, even $O(m\log{\Delta})$ time—with high probability, giving a near-optimal algorithm for this fundamental problem.
1 Introduction
Given a simple undirected graph $G=(V,E)$ on $n$ vertices and $m$ edges, as well as an integer $\kappa\in\mathbb{N}^{+}$, a $\kappa$-edge coloring $\chi:E\rightarrow\{1,2,\ldots,\kappa\}$ of $G$ assigns a color $\chi(e)$ to each edge $e\in E$ so that any two adjacent edges receive distinct colors. The minimum possible value of $\kappa$ for which a $\kappa$-edge coloring exists in $G$ is known as the edge chromatic number of $G$. If $G$ has maximum vertex degree $\Delta$, any proper edge coloring would require at least $\Delta$ different colors. A classical theorem by Vizing shows that $\Delta+1$ colors are always sufficient [Viz64]. Moreover, it was proven by [Hol81] that it is NP-complete to distinguish whether the edge chromatic number of a given graph is $\Delta$ or $\Delta+1$, and therefore $\Delta+1$ is the best bound we can hope for with polynomial time algorithms.
Vizing’s original proof easily extends to an $O(mn)$ time algorithm, which was improved to $\tilde{O}(m\sqrt{n})$ in the 1980s by Arjomandi [Arj82] and Gabow et al. [GNK+85] independently. More recently, Sinnamon [Sin19] simplified the algorithms of [Arj82, GNK+85] by using randomization, while shaving off extra logarithmic factors from their runtime complexities, achieving a clean $O(m\sqrt{n})$ runtime bound. Very recently, this longstanding $O(m\sqrt{n})$ time barrier was bypassed in two recent concurrent works by Assadi [Ass24] and Bhattacharya et al. [BCC+24] which improve the runtime bound to two incomparable bounds of $\tilde{O}(n^{2})$ and $\tilde{O}(mn^{1/3})$, respectively. In a follow-up work, Bhattacharya et al. [BCSZ25] improved the $\tilde{O}(mn^{1/3})$ time bound of [BCC+24] to $\tilde{O}(mn^{1/4})$.
In this work, we resolve the time complexity of randomized $(\Delta+1)$-edge coloring (up to small logarithmic factors) by presenting a near-linear time algorithm for this problem.
Theorem 1.1.
There is a randomized algorithm that, given any simple undirected graph $G=(V,E)$ on $n$ vertices and $m$ edges with maximum degree $\Delta$, finds a $(\Delta+1)$-edge coloring of $G$ in $\tilde{O}(m)$ time with high probability.
Remarks:
  1. 1.
    Our algorithm can be implemented to run in $O(m\log{\Delta})$ time with high probability. Nonetheless, our main focus is on polynomial runtime improvements in general, and achieving a near-linear time algorithm in particular. For this reason as well as for the sake of simplicity and transparency of our techniques, the main body of the paper presents and analyzes an algorithm that runs in $O(m\log^{3}n)$ time with high probability, which showcases our most novel ideas. A more careful implementation of the same exact algorithm, with proper adaptations, achieves $O(m\log\Delta)$ time; we defer the details of this part to Appendix A.
  2. 2.
    The $O(m\log\Delta)$ time bound that our final algorithm achieves matches the longstanding time bound for $\Delta$-edge coloring bipartite graphs [COS01] (which, to the best of our knowledge, is also the best known runtime for $(\Delta+1)$-edge coloring bipartite graphs). Our algorithm is also related to a recent line of work in [BD23, BD24, Dha24] that focuses on the $\Delta=n^{o(1)}$ case, which in particular gives a randomized $(\Delta+1)$-coloring that runs in $O(m\Delta^{4}\log\Delta)$ time with high probability.
1.1 Related Work
There has been a long line of work on fast algorithms for edge coloring that use more than $\Delta+1$ colors. It was first shown in [KS87] that an edge coloring can be computed in $\tilde{O}(m)$ time when we have $\Delta+\tilde{O}(\sqrt{\Delta})$ different colors. In addition, there are algorithms which run in linear or near-linear time for $(1+\epsilon)\Delta$-edge coloring [DHZ19, BCPS24c, EK24, BD24, Dha24] when $\epsilon\in(0,1)$ is a constant. Most recently, it was shown in [Ass24] that even a $(\Delta+O(\log n))$-edge coloring can be computed in near-linear time.
There are other studies on restricted graph classes. In bipartite graphs, a $\Delta$-edge coloring can be computed in $\tilde{O}(m)$ time [CH82, COS01]. In bounded degree graphs, one can compute a $(\Delta+1)$-edge coloring in $\tilde{O}(m\Delta)$ time [GNK+85], and it was generalized recently for bounded arboricity graphs [BCPS24b]; see also [BCPS24a, CRV24, Kow24] for further recent results on edge coloring in bounded arboricity graphs. Subfamilies of bounded arboricity graphs, including planar graphs, bounded tree-width graphs and bounded genus graphs, were studied in [CY89, CN90, CK08].
Beside the literature in the classical sequential model, considerable effort has been devoted to the study of edge coloring in various computational models in the past few years, including dynamic [BM17, BCHN18, DHZ19, Chr23, BCPS24c, Chr24], online [CPW19, BGW21, SW21, KLS+22, BSVW24, DGS24], distributed [PR01, EPS14, FGK17, GKMU18, BBKO22, CHL+20, Ber22, Chr23, Dav23], and streaming [BDH+19, SB24, CMZ24, GS24] models, among others.
1.2 Technical Overview
We now present an overview of prior approaches to $(\Delta+1)$-edge coloring and describe our techniques at a high level. For the following discussions, we will assume basic familiarity with Vizing’s proof and its underlying algorithm; see Section 2.2 for more details on Vizing’s algorithm.
Prior Approaches: A generic approach for $(\Delta+1)$-edge coloring, dating back to Vizing’s proof itself, is to extend a partial $(\Delta+1)$-edge coloring of the graph one edge at a time, possibly by recoloring some subset of edges, until the entire graph becomes colored. As expected, the main bottleneck in the runtime of this approach comes from extending the coloring to the last few uncolored edges. For instance, given a graph $G=(V,E)$ with maximum degree $\Delta$, we can apply Eulerian partition to divide $G$ into two edge-disjoint subgraphs $G=G_{1}\cup G_{2}$ with maximum degrees at most $\left\lceil\Delta/2\right\rceil$. We then find $(\left\lceil\Delta/2\right\rceil+1)$-edge colorings of the subgraphs $G_{1}$ and $G_{2}$ recursively. Directly combining the colorings of $G_{1}$ and $G_{2}$ gives a coloring of $G$ with $\Delta+3$ colors, so we have to uncolor a $2/(\Delta+3)$ fraction of the edges—amounting to $O(m/\Delta)$ edges—and try to extend the current partial $(\Delta+1)$-edge coloring to these edges. Thus, the “only” remaining part is to figure out a way to color these final $O(m/\Delta)$ edges, and let the above approach take care of the rest.
This task of extending the coloring to the last $\Theta(m/\Delta)$ edges is the common runtime bottleneck of all previous algorithms. Vizing’s original algorithm [Viz64] gives a procedure to extend any partial coloring to an arbitrary uncolored edge $(u,v)$ by rotating some colors around $u$ and flipping the colors of an alternating path starting at $u$. The runtime of this procedure would be proportional to the size of the rotation, usually called a Vizing fan, and the length of the alternating path, usually called a Vizing chain, which are bounded by $\Delta$ and $n$ respectively. As such, the total runtime for coloring the remaining $O(m/\Delta)$ edges using Vizing fans and Vizing chains will be $O(mn/\Delta)$ time.
As one can see, flipping long alternating paths is the major challenge in Vizing’s approach. To improve this part of the runtime, [GNK+85] designed an algorithm that groups all uncolored edges into $O(\Delta^{2})$ types depending on the two colors of the Vizing chain induced by this edge. Since all the Vizing chains of the same type are vertex-disjoint, they can be flipped simultaneously and their total length is only $O(n)$. This means that the runtime of coloring all edges of a single type can be bounded by $O(n)$ time as well. This leads to an $O(n\Delta^{2})$ time algorithm for handling all $O(\Delta^{2})$ types; a more careful analysis can bound this even with $O(m\Delta)$ time. Finally, balancing the two different bounds of $O(mn/\Delta)$ and $O(m\Delta)$ yields a runtime bound of $O(m\sqrt{n})$ for coloring $O(m/\Delta)$ edges, which leads to an $\tilde{O}(m\sqrt{n})$ time algorithm using the above framework.
There has been some very recent progress that broke through this classical barrier of $O(m\sqrt{n})$ time in [GNK+85, Arj82]. In [BCC+24], the authors proved a new useful condition for extending the coloring to uncolored edges: if the uncolored edges admit a small vertex cover, then there is also a faster algorithm for extending the coloring to them. They then show how to precondition the problem so that uncolored edges do indeed admit a small vertex cover, leading to an algorithm with $\tilde{O}(mn^{1/3})$ time. In [Ass24], the author avoided the need for Eulerian partition and recursion altogether by instead designing a new near-linear time algorithm for $(\Delta+O(\log n))$-edge coloring. This algorithm borrows insights from sublinear matching algorithms in regular bipartite graphs by [GKK10] and is thus completely different from other edge coloring algorithms mentioned above. By using this algorithm, finding a $(\Delta+1)$-edge coloring directly reduces to the color extension problem with $O((m\log n)/\Delta)$ uncolored edges (by removing the colors of a $\Theta(\log{n})/\Delta$ fraction of the edges in the $(\Delta+O(\log{n}))$-edge coloring to obtain a partial $(\Delta+1)$-edge coloring first). Applying Vizing’s procedure for these uncolored edges takes additional $\tilde{O}(mn/\Delta)=\tilde{O}(n^{2})$ time, leading to an $\tilde{O}(n^{2})$ time algorithm for $(\Delta+1)$-edge coloring. Finally, in [BCSZ25], the authors showed that a $(\Delta+1)$-coloring can be computed in $\tilde{O}(mn^{1/4})$ time, by using the algorithm of [Ass24] for initial coloring of the graph and then presenting an improved color extension subroutine with a runtime of $\tilde{O}(\Delta^{2}+\sqrt{\Delta n})$ for coloring each remaining edge; the best previous color extension time bounds were either the trivial $O(n)$ bound or the bound $\tilde{O}(\Delta^{6})$ by [Ber22].
Our Approach: A Near-Linear Time Color Extension Algorithm
Our algorithm takes an entirely different path compared to previous work on this problem; specifically, we no longer attempt to design a faster color extension for a single edge, and instead color them in large batches, which allows for a much better amortization of runtime in coloring multiple edges. This ultimately leads to our main contribution — a new randomized algorithm for solving the aforementioned color extension problem for the last $O(m/\Delta)$ edges in $\tilde{O}(m)$ time. With this algorithm at hand, we can simply follow the Eulerian partition approach mentioned earlier and obtain an algorithm for $(\Delta+1)$-edge coloring whose runtime $T(m)$ follows the recursion $T(m)\leq 2T(m/2)+\tilde{O}(m)$; this implies that $T(m)=\tilde{O}(m)$, hence, giving us a near-linear time algorithm for $(\Delta+1)$-edge coloring. This way, we will not even need to rely on the $(\Delta+O(\log{n}))$-edge coloring algorithm of [Ass24] to color the earlier parts of the graph (although one can use that algorithm instead of Eulerian partition approach to the same effect).
We now discuss the main ideas behind our color extension algorithm. In the following, it helps to think of the input graph as being near-regular (meaning that the degree of each vertex is $\Theta(\Delta)$), and thus the total number of edges will be $m=\Theta(n\Delta)$; this assumption is not needed for our algorithm and is only made here to simplify the exposition.
Color Type Reduction: Recall that the runtime of $O(n\Delta^{2})$ for the color extension algorithm of [GNK+85] is due to the fact that there are generally $O(\Delta^{2})$ types of alternating paths in the graph and that the total length of the paths of each color type is bounded by $O(n)$ edges. However, if it so happens that the existing partial coloring of the graph only involves $O(\Delta)$ color types instead, then the same algorithm will only take $O(n\Delta)=O(m)$ time (by the near-regularity assumption) and we will be done. As such, the underlying idea behind our algorithm is to modify the current partial edge coloring (without decreasing the number of uncolored edges) so that the total number of color types reduces from $O(\Delta^{2})$ to $O(\Delta)$ only.
To explore this direction, let us assume for now the input graph is bipartite, which greatly simplifies the structure of Vizing fans and Vizing chains, thus allowing us to convey the key ideas more clearly; later we highlight some of the key challenges that arise when dealing with general graphs. We shall note that it has been known since the 80s that one can $\Delta$-edge color bipartite graphs in $\tilde{O}(m)$ time [CH82, COS01]. However, the algorithms for bipartite graphs use techniques that are entirely different from Vizing fans and Vizing chains, and which do not involve solving the color extension problem at all. In particular, prior to this work, it was unclear whether one can efficiently solve the color extension problem in bipartite graphs. Therefore, the assumption of a bipartite input graph does not trivialize our goal of using Vizing fans and Vizing chains for efficiently solving the color extension problem. Additionally, we can assume that in the color extension problem, the last $O(m/\Delta)$ edges to be colored can be partitioned into $O(1)$ matchings (this guarantee follows immediately from the recursive framework we outlined earlier), and that we deal with each of these matchings separately. In other words, we can also assume that the uncolored edges are vertex-disjoint.
Let $\chi$ be a partial $(\Delta+1)$-edge coloring, and for any vertex $w\in V$, let $\mathsf{miss}_{\chi}(w)\subseteq[\Delta+1]$ be the set of colors missing from the edges incident to $w$ under $\chi$. Given any uncolored edge $(u,v)$, the color type of this edge $(u,v)$ would be $\{c_{u},c_{v}\}$ for some arbitrary choices of $c_{u}\in\mathsf{miss}_{\chi}(u)$ and $c_{v}\in\mathsf{miss}_{\chi}(v)$ (it is possible for an edge to be able to choose more than one color type, but we fix one arbitrary choice among them); in other words, if we flip the $\{c_{u},c_{v}\}$-alternating path starting at $u$, then we can assign $\chi(u,v)$ to be $c_{v}$. To reduce the total number of different color types to $O(\Delta)$, we would have to make some color types much more popular: at the beginning, a type spans an $\Omega(1/\Delta^{2})$ proportion of the uncolored edges but we would like to have a type spanning an $\Omega(1/\Delta)$ proportion. For this purpose, we fix an arbitrary color type $\{\alpha,\beta\}$, and want to modify $\chi$ to transform the type of an arbitrary uncolored edge $(u,v)$ from $\{c_{u},c_{v}\}$ to $\{\alpha,\beta\}$ – we call this popularizing the edge $(u,v)$. To do this, we can simply flip the $\{\alpha,c_{u}\}$-alternating path $P_{u}$ starting at $u$ and the $\{\beta,c_{v}\}$-alternating path $P_{v}$ starting at $v$.
There are two technical issues regarding this path-flipping approach. Firstly, the alternating paths $P_{u}$ and $P_{v}$ could be very long, and require a long time for being flipped. More importantly, flipping $P_{u}$ and $P_{v}$ could possibly damage other $\{\alpha,\beta\}$-type (uncolored) edges that we popularized before. More specifically, say that we have popularized a set $\Phi$ of uncolored edges. However, when popularizing the next uncolored edge $(u,v)$, it could be the case that the $\{\alpha,c_{u}\}$-alternating path $P_{u}$ is ending at a vertex $u^{\prime}$ for some edge $(u^{\prime},v^{\prime})\in\Phi$. If we flip the path $P_{u}$, then $(u^{\prime},v^{\prime})$ would no longer be of $\{\alpha,\beta\}$-type as $\alpha$ would not be missing at $u^{\prime}$ anymore.
Our key observation is that when $|\Phi|$ is relatively small, most choices for an alternating path $P_{u}$ cannot be ending at edges in $\Phi$. Consider the above bad example where $P_{u}$ is a $\{c_{u},\alpha\}$-alternating path ending at $u^{\prime}$ for some $(u^{\prime},v^{\prime})\in\Phi$. Let us instead look at this from the perspective of the $\{\alpha,\beta\}$-type edge $(u^{\prime},v^{\prime})\in\Phi$. For any $(u^{\prime},v^{\prime})\in\Phi$ and any color $\gamma$, there can be at most one uncolored edge $(u,v)\notin\Phi$ whose corresponding path $P_{u}$ is the same $\{\gamma,\alpha\}$-alternating path starting at $u^{\prime}$; this is because any vertex belongs to at most one alternating path of a fixed type (here, the type $\{\gamma,\alpha\}$) and each vertex belongs to at most one uncolored edge (recall that the uncolored edges form a matching). This is also true for $\{\gamma,\beta\}$-type edges for the same exact reason. As such, ranging over all possible choices of $\gamma\in[\Delta+1]$, there are at most $O(|\Phi|\Delta)$ uncolored edges $(u,v)$ whose alternating paths could damage the set $\Phi$. Therefore, as long as the size of $\Phi$ is a $o(1/\Delta)$ fraction (or more precisely, a $O(1/\Delta)$ fraction, for a sufficiently small constant hiding in the $O$-notation) of all uncolored edges, the following property holds: a constant fraction of uncolored edges $e\notin\Phi$ can be popularized using the above method without damaging $\Phi$.
This property resolves both technical issues raised earlier simultaneously. Let $\lambda$ denote the number of uncolored edges. For the second issue, as long as $|{\Phi}|=o(\lambda/\Delta)$, we can take a random uncolored edge $(u,v)\notin\Phi$ and flip $P_{u}$ and $P_{v}$ if this does not damage any already popularized edge in $\Phi$; by the observation above, this happens with constant probability. For the first issue, we can show that the expected length of alternating paths $P_{u}$ and $P_{v}$, for the random uncolored edge picked above, is $O(m/\lambda)$; indeed, this is because the number of edges colored $\alpha$ or $\beta$ is $O(n)$, hence the total length of all alternating paths with one color being fixed to $\alpha$ or $\beta$ is $O(m)$. All in all, the total time we spend to popularize a single type $\{\alpha,\beta\}$ to become a $\Theta(1/\Delta)$ fraction of all uncolored edges is $O(\lambda/\Delta\cdot m/\lambda)=O(m/\Delta)$. Since coloring edges of a single type takes $O(n)=O(m/\Delta)$ time by our earlier discussion, we can color in this way a $\Theta(1/\Delta)$ fraction of all uncolored edges in $O(m/\Delta)$ expected time. As a direct corollary, we can color all uncolored edges in $O(m\log{n})$ time, hence solving the color extension problem, in (near-regular) bipartite graphs, in near-linear time. The above argument will be detailed in the proof of Lemma 4.2 in Section 6.
Collecting U-Fans in General Graphs: We now discuss the generalization of our scheme above to non-bipartite graphs. Existence of odd cycles in non-bipartite graphs implies that we can no longer assign a color type $\{c_{u},c_{v}\}$ to an uncolored edge $(u,v)$ for $c_{u}\in\mathsf{miss}_{\chi}(u),c_{v}\in\mathsf{miss}_{\chi}(v)$, and hope that flipping the $\{c_{u},c_{v}\}$-alternating path from $u$ allows us to color the edge $(u,v)$ with $c_{v}$ (because the path may end at $v$, and thus after flipping it $c_{v}$ will no longer be missing from $v$). This is where Vizing fans and Vizing chains come into play: in non-bipartite graphs, a color type of an uncolored edge $(u,v)$ is $\{c_{u},\gamma_{u,v}\}$ where $c_{u}$ is an arbitrary color in $\in\mathsf{miss}_{\chi}(u)$ but $\gamma_{u,v}$ is determined by the Vizing fan around $u$ and the alternating path that we take for coloring $(u,v)$ (or the Vizing chain of $(u,v)$). Thus, while as before we can switch $c_{u}$ with some fixed color $\alpha$, it is unclear how to flip alternating paths to change $\gamma_{u,v}$ to $\beta$ also, in order to popularize the edge $(u,v)$ to be of some designated type $\{\alpha,\beta\}$ —without damaging another popularized edge as a result.
To address this challenge, we rely on the notion of a u-fan, introduced by [GNK+85], which is the non-bipartite graph analogue of an uncolored edge in bipartite graphs. A u-fan of type $\{\alpha,\beta\}$ is a pair of uncolored edges $(u,v),(u,w)$ such that $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(u)\cap\mathsf{miss}_{\chi}(v)$. Consider the $\{\alpha,\beta\}$-alternating path starting at $u$. As at least one of $v$ or $w$ is not the other endpoint of this alternating path (say $v$), flipping this path allows us to assign the color $\beta$ to $(u,v)$. Consequently, as u-fans are similar to edges in bipartite graphs, we can still essentially (but not exactly) apply our color type reduction approach if all the uncolored edges are paired as u-fans. Therefore, it suffices to modify $\chi$ to pair all uncolored edges together to form u-fans.
In order to pair different uncolored edges together and form $u$-fans, we should first be able to move uncolored edges around. Such operations already appeared in some previous work on dynamic edge coloring [DHZ19, Chr23, Chr24]. Basically, for any uncolored edge $(u,v)$, we can modify $\chi$ to shift this uncolored edge to any position on its Vizing chain. This naturally leads to a win-win situation: if the Vizing chain is short, then $(u,v)$ can be colored efficiently using Vizing’s procedure; otherwise if most Vizing chains are long, then there must be a pair of Vizing chains meeting together after a few steps, so we can shift two uncolored edges to form a u-fan efficiently.
Let us make this a bit more precise. Fix an arbitrary color $\alpha$ and consider the set $U_{\alpha}$ of all the uncolored edges $(u,v)$ such that $\alpha\in\mathsf{miss}_{\chi}(u)$ and the respective Vizing chain is of type $\{\alpha,\cdot\}$. Also, assume there are $m_{\alpha}$ edges colored $\alpha$ under $\chi$. If most $\{\alpha,\cdot\}$-Vizing chains have length larger than $\Omega(m_{\alpha}/|U_{\alpha}|)$, then on average, two Vizing chains should meet within the first $O(m_{\alpha}/|U_{\alpha}|)$ steps; in this case, we can repeatedly pick two intersecting Vizing chains and create a u-fan by shifting their initial uncolored edges to the intersection of these Vizing chains. Given the length of the chains, this takes $O(m_{\alpha}/|U_{\alpha}|)$ time. Otherwise, the average cost of applying Vizing’s color extension procedure is $O(m_{\alpha}/|U_{\alpha}|)$, and in this case we can directly color all those edges in $O(m_{\alpha})$ time. Summing over all different $\alpha\in[\Delta+1]$ gives a near-linear runtime. The above argument will be detailed in the proof of Lemma 4.1 in Section 5.
The above discussion leaves out various technical challenges. For instance, moving around uncolored edges as described above breaks the assumption that the uncolored edges form a matching. Handling this requires dedicating different colors from $\mathsf{miss}_{\chi}(u)$ for every uncolored edge incident on a vertex $u$. This is formalized via the notion of separability in Section 3.1. Additionally, we have ignored all algorithmic aspects of (efficiently) finding pairs of intersecting Vizing chains, as well as the corner cases of Vizing fan intersections and fan-chain intersections. We defer the discussions on these details to the main body of the paper.
2 Preliminaries
We begin by describing the basic notation used throughout the paper in Section 2.1 and then give the classic definitions of Vizing fans and Vizing chains in Section 2.2. In Section 3, we introduce the new definitions that provide the basic building blocks for our algorithm.
2.1 Notation
Let $G=(V,E)$ be graph on $n$ vertices with $m$ edges and maximum degree $\Delta$ and let $\chi:E\longrightarrow[\Delta+1]\cup\{\bot\}$ be a (partial) $(\Delta+1)$-edge coloring of $G$. We refer to edges $e\in E$ with $\chi(e)=\bot$ as uncolored. Given a vertex $u\in V$, we denote the set of colors that are not assigned to any edge incident on $u$ by $\mathsf{miss}_{\chi}(u)$. We sometimes refer to $\mathsf{miss}_{\chi}(u)$ as the palette of $u$. We say that the colors in $\mathsf{miss}_{\chi}(u)$ are missing (or available) at $u$.
Given a path $P=e_{1},\dots,e_{k}$ in $G$, we say that $P$ is an $\{\alpha,\beta\}$-alternating path if $\chi(e_{i})=\alpha$ whenever $i$ is odd and $\chi(e_{i})=\beta$ whenever $i$ is even (or vice versa). We say that the alternating path $P$ is maximal if one of the colors $\alpha$ or $\beta$ is missing at each of the endpoints of $P$. We refer to the process of changing the color of each edge $e_{i}\in P$ with color $\alpha$ (resp. $\beta$) to $\beta$ (resp. $\alpha$) as flipping the path $P$. We denote by $|P|$ the length (i.e., the number of edges) of the alternating path $P$. We define the length $i$ prefix of the path $P$ to be the path $P_{\leq i}:=e_{1},\dots,e_{i}$.
Given an uncolored edge $e$ under $\chi$, by extending $\chi$ to $e$ we mean modifying $\chi$ so that $\chi(e)\neq\bot$ without creating new uncolored edges. Our algorithm will always work by modifying a partial coloring $\chi$; unless explicitly specified otherwise, every new concept we define (such as u-fans and separable collection) will be defined with respect to this particular partial coloring $\chi$.
2.2 Vizing fans and chains
Vizing Fans: We now define the notion of Vizing fans, which has been used extensively in the edge coloring literature [Viz64, GNK+85, Sin19].
Definition 2.1 (Vizing fan).
A Vizing fan is a sequence $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ where $u,v_{1},\dots,v_{k}$ are distinct vertices and $c_{1},\dots,c_{k}$ are colors such that
  1. 1.
    $\alpha\in\mathsf{miss}_{\chi}(u)$ and $c_{i}\in\mathsf{miss}_{\chi}(v_{i})$ for all $i\in[k]$.
  2. 2.
    $v_{1},\dots,v_{k}$ are distinct neighbours of $u$.
  3. 3.
    $\chi(u,v_{1})=\bot$ and $\chi(u,v_{i})=c_{i}$ for all $i>1$.
  4. 4.
    Either $c_{k}\in\mathsf{miss}_{\chi}(u)$ or $c_{k}\in\{c_{1},\dots,c_{k-1}\}$.
We say that the Vizing fan $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{0},c_{0}),\dots,(v_{k},c_{k})$ is $\alpha$-primed, has center $u$ and leaves $v_{1},\dots,v_{k}$. We refer to $c_{i}$ as the color of $v_{i}$ within $\operatorname*{\mathbf{F}}$. A crucial property is that we can rotate colors around the Vizing fan $\operatorname*{\mathbf{F}}$ by setting $\chi(u,v_{1})\leftarrow c_{1},\dots,\chi(u,v_{i-1})\leftarrow c_{i-1},\chi(u,v_{i})\leftarrow\bot$ for any $i\in[k]$. We say that $\operatorname*{\mathbf{F}}$ is a trivial Vizing fan if $c_{k}\in\mathsf{miss}_{\chi}(u)$. Note that, if $\operatorname*{\mathbf{F}}$ is trivial, we can immediately extend the coloring $\chi$ to $(u,v_{1})$ by rotating all the colors around $\operatorname*{\mathbf{F}}$ and setting $\chi(u,v_{k})\leftarrow c_{k}$.
Algorithm 1 describes the standard procedure used to construct Vizing fans. As input, it takes a vertex $u$ and a color $\alpha\in\mathsf{miss}_{\chi}(u)$, and returns an $\alpha$-primed Vizing fan with center $u$.
1 For each $x\in V$, let $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)$
2 $k\leftarrow 1$ and $v_{1}\leftarrow v$
3 $c_{1}\leftarrow\mathsf{clr}(v_{1})$
4 while $c_{k}\notin\{c_{1},\dots,c_{k-1}\}$ and $c_{k}\notin\mathsf{miss}_{\chi}(u)$ do
5       Let $(u,v_{k+1})$ be the edge with color $\chi(u,v_{k+1})=c_{k}$
6       $c_{k+1}\leftarrow\mathsf{clr}(v_{k+1})$
7       $k\leftarrow k+1$
8
return $(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$
[…]
The Algorithm Vizing: We now describe the algorithm Vizing that, given a Vizing fan $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ as input, extends the coloring $\chi$ to the edge $(u,v_{1})$ by building a Vizing chain. Algorithm 2 gives a formal description of this procedure.
1 if $\operatorname*{\mathbf{F}}$ is trivial then
2       $\chi(u,v_{i})\leftarrow c_{i}$ for all $i\in[k]$
3       return
4
5Let $P$ denote the maximal $\{\alpha,c_{k}\}$-alternating path starting at $u$
6 Extend $\chi$ to $(u,v_{1})$ by flipping the path $P$ and rotating colors in $\operatorname*{\mathbf{F}}$ (details in Lemma 2.2)
Algorithm 2 Vizing$(\operatorname*{\mathbf{F}})$
Thus, running $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ extends the coloring $\chi$ to the uncolored edge $(u,v_{1})$ by rotating colors in the Vizing fan $\operatorname*{\mathbf{F}}$ and flipping the colors of the alternating path $P$. We sometimes refer to the process of running $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ as activating the Vizing fan $\operatorname*{\mathbf{F}}$.
Lemma 2.2.
Algorithm 2 extends the coloring $\chi$ to the edge $(u,v_{1})$ in time $O(\Delta+|P|)$.
Proof.
To see that the path $P$ is well-defined, note that $\alpha\in\mathsf{miss}_{\chi}(u)$ and $c_{k}\notin\mathsf{miss}_{\chi}(u)$, so there is an $\{\alpha,c_{k}\}$-alternating path starting at $u$.
Extending the coloring $\chi$: By the definition of a Vizing fan, $c_{k}$ is the first repetition of some color in $\{c_{1},\ldots,c_{k-1}\}$, so there is a unique index $j\in[k-1]$ such that $c_{j}=c_{k}$, and $P$ has $(u,v_{j+1})$ as its first edge with color $c_{j}$. We consider the cases where the path $P$ does or does not have $v_{j}$ as an endpoint. If $P$ does not end at $v_{j}$, then we can rotate the colors of the first $j+1$ edges in the fan by setting $\chi(u,v_{1})\leftarrow c_{1},\dots,\chi(u,v_{j})\leftarrow c_{j}$ and flip the colors of the alternating path $P$. If $P$ does end at $v_{j}$, then we flip the colors of the alternating path $P$, rotate the colors of the fan by setting $\chi(u,v_{1})\leftarrow\chi(u,v_{2}),\dots,\chi(u,v_{k-1})\leftarrow\chi(u,v_{k})$, and set $\chi(u,v_{k})\leftarrow c_{k}$. Note that, while rotating the fan in this last step, we have $\chi(u,v_{j+1})=\alpha\neq c_{k}$.
Using standard data structures (see Section 7), we can implement this algorithm to run in time proportional to the number of edges that change their colors, which is $O(\Delta+|P|)$. ∎
[…]
3 Basic Building Blocks
In this section, we introduce the notation of u-fans, u-edges and separable collections, which are the definitions that work as the basic building blocks for our algorithms.
3.1 U-Fans and U-Edges
We begin by defining the notion of u-fans that was used by [GNK+85]. [^2] The term ‘u-fan’ was originally introduced by [GNK+85] as an abbreviation for ‘uncolored fan’.
Definition 3.1 (u-fan, [GNK+85]).
A u-fan is a tuple $\operatorname*{\boldsymbol{f}}=(u,v,w,\alpha,\beta)$ where $u$, $v$ and $w$ are distinct vertices and $\alpha$ and $\beta$ are distinct colors such that:
  1. 1.
    $(u,v)$ and $(u,w)$ are uncolored edges.
  2. 2.
    $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)\cap\mathsf{miss}_{\chi}(w)$.
We say that $u$ is the center of $\operatorname*{\boldsymbol{f}}$ and that $v$ and $w$ are the leaves of $\operatorname*{\boldsymbol{f}}$. We also say that the u-fan $\operatorname*{\boldsymbol{f}}$ is $\{\alpha,\beta\}$-primed and define $c_{\operatorname*{\boldsymbol{f}}}(u):=\alpha$, $c_{\operatorname*{\boldsymbol{f}}}(v):=\beta$ and $c_{\operatorname*{\boldsymbol{f}}}(w):=\beta$ (i.e. given a vertex $x\in\operatorname*{\boldsymbol{f}}$, $c_{\operatorname*{\boldsymbol{f}}}(x)$ is the available color that $\operatorname*{\boldsymbol{f}}$ ‘assigns’ to $x$).
Activating U-Fans: Let $\operatorname*{\boldsymbol{f}}$ be an $\{\alpha,\beta\}$-primed u-fan with center $u$ and leaves $v$ and $w$. The key property of u-fans is that at most one of the $\{\alpha,\beta\}$-alternating paths starting at $v$ or $w$ ends at $u$. Say that the $\{\alpha,\beta\}$-alternating path starting at $v$ does not end at $u$. Then, after flipping this $\{\alpha,\beta\}$-alternating path, both $u,v$ are missing color $\alpha$. Thus, we can extend the coloring $\chi$ by assigning $\chi(u,v)\leftarrow\alpha$. We refer to this as activating the u-fan $\operatorname*{\boldsymbol{f}}$.
We also define the notion of a u-edge similarly to u-fans.
Definition 3.2 (u-edge).
A u-edge is a tuple $\operatorname*{\boldsymbol{e}}=(u,v,\alpha)$ where $(u,v)$ is an uncolored edge and $\alpha$ is a color such that $\alpha\in\mathsf{miss}_{\chi}(u)$.
We say that $u$ is the center of $\operatorname*{\boldsymbol{e}}$ and that $\alpha$ is the active color of $\operatorname*{\boldsymbol{e}}$. For notational convenience, we also say that the u-edge $\operatorname*{\boldsymbol{e}}$ is $\alpha$-primed and define $c_{\operatorname*{\boldsymbol{e}}}(u):=\alpha$ and $c_{\operatorname*{\boldsymbol{e}}}(v)=\bot$. [^3] Whenever we refer to an ‘uncolored edge $e$’, we are referring to an edge $e\in E$ such that $\chi(e)=\bot$, whereas a ‘u-edge $\operatorname*{\boldsymbol{e}}$’ always refers to the object from LABEL:{def:u-edge} and is denoted in bold.
Collections of U-Components: While working with both u-fans and u-edges simultaneously, we sometimes refer to some $\operatorname*{\boldsymbol{g}}$ that is either a u-fan or a u-edge as a u-component. Throughout this paper, we often consider collections of u-components $\operatorname*{\mathcal{U}}$. We only use the term ‘collection’ in this context, so we abbreviate ‘collection of u-components’ by just ‘collection’. We will be particularly interested in collections satisfying the following useful property, which we refer to as separability.
Definition 3.3 (Separable Collection).
Let $\chi$ be a partial $(\Delta+1)$-edge coloring of $G$ and $\mathcal{U}$ be a collection of u-components (i.e. u-fans and u-edges). We say that the collection $\mathcal{U}$ is separable if the following holds:
  1. 1.
    All of the u-components in $\mathcal{U}$ are edge-disjoint.
  2. 2.
    For each $x\in V$, the colors in the multi-set $C_{\mathcal{U}}(x):=\{c_{\operatorname*{\boldsymbol{g}}}(x)\mid\operatorname*{\boldsymbol{g}}\in\mathcal{U},\,x\in\operatorname*{\boldsymbol{g}}\}$ are distinct.
Damaged U-Components: Suppose we have a partial $(\Delta+1)$-edge coloring $\chi$ and a separable collection $\operatorname*{\mathcal{U}}$ w.r.t. $\chi$. Now, suppose that we modify the coloring $\chi$. We say that a u-component $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$ is damaged if it is no longer a u-component w.r.t. the new coloring $\chi$.
We note that this can only happen for one of the following two reasons: (1) one of the uncolored edges in $\operatorname*{\boldsymbol{g}}$ is now colored, or (2) there is a vertex $x\in\operatorname*{\boldsymbol{g}}$ such the color $c_{\operatorname*{\boldsymbol{g}}}(x)$ that $\operatorname*{\boldsymbol{g}}$ assigns to $x$ is no longer missing at $x$.
The following lemma shows that flipping the colors of an alternating path cannot damage many u-components in a separable collection $\operatorname*{\mathcal{U}}$.
Lemma 3.4.
Let $\chi$ be a partial $(\Delta+1)$-edge coloring of a graph $G$, $\operatorname*{\mathcal{U}}$ a separable collection and $P$ a maximal alternating path in $\chi$. Then flipping the colors of the alternating path $P$ damages at most $2$ u-components in $\operatorname*{\mathcal{U}}$ (corresponding to the two endpoints of the path).
Proof.
Let $x$ and $y$ be the endpoints of the path $P$. Since only the palettes of $x$ and $y$ change after flipping the colors of $P$, only u-components $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$ that contain $x$ and $y$ may be damaged. Since the palette of $x$ changes by $1$ and the collection $\operatorname*{\mathcal{U}}$ is separable, only one $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$ containing $x$ may be damaged by the color $c_{\operatorname*{\boldsymbol{g}}}(x)$ no longer being available at $x$. The same holds for the vertex $y$. Thus, at most 2 u-components in $\operatorname*{\mathcal{U}}$ are damaged by this operation. ∎
3.2 Data Structures
In Section 7, we describe the data structures that we use to implement our algorithm. On top of the standard data structures used to maintain the edge coloring $\chi$, which allows us to implement Algorithms 1 and 2 efficiently, we also use data structures that allow us to efficiently maintain a separable collection $\operatorname*{\mathcal{U}}$. More specifically, the data structures that we use to implement a separable collection $\operatorname*{\mathcal{U}}$ support the following queries.
  • $\textsc{Insert}_{\operatorname*{\mathcal{U}}}(\operatorname*{\boldsymbol{g}})$: The input to this query is a u-component $\operatorname*{\boldsymbol{g}}$. In response, the data structure adds $\operatorname*{\boldsymbol{g}}$ to $\operatorname*{\mathcal{U}}$ if $\operatorname*{\mathcal{U}}\cup\{\operatorname*{\boldsymbol{g}}\}$ is separable and outputs fail otherwise.
  • $\textsc{Delete}_{\operatorname*{\mathcal{U}}}(\operatorname*{\boldsymbol{g}})$: The input to this query is a u-component $\operatorname*{\boldsymbol{g}}$. In response, the data structure removes $\operatorname*{\boldsymbol{g}}$ from $\operatorname*{\mathcal{U}}$ if $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$ and outputs fail otherwise.
[…]
The following claim shows that it is always possible to answer a Missing-Color query.
Claim 3.5.
For each $x\in V$, the set $\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)$ is non-empty.
Proof.
Let $d$ denote the number of uncolored edges incident on $x$. Since the collection $\operatorname*{\mathcal{U}}$ is separable, we have that $|C_{\operatorname*{\mathcal{U}}}(x)|\leq d$. Since $|\mathsf{miss}_{\chi}(x)|\geq d+1$, it follows that $|C_{\operatorname*{\mathcal{U}}}(x)|<|\mathsf{miss}_{\chi}(x)|$ and so $\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)\neq\varnothing$. ∎
Furthermore, the data structure supports the following initialization operation.
  • $\textsc{Initialize}(G,\chi)$: Given a graph $G$ and an edge coloring $\chi$ of $G$, we can initialize the data structure with an empty separable collection $\operatorname*{\mathcal{U}}=\varnothing$.
In Section 7, we show how to implement the initialization operation in $O(m)$ time and each of these queries in $O(1)$ time with the appropriate data structures. These queries provide the ‘interface’ via which our algorithm will interact with the u-components.
A Remark on Randomization: In order to get good space complexity and running time simultaneously, we implement the standard data structures (used for Algorithms 1 and 2) and the data structure supporting the preceding queries using hashmaps (see Section 7). The following proposition describes the hashmaps used by our data structures. [^5] We could have even used the construction of [Kus22] to obtain exponentially small error probability in this case, but since we do not need this stronger guarantee we stick with the simpler work of [DadH90].
Proposition 3.6 ([DadH90]).
There exists a dynamic dictionary that, given a parameter $k$, can handle $k$ insertions, deletions, and map operations, uses $O(k)$ space and takes $O(1)$ worst-case time per operation with probability at least $1-O(1/k^{7})$.
This implementation gives us the guarantee that across the run of the entire algorithm, every query made to one of these data structures takes $O(1)$ time and each initialization operation takes $O(m)$ time, with high probability. Since the randomization used for the hashmaps is independent of the randomization used in the rest of the algorithm, we implicitly condition on the high probability event that, every operation performed using a hashmap runs in $O(1)$ time throughout the rest of the paper. [^6] Alternatively, one can replace these hashmaps with balanced search trees to make these data structures deterministic, incurring $O(\log n)$ overhead in the running time of each operation.
3.3 Vizing Fans in Separable Collections
Within our algorithm, we only construct Vizing fans and Vizing chains in a setting where there is some underlying separable collection $\operatorname*{\mathcal{U}}$. To ensure that activating and rotating colors around Vizing fans does not damage too many u-components, we choose the missing colors involved in Vizing fan constructions so that they ‘avoid’ the colors assigned to the u-components in $\operatorname*{\mathcal{U}}$.
Definition 3.7.
Let $\operatorname*{\mathcal{U}}$ be a separable collection and $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ be a Vizing fan. We say that the Vizing fan $\operatorname*{\mathbf{F}}$ if $\operatorname*{\mathcal{U}}$-avoiding if $c_{i}\in\mathsf{miss}_{\chi}(v_{i})\setminus C_{\operatorname*{\mathcal{U}}}(v_{i})$ for each leaf $v_{i}\in\operatorname*{\mathbf{F}}$.
We say that a Vizing fan $\operatorname*{\mathbf{F}}$ is a Vizing fan of the u-edge $\operatorname*{\boldsymbol{e}}=(u,v,\alpha)$ if $\operatorname*{\mathbf{F}}$ is $\alpha$-primed, has center $u$ and its first leaf is $v$. The following lemma shows that we can always find a $\operatorname*{\mathcal{U}}$-avoiding Vizing fan for a u-edge.
Lemma 3.8.
Given a u-edge $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{U}}$, there exists a $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}$ of $\operatorname*{\boldsymbol{e}}$. Furthermore, we can compute such a Vizing fan in $O(\Delta)$ time.
Proof.
By 3.5, we can always find a collection of colors $\{\mathsf{clr}(x)\}_{x\in V}$ such that $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)$ for each $x\in V$. If we construct a Vizing fan $\operatorname*{\mathbf{F}}$ by calling Vizing-Fan and using such a collection of colors within the algorithm, then it follows that the Vizing fan $\operatorname*{\mathbf{F}}$ is $\operatorname*{\mathcal{U}}$-avoiding. By combining the standard implementation of Algorithm 1 with the queries described in Section 3.2 we can do this in $O(\Delta)$ time. ∎
The following lemma describes some crucial properties of $\operatorname*{\mathcal{U}}$-avoiding Vizing fans.
Lemma 3.9.
Let $\chi$ be a $(\Delta+1)$-edge coloring of a graph $G$ and $\operatorname*{\mathcal{U}}$ be a separable collection. For any u-edge $\operatorname*{\boldsymbol{e}}=(u,v,\alpha)\in\operatorname*{\mathcal{U}}$ with a $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}$, we have the following:
  1. 1.
    Rotating colors around $\operatorname*{\mathbf{F}}$ does not damage any u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}\}$.
  2. 2.
    Calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ damages at most one u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}\}$. Furthermore, we can identify the damaged u-component in $O(1)$ time.
Proof.
Let $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$. Rotating colors around $\operatorname*{\mathbf{F}}$ will only remove the color $c_{i}$ from the palette of a leaf $v_{i}$ appearing in $\operatorname*{\mathbf{F}}$. Since $c_{i}\notin C_{\operatorname*{\mathcal{U}}}(v_{i})$, these changes to the palettes will not damage any u-component in $\operatorname*{\mathcal{U}}$. However, rotating colors around $\operatorname*{\mathbf{F}}$ will color the edge $(u,v)$, damaging the u-edge $\operatorname*{\boldsymbol{e}}$.
Now, suppose we call $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$. If the Vizing fan $\operatorname*{\mathbf{F}}$ is trivial, then the algorithm rotates all the colors in $\operatorname*{\mathbf{F}}$ and sets $\chi(u,v_{k})\leftarrow c_{k}$. Rotating the colors can only damage $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{U}}$, and removing $c_{k}$ from the palette of $u$ can damage at most one other u-component in $\operatorname*{\mathcal{U}}$ since $\operatorname*{\mathcal{U}}$ is separable. Using the queries from Section 3.2, we can check if such a u-component exist and return it in $O(1)$ time. If $\operatorname*{\mathbf{F}}$ is not trivial, then the algorithm flips the $\{\alpha,c_{k}\}$-alternating path $P=\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})$ and rotates colors around $\operatorname*{\mathbf{F}}$. By similar arguments, rotating colors around $\operatorname*{\mathbf{F}}$ can only damage $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{U}}$. Let $x$ denote the endpoint of $P$ that is not $u$. Flipping the path $P$ might remove either the color $\alpha$ or $c_{k}$ from the palette of $x$, which might damage at most one u-component in $\operatorname*{\mathcal{U}}\setminus\{e\}$ since $\operatorname*{\mathcal{U}}$ is separable. We can again identify this u-component in $O(1)$ time using the queries from Section 3.2. It also removes the color $\alpha$ from the palette of $u$, but this cannot damage any u-component other than $\operatorname*{\boldsymbol{e}}$ (again, since $\operatorname*{\mathcal{U}}$ is separable). ∎
4 The Main Algorithm
Our final algorithm consists of two main components. The first component is an algorithm called Construct-U-Fans that takes a partial $(\Delta+1)$-edge coloring $\chi$ with $\lambda$ uncolored edges and either extends $\chi$ to $\Omega(\lambda)$ of these edges or modifies the coloring to construct a separable collection of $\Omega(\lambda)$ u-fans. Lemma 4.1 summarizes the behaviour of this algorithm.
Lemma 4.1.
Given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ and a set of $\lambda$ uncolored edges $U$, the algorithm Construct-U-Fans does one of the following in $O((m+\Delta\lambda)\log n)$ time:
  1. 1.
    Extends the coloring to $\Omega(\lambda)$ uncolored edges.
  2. 2.
    Modifies $\chi$ to obtain a separable collection of $\Omega(\lambda)$ u-fans $\mathcal{U}$.
The second component is an algorithm called Color-U-Fans that takes a collection of $\lambda$ u-fans and extends the coloring to $\Omega(\lambda)$ of the edges in the u-fans. Lemma 4.2 summarizes the behaviour of this algorithm.
Lemma 4.2.
Given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ and a separable collection of $\lambda$ u-fans $\mathcal{U}$, the algorithm Color-U-Fans extends $\chi$ to $\Omega(\lambda)$ edges in $O(m\log n)$ time w.h.p.
In Sections 5 and 6, we describe these algorithms and prove Lemmas 4.1 and 4.2 respectively. Using these lemmas, we now show how to efficiently extend an edge coloring $\chi$ to the remaining uncolored edges.
Lemma 4.3.
Given a graph $G$ and a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges $U$, we can extend $\chi$ to the remaining uncolored edges in time $O((m+\Delta\lambda)\log^{2}n)$ w.h.p.
Proof.
Let $U$ denote the set of edges that are uncolored by $\chi$. We can then apply Lemma 4.1 to either extend $\chi$ to a constant fraction of the edges in $U$ or construct a separable collection of $\Omega(\lambda)$ u-fans $\mathcal{U}$ in $O((m+\Delta\lambda)\log n)$ time. In the second case, we can then apply Lemma 4.2 to color $\Omega(\lambda)$ of the edges contained in these u-fans in $O(m\log n)$ time w.h.p. Thus, we can extend $\chi$ to a constant fraction of the uncolored edges in $O((m+\Delta\lambda)\log n)$ time w.h.p. After repeating this process $O(\log\lambda)\leq O(\log n)$ many times, no edges remain uncolored. Thus, we can extend the coloring $\chi$ to the entire graph in $O((m+\Delta\lambda)\log^{2}n)$ time w.h.p. ∎
We now use this lemma to prove Theorem 1.1.
Proof of Theorem 1.1..
We prove this by applying Lemma 4.3 to the standard Euler partition framework [GNK+85, Sin19]. Given a graph $G$, we partition it into two edge-disjoint subgraphs $G_{1}$ and $G_{2}$ on the same vertex set such that $\Delta(G_{i})\leq\lceil\Delta/2\rceil$ for each $G_{i}$, where $\Delta(G_{i})$ denotes the maximum degree of $G_{i}$. We then recursively compute a $(\Delta(G_{i})+1)$-edge coloring $\chi_{i}$ for each $G_{i}$. Combining $\chi_{1}$ and $\chi_{2}$, we obtain a $(\Delta+3)$-edge coloring $\chi$ of $G$. We then uncolor the two smallest color classes in $\chi$, which contain $O(m/\Delta)$ edges, and apply Lemma 4.3 to recolor all of the uncolored edges in $\chi$ using only $\Delta+1$ colors in $O(m\log^{2}n)$ time w.h.p.
To show that the total running time of the algorithm is $O(m\log^{3}n)$, first note that the depth of the recursion tree is $O(\log\Delta)$. Next, consider the $i^{th}$ level of the recursion tree, for an arbitrary $i=O(\log\Delta)$: we have $2^{i}$ edge-disjoint subgraphs $G_{1},\dots,G_{2^{i}}$ such that $\Delta(G_{j})\leq O(\Delta/2^{i})$ and $\sum_{j=1}^{2^{i}}|E(G_{j})|=m$. Since the total running time at recursion level $i$ is $O(m\log^{2}n)$ and the depth of the recursion tree is $O(\log\Delta)$, it follows that the total running time is $O(m\log^{3}n)$ w.h.p. ∎
5 The Algorithm Construct-U-Fans (Proof of Lemma 4.1)
As input, the algorithm Construct-U-Fans is given a graph $G$ and a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges. It begins by taking the $\lambda$ uncolored edges in $\chi$ and using them to construct a separable collection $\operatorname*{\mathcal{U}}$ of $\lambda$ u-edges in an obvious way, which we describe in Lemma 5.14. It then uses two subroutines, Prune-Vizing-Fans and Reduce-U-Edges, to reduce the number of u-edges in $\operatorname*{\mathcal{U}}$, either by coloring them or modifying $\chi$ to turn them into u-fans. More specifically, for each color $\alpha\in[\Delta+1]$, the algorithm considers the collection $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ of u-edges that are $\alpha$-primed and (1) calls the subroutine Prune-Vizing-Fans to ensure that the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ have vertex-disjoint Vizing fans by either coloring the u-edges with overlapping Vizing fans or using them to create u-fans, and (2) calls the subroutine Reduce-U-Edges which either extends the coloring to the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ or uses them to create u-fans. Algorithm 3 gives the pseudocode for this algorithm.
1 Construct a separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$
2 for each $\alpha\in[\Delta+1]$ do
3       $\mathcal{F}\leftarrow\textnormal{{Prune-Vizing-Fans}}(\operatorname*{\mathcal{U}},\alpha)$
[…]
5
return $\operatorname*{\mathcal{U}}$
Algorithm 3 Construct-U-Fans
Organization of Section 5: We first describe and analyze the subroutines Prune-Vizing-Fans and Reduce-U-Edges used by the algorithm Construct-U-Fans before proving Lemma 4.1 in Section 5.3.
5.1 The Subroutine Prune-Vizing-Fans
As input, the subroutine Prune-Vizing-Fans is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, a color $\alpha$ and a separable collection $\operatorname*{\mathcal{U}}$ with $\lambda_{\alpha}$ $\alpha$-primed u-edges. We let $\operatorname*{\mathcal{E}}(\operatorname*{\mathcal{U}})\subseteq\operatorname*{\mathcal{U}}$ denote the set of u-edges in $\operatorname*{\mathcal{U}}$ and $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})\subseteq\operatorname*{\mathcal{E}}(\operatorname*{\mathcal{U}})$ denote the set of u-edges in $\operatorname*{\mathcal{U}}$ that are $\alpha$-primed. The subroutine then modifies the coloring $\chi$ and collection $\operatorname*{\mathcal{U}}$ to ensure that the Vizing fans of the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ are vertex-disjoint.
Algorithm Description: The subroutine Prune-Vizing-Fans scans through the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and constructs $\operatorname*{\mathcal{U}}$-avoiding Vizing fans at these u-edges (see Lemma 3.8), maintaining a vertex-disjoint subset $\mathcal{F}$ of these Vizing fans throughout the process. After constructing a Vizing fan $\operatorname*{\mathbf{F}}$ for a u-edge $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$, the subroutine checks if $F$ intersects some other fan in $\mathcal{F}$. If not, it adds $\operatorname*{\mathbf{F}}$ to $\mathcal{F}$. Otherwise, it uses these intersecting Vizing fans to either extend the coloring $\chi$ or construct a u-fan, removing the corresponding Vizing fans and u-edges from $\mathcal{F}$ and $\operatorname*{\mathcal{U}}$ respectively.
More formally, let $\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{\ell}$ denote the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ at the time when the subroutine is called and let $\operatorname*{\boldsymbol{e}}_{i}=(u_{i},v_{i},\alpha)$. Initialize an empty set of Vizing fans $\mathcal{F}\leftarrow\varnothing$. Then, for each $i\in[\ell]$, the subroutine constructs a $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}_{i}$ of $\operatorname*{\boldsymbol{e}}_{i}$ and checks if any of the vertices in $\operatorname*{\mathbf{F}}_{i}$ (i.e. any of the leaves or the center of $\operatorname*{\mathbf{F}}_{i}$) are contained in any of the Vizing fans in $\mathcal{F}$. Note that, since $\operatorname*{\mathcal{U}}$ is separable and $u_{1},\dots,u_{\ell}$ are the centers of $\alpha$-primed u-edges, these vertices are all distinct. If $\operatorname*{\mathbf{F}}_{i}$ is vertex-disjoint from the Vizing fans in $\mathcal{F}$, we add $\operatorname*{\mathbf{F}}_{i}$ to $\mathcal{F}$. Otherwise, consider the following 3 cases.
  • If $u_{i}$ is a leaf in $\operatorname*{\mathbf{F}}_{j}$ for some $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$: Then we rotate the fan $\operatorname*{\mathbf{F}}_{j}$ so that $(u_{i},u_{j})$ is uncolored, set $\chi(u_{i},u_{j})\leftarrow\alpha$, remove $\operatorname*{\mathbf{F}}_{j}$ from $\mathcal{F}$ and remove $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$.
If this is not the case, then some leaf of $\operatorname*{\mathbf{F}}_{i}$ appears in a Vizing fan in $\mathcal{F}$. Let $w$ be the first such leaf in $\operatorname*{\mathbf{F}}_{i}$.
  • If $w=u_{j}$ for some $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$: Then we rotate the fan $\operatorname*{\mathbf{F}}_{i}$ so that $(u_{i},w)$ is uncolored, set $\chi(u_{i},w)\leftarrow\alpha$, remove $\operatorname*{\mathbf{F}}_{j}$ from $\mathcal{F}$ and remove $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$.
  • If $w$ is a leaf in $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$: Then we rotate the fans $\operatorname*{\mathbf{F}}_{i}$ and $\operatorname*{\mathbf{F}}_{j}$ so that $(u_{i},w)$ and $(u_{j},w)$ are uncolored, create a u-fan $\operatorname*{\boldsymbol{f}}=(w,u_{i},u_{j},\beta,\alpha)$ for an arbitrary $\beta\in\mathsf{miss}_{\chi}(w)\setminus C_{\operatorname*{\mathcal{U}}}(w)$, remove $\operatorname*{\mathbf{F}}_{j}$ from $\mathcal{F}$, add $\operatorname*{\boldsymbol{f}}$ to $\operatorname*{\mathcal{U}}$ and remove $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. [^7] Note that, since the vertices $u_{1},\dots,u_{\ell}$ are all distinct, these cases are exhaustive.
The subroutine then returns the set $\mathcal{F}$ of vertex-disjoint Vizing fans.
Analysis of Prune-Vizing-Fans
The following lemmas summarize the main properties of the subroutine Prune-Vizing-Fans.
Lemma 5.1.
Prune-Vizing-Fans returns a set $\mathcal{F}$ of vertex-disjoint $\operatorname*{\mathcal{U}}$-avoiding Vizing fans for all of the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$.
Proof.
We can show by induction that, after the subroutine finishes scanning the u-edge $\operatorname*{\boldsymbol{e}}_{i}$, the set $\mathcal{F}$ consists of vertex-disjoint $\operatorname*{\mathcal{U}}$-avoiding Vizing fans for the u-edges in $\{\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{i}\}\cap\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. [^8] Since we remove edges from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ throughout this process, $\{\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{i-1}\}$ might not be contained in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. This is true trivially for $i=1$. Now, suppose that this is true for $i-1$. If the $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}_{i}$ of $\operatorname*{\boldsymbol{e}}_{i}$ does not intersect any of the Vizing fans in $\mathcal{F}$, then we add $\operatorname*{\mathbf{F}}_{i}$ to $\mathcal{F}$ and are done.
Thus, we assume this is not the case. Let $x_{0}$ denote $u_{i}$ and $x_{1},\dots,x_{k}$ denote the leaves of the Vizing fan $\operatorname*{\mathbf{F}}_{i}$ (in order). Let $x_{p}$ be the first vertex in the sequence $x_{0},x_{1},\dots,x_{k}$ which appears in some Vizing fan $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$. It follows from our algorithm that we remove $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ while only affecting the palettes of vertices in $\operatorname*{\mathbf{F}}_{j}$ and $x_{0},\dots,x_{p}$ by rotating the fans $\operatorname*{\mathbf{F}}_{i}$ and $\operatorname*{\mathbf{F}}_{j}$. Since all of the Vizing fans in $\mathcal{F}$ were vertex-disjoint, changing the palettes of vertices in $\operatorname*{\mathbf{F}}_{j}$ cannot affect the palettes of vertices in any of the Vizing fans in $\mathcal{F}\setminus\{\operatorname*{\mathbf{F}}_{j}\}$, and hence they are still valid. Similarly, none of the vertices $x_{0},\dots,x_{p}$ are contained in any of the Vizing fans in $\mathcal{F}\setminus\{\operatorname*{\mathbf{F}}_{j}\}$, so changing their palettes does not affect the validity of any other Vizing fans in $\mathcal{F}$. Thus, after removing $\operatorname*{\mathbf{F}}_{j}$ from $\mathcal{F}$, removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and making the changes to $\chi$, the set $\mathcal{F}$ consists of vertex-disjoint Vizing fans of the u-edges in $\{\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{i}\}\cap\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. To see why all of these Vizing fans also remain $\operatorname*{\mathcal{U}}$-avoiding, note that any u-fan $\operatorname*{\boldsymbol{f}}$ added to $\operatorname*{\mathcal{U}}$ is vertex-disjoint from these Vizing fans by an analogous argument. ∎
Lemma 5.2.
The subroutine Prune-Vizing-Fans maintains the invariant that $\operatorname*{\mathcal{U}}$ is separable.
Proof.
We can show by induction that, after the subroutine finishes scanning the u-edge $\operatorname*{\boldsymbol{e}}_{i}$, the collection $\operatorname*{\mathcal{U}}$ is separable. This is true trivially for $i=1$, since the subroutine does not change $\chi$ or $\operatorname*{\mathcal{U}}$ while scanning this u-edge. Now, suppose that this is true for $i-1$. It follows from the inductive argument in Lemma 5.1 that the Vizing fans currently in $\mathcal{F}$ are all $\operatorname*{\mathcal{U}}$-avoiding. If we do not change $\chi$ or $\operatorname*{\mathcal{U}}$ while scanning $\operatorname*{\boldsymbol{e}}_{i}$, then we are done. Thus, we assume that this is not the case. The subroutine can modify $\chi$ and $\operatorname*{\mathcal{U}}$ in the following ways:
  • Rotating colors around some $\operatorname*{\mathbf{F}}_{j}$ so that $(u_{j},u_{i})$ is uncolored, setting $\chi(u_{j},u_{i})\leftarrow\alpha$ and removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$.
  • Rotating colors around $\operatorname*{\mathbf{F}}_{i}$ so that $(u_{i},u_{j})$ is uncolored for some $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$, setting $\chi(u_{j},u_{i})\leftarrow\alpha$ and removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$.
  • Rotating colors around $\operatorname*{\mathbf{F}}_{i}$ and some $\operatorname*{\mathbf{F}}_{j}\in\mathcal{F}$ so that $(u_{i},u_{j})$ is uncolored, removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$ and adding a u-fan $\operatorname*{\boldsymbol{f}}$ to $\operatorname*{\mathcal{U}}$.
In the first case, it follows from Lemma 3.9 (since the Vizing fans are all $\operatorname*{\mathcal{U}}$-avoiding) that rotating colors around $\operatorname*{\mathbf{F}}_{j}$ does not damage any u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}_{j}\}$. Furthermore, since $\operatorname*{\mathcal{U}}$ is separable, setting $\chi(u_{j},u_{i})\leftarrow\alpha$ does not damage any u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}_{i},\operatorname*{\boldsymbol{e}}_{j}\}$. Since we remove $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$, we ensure that $\operatorname*{\mathcal{U}}$ remains separable. The same argument extends to the second case.
Finally, for the third case, it follows from Lemma 3.9 that rotating colors around $\operatorname*{\mathbf{F}}_{i}$ and $\operatorname*{\mathbf{F}}_{j}$ does not damage an u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}_{i},\operatorname*{\boldsymbol{e}}_{j}\}$, so after removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$ we have that $\operatorname*{\mathcal{U}}$ is separable. Since we have that $\alpha\in C_{\operatorname*{\mathcal{U}}}(u_{i})\cap C_{\operatorname*{\mathcal{U}}}(u_{j})$ after removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$, we can see that $\operatorname*{\boldsymbol{f}}$ is indeed a u-fan and that $\operatorname*{\mathcal{U}}\cup\{\operatorname*{\boldsymbol{f}}\}$ is separable. [^9] Note that, before removing $\operatorname*{\boldsymbol{e}}_{i}$ and $\operatorname*{\boldsymbol{e}}_{j}$ from $\operatorname*{\mathcal{U}}$, $c_{\operatorname*{\boldsymbol{e}}_{i}}(u_{i})=c_{\operatorname*{\boldsymbol{e}}_{j}}(u_{j})=\alpha$.
Lemma 5.3.
Each time Prune-Vizing-Fans modifies the coloring $\chi$, it removes at most $2$ u-edges from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and either (1) extends the coloring $\chi$ to one more edge, or (2) adds a u-fan to $\operatorname*{\mathcal{U}}$.
Proof.
This follows immediately from the description of the subroutine. ∎
The following lemma shows that this subroutine can be implemented efficiently.
Lemma 5.4.
We can implement Prune-Vizing-Fans to run in $O(\Delta\lambda_{\alpha})$ time.
Proof.
By Lemma 3.8, we can construct each of the $\operatorname*{\mathcal{U}}$-avoiding Vizing fans $\operatorname*{\mathbf{F}}_{i}$ in $O(\Delta)$ time. As we scan through the u-edges $\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{\ell}$, we can maintain the set of vertices $S$ that are contained in the Vizing fans in $\mathcal{F}$, along with pointers to the corresponding Vizing fan. This allows us to either verify that $\operatorname*{\mathbf{F}}_{i}$ is vertex-disjoint from the Vizing fans in $\mathcal{F}$ or to find the first vertex in $\operatorname*{\mathbf{F}}_{i}$ that is contained in one of these fans in $O(\Delta)$ time. We can then appropriately update the coloring $\chi$ and the collection $\operatorname*{\mathcal{U}}$ in $O(\Delta)$ time using the data structures outlined in Section 3.2. It follows that the entire subroutine takes $O(\Delta\lambda_{\alpha})$ time. ∎
5.2 The Subroutine Reduce-U-Edges
As input, the subroutine Reduce-U-Edges is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, a color $\alpha$, a separable collection $\operatorname*{\mathcal{U}}$ with $\lambda_{\alpha}$ $\alpha$-primed u-edges and a set $\mathcal{F}$ of vertex-disjoint $\operatorname*{\mathcal{U}}$-avoiding Vizing fans for all of the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. Let $m_{\alpha}$ denote the number of edges with color $\alpha$ in the input coloring $\chi$. For each u-edge $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$, let $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}$ denote its Vizing fan in $\mathcal{F}$. Similarly, for each $\operatorname*{\mathbf{F}}\in\mathcal{F}$, let $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ denote its corresponding u-edge in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. The subroutine then either extends the coloring to the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ or uses them to construct u-fans.
[…]
More formally, the subroutine begins by scanning through all of the trivial Vizing fans $\operatorname*{\mathbf{F}}\in\mathcal{F}$ and calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$, removing the $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ from $\operatorname*{\mathcal{U}}$, and removing any other u-component damaged by calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ from $\operatorname*{\mathcal{U}}$ (see Lemma 3.9). Let $L\leftarrow 0$, $S\leftarrow\varnothing$ and $P_{\operatorname*{\mathbf{F}}}\leftarrow\varnothing$ for all $\operatorname*{\mathbf{F}}\in\mathcal{F}$. The subroutine then proceeds in rounds which consist of updating each of the paths $P_{\operatorname*{\mathbf{F}}}$ while maintaining the following invariant.
Invariant 5.5.
After updating a path $P_{\operatorname*{\mathbf{F}}}$, the following hold:
  1. 1.
[…]
  2. 2.
    The prefix paths $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$ are all edge-disjoint.
  3. 3.
    The set $S$ is the union of all the edges in the prefix paths $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$.
  4. 4.
    The collection $\operatorname*{\mathcal{U}}$ is separable and $|\mathcal{F}|=|\mathcal{E}_{\alpha}(\operatorname*{\mathcal{U}})|$.
  5. 5.
    The Vizing fans in $\mathcal{F}$ are $\operatorname*{\mathcal{U}}$-avoiding and vertex-disjoint.
The prefix paths $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$ maintained by the algorithm also satisfy the following invariant.
Invariant 5.6.
At the start of each round, we have that $|P_{\operatorname*{\mathbf{F}}}|=L$ for all $\operatorname*{\mathbf{F}}\in\mathcal{F}$.
[…]
1 $P_{\operatorname*{\mathbf{F}}}\leftarrow\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})_{\leq|P_{\operatorname*{\mathbf{F}}}|+1}$
2 Let $(x,y)$ be the $(L+1)^{th}$ edge in $P_{\operatorname*{\mathbf{F}}}$ and $y$ be an endpoint of $P_{\operatorname*{\mathbf{F}}}$
3 if $(x,y)\in S$ then
4       Let $\operatorname*{\mathbf{F}}^{\prime}\in\mathcal{F}$ be the Vizing fan such that $(x,y)\in P_{\operatorname*{\mathbf{F}}^{\prime}}$
5       if $(x,y)$ appears in the same orientation in $P_{\operatorname*{\mathbf{F}}}$ and $P_{\operatorname*{\mathbf{F}}^{\prime}}$ then
6             Let $(z,x)$ and $(z^{\prime},x)$ be the edges appearing before $(x,y)$ in $P_{\operatorname*{\mathbf{F}}}$ and $P_{\operatorname*{\mathbf{F}}^{\prime}}$ respectively
7             $\beta\leftarrow\chi(z,x)$
8             $\chi(z,x)\leftarrow\bot$ and $\chi(z^{\prime},x)\leftarrow\bot$
9             $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$
10             $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}}^{\prime})$
11             Remove $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ from $\operatorname*{\mathcal{U}}$
12             Add the u-fan $\operatorname*{\boldsymbol{f}}=(x,z,z^{\prime},\beta,\alpha)$ to $\operatorname*{\mathcal{U}}$
13
14      else
15             $\chi(x,y)\leftarrow\bot$
16             $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$
17             $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}}^{\prime})$
18             Remove $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ from $\operatorname*{\mathcal{U}}$
19
20      $S\leftarrow S\setminus(P_{\operatorname*{\mathbf{F}}}\cup P_{\operatorname*{\mathbf{F}}^{\prime}})$
21       return
[…]
23 if $P_{\operatorname*{\mathbf{F}}}$ is maximal  then
24       $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$
25       Remove $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ from $\operatorname*{\mathcal{U}}$
26       $S\leftarrow S\setminus P_{\operatorname*{\mathbf{F}}}$
27       if there exists $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$ with $c_{\operatorname*{\boldsymbol{g}}}(y)=\chi(x,y)$ then
28             Remove $\operatorname*{\boldsymbol{g}}$ from $\operatorname*{\mathcal{U}}$ (see Lemma 3.9)
29
30      if there exists $\operatorname*{\mathbf{F}}^{\prime}\in\mathcal{F}$ such that $y\in\operatorname*{\mathbf{F}}$ then
31             Remove $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ from $\operatorname*{\mathcal{U}}$ // $\operatorname*{\boldsymbol{e}}_{F^{\prime}}$ might also get removed in Algorithm 4
32             $S\leftarrow S\setminus P_{\operatorname*{\mathbf{F}}^{\prime}}$
33
[…]
Analysis of Reduce-U-Edges
The following lemmas summarize the main properties of the subroutine Reduce-U-Edges.
Lemma 5.7.
The subroutine Reduce-U-Edges satisfies 5.5.
Proof.
[…]
After activating and removing all of the trivial Vizing fans in $\mathcal{F}$, Conditions 1-3 are clearly still satisfied (since the paths are all empty). For Condition 4, note that the subroutine removes any damaged u-components from $\operatorname*{\mathcal{U}}$, removes $\operatorname*{\boldsymbol{e}}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ if it removes $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}$ from $\mathcal{F}$, and that activating some trivial $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}\in\mathcal{F}$ cannot damage any $\operatorname*{\boldsymbol{e}}^{\prime}\in\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})\setminus\{\operatorname*{\boldsymbol{e}}\}$ since their Vizing fans are vertex-disjoint. For Condition 5, note that the subroutine does not add any u-component to $\operatorname*{\mathcal{U}}$, so the vertex-disjoint Vizing fans remain $\operatorname*{\mathcal{U}}$-avoiding. This establishes the base case of the induction.
[…]
Conditions 2 and 3: After extending $P_{\operatorname*{\mathbf{F}}}$ by one more edge, we check if there is some other path $P_{\operatorname*{\mathbf{F}}^{\prime}}$ that intersects the updated path $P_{\operatorname*{\mathbf{F}}}$ at this new edge. If so, we remove both $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$, ensuring that the remaining paths are edge-disjoint. Otherwise, the paths are all edge-disjoint. It’s straightforward to verify that $S$ is updated correctly in each case.
Condition 4: We first show that the collection $\operatorname*{\mathcal{U}}$ remains separable. If this call to Update-Path does not change $\chi$ or $\operatorname*{\mathcal{U}}$, then clearly $\operatorname*{\mathcal{U}}$ remains separable. Note that the Vizing fans in $\mathcal{F}$ are all $\operatorname*{\mathcal{U}}$-avoiding and recall Lemma 3.9. We now consider the following three cases.
(1) If we enter the if statement on Algorithm 4, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ and $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}}^{\prime})$ after uncoloring edges on their Vizing chains does not damage any u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$, which are both removed from $\operatorname*{\mathcal{U}}$. This is because shifting colors around the Vizing fans $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ does not damage any u-components apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ (see Lemma 3.9) and truncating an alternating path before flipping it ensures that we do not remove any colors from the palettes of the vertices on that path. Finally, note that the u-fan $\operatorname*{\boldsymbol{f}}$ that we add to $\operatorname*{\mathcal{U}}$ only uses colors that were previously unavailable. (2) If we enter the else statement on Algorithm 4, then the argument is completely analogous to the previous case, except that we do not add any u-components to $\operatorname*{\mathcal{U}}$ (see Lemma 3.9). (3) If we enter the if statement on Algorithm 4, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ damages at most one u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ (see Lemma 3.9). In particular, if there is such a u-component $\operatorname*{\boldsymbol{g}}$, it will contain the vertex $y$ and have $c_{\operatorname*{\boldsymbol{g}}}(x)=\chi(x,y)$. Thus, we check if such a u-component exists and remove it from $\operatorname*{\mathcal{U}}$.
Finally, to see that $|\mathcal{F}|=|\mathcal{E}_{\alpha}(\operatorname*{\mathcal{U}})|$, we can verify that we remove $\operatorname*{\boldsymbol{e}}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ if and only if we remove $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}$ from $\mathcal{F}$. It’s clear that whenever we remove some $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$, we also remove $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$. Similarly, if we remove a u-edge $\operatorname*{\boldsymbol{e}}$ from $\mathcal{E}_{\alpha}(\operatorname*{\mathcal{U}})$ in Algorithm 4, then we can see that we then remove $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}$ from $\mathcal{F}$ in Algorithm 4 immediately afterwards.
Condition 5: Let $\operatorname*{\mathbf{F}}^{\prime}$ be some Vizing fan that remains in $\mathcal{F}$ after handling this call. Then, for any Vizing fan $\operatorname*{\mathbf{F}}^{\prime\prime}$ that was activated during the call, we know that its corresponding Vizing chain did not end at $\operatorname*{\mathbf{F}}^{\prime}$, otherwise $\operatorname*{\mathbf{F}}^{\prime}$ would have been removed from $\mathcal{F}$ (see Algorithm 4). Thus, activating $\operatorname*{\mathbf{F}}^{\prime\prime}$ (which is vertex-disjoint from $\operatorname*{\mathbf{F}}^{\prime}$) does not change the palettes of any vertices in $\operatorname*{\mathbf{F}}^{\prime}$. Note that, since the center of $\operatorname*{\mathbf{F}}^{\prime}$ is missing $\alpha$, it is not possible for the Vizing chain of $\operatorname*{\mathbf{F}}^{\prime\prime}$ to contain any edges in $\operatorname*{\mathbf{F}}^{\prime}$ without ending at $\operatorname*{\mathbf{F}}^{\prime}$. Hence, the Vizing fans remaining in $\mathcal{F}$ are still valid Vizing fans and remain vertex-disjoint. To see that they also remain $\operatorname*{\mathcal{U}}$-avoiding, note that if we add a u-fan $\operatorname*{\boldsymbol{f}}$ to $\operatorname*{\mathcal{U}}$ (see Algorithm 4), the color that $\operatorname*{\boldsymbol{f}}$ assigns to a vertex $x\in\operatorname*{\boldsymbol{f}}$ was previously not available at $x$ before the start of the call to Update-Path. Thus, if $x$ is contained in the Vizing fan $\operatorname*{\mathbf{F}}^{\prime}$, its color within $\operatorname*{\mathbf{F}}^{\prime}$ is not $c_{\operatorname*{\boldsymbol{f}}}(x)$.
[…]
Lemma 5.8.
The subroutine Reduce-U-Edges satisfies 5.6.
Proof.
[…]
Lemma 5.9.
Each time Reduce-U-Edges modifies the coloring $\chi$, it removes at most $2$ u-edges from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and either (1) extends the coloring $\chi$ to one more edge and removes at most one other u-component from $\operatorname*{\mathcal{U}}$, or (2) adds a u-fan to $\operatorname*{\mathcal{U}}$.
Proof.
[…]
Running Time: We now show how to implement this subroutine and analyse its running time. We begin with the following claim which shows that Update-Path can be implemented efficiently.
Claim 5.10.
[…]
Proof.
[…]
If this call does not remove $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$, then we know that $P_{\operatorname*{\mathbf{F}}}$ is not maximal and is also edge-disjoint from the other prefix paths maintained by the subroutine. In this case, we update $S$ in $O(1)$ time and do not modify $\chi$ or $\operatorname*{\mathcal{U}}$.
On the other hand, if this call does remove $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$, then we need to activate $O(1)$ many Vizing chains of length $O(L)$, which can be done in $O(\Delta+L)$ time. Removing the edges of the corresponding paths from $S$ can also be done in $O(L)$ time. Finally, using the data structures outlined in Section 3.2, we can update $\mathcal{F}$ and $\operatorname*{\mathcal{U}}$ in $O(1)$ time. ∎
Let $\mathcal{F}^{\star}$ denote the set $\mathcal{F}$ when we first call the subroutine Reduce-U-Edges. For each $\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}$, let $L_{\operatorname*{\mathbf{F}}}$ denote the value of $L$ at the time that $\operatorname*{\mathbf{F}}$ is removed from $\mathcal{F}$.
Claim 5.11.
The total time spent handling calls to Update-Path is at most
$O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}L_{\operatorname*{\mathbf{F}}}.$ (1)
Proof.
[…]
$\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}O(\Delta+L_{\operatorname*{\mathbf{F}}})\leq O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}L_{\operatorname*{\mathbf{F}}}.\qed$
Now, assume w.l.o.g. that =$\mathcal{F}^{\star}=\{\operatorname*{\mathbf{F}}_{1},\dots,\operatorname*{\mathbf{F}}_{\lambda_{\alpha}}\}$ and $L_{\lambda_{\alpha}}\leq\dots\leq L_{1}$, where $L_{i}:=L_{\operatorname*{\mathbf{F}}_{i}}$. Recall that $m_{\alpha}$ denotes the number of edges with color $\alpha$ when we first call the subroutine.
Claim 5.12.
For each $i\in[\lambda_{\alpha}]$, we have that $L_{i}\leq 4(m_{\alpha}+2\lambda_{\alpha})/i$.
Proof.
Let $i\in[\lambda_{\alpha}]$, $\tau=4(m_{\alpha}+2\lambda_{\alpha})/i$ and suppose that $\tau<L_{i}$. Then, at the start of iteration $\tau+1$, the Vizing fans $\operatorname*{\mathbf{F}}_{1},\dots\operatorname*{\mathbf{F}}_{i}$ are all contained in $\mathcal{F}$. Furthermore, it follows from Invariants 5.5 and 5.6 that, at the start of iteration $\tau+1$, the alternating paths $P_{\operatorname*{\mathbf{F}}_{1}},\dots,P_{\operatorname*{\mathbf{F}}_{i}}$ form a collection of $i$ edge-disjoint $\{\alpha,\cdot\}$-alternating paths of length $\tau$ in $G$. Let $T$ denote the total length of these paths. We can observe that $T$ is at most $3m_{\alpha}^{\prime}$, where $m_{\alpha}^{\prime}$ is the number of edges that currently have color $\alpha$, since at least a third of the edges in each of these paths has color $\alpha$ (note that any $\{\alpha,\cdot\}$-alternating path of length $k$ has at least $\lfloor k/2\rfloor$ edges with color $\alpha$). We can also observe that $m_{\alpha}^{\prime}\leq m_{\alpha}+2\lambda_{\alpha}$ since the number of edges with color $\alpha$ increases by at most $2$ each time we activate a Vizing fan in $\mathcal{F}$. Thus, it follows that
$4(m_{\alpha}+2\lambda_{\alpha})=i\tau\leq T\leq 3m_{\alpha}^{\prime}\leq 3(m_{\alpha}+2\lambda_{\alpha}),$
giving a contradiction. ∎
Lemma 5.13.
We can implement Reduce-U-Edges to run in $O((m_{\alpha}+\Delta\lambda_{\alpha})\log n)$ time.
Proof.
Activating the trivial Vizing fans in $\mathcal{F}$ when the subroutine is first called can be done in $O(\Delta\lambda_{\alpha})$ time. The running time of the rest of the subroutine is dominated by the time taken to handle the call to Update-Path. Combining Claims 5.11 and 5.12, it follows that the total time spent handling calls to Update-Path is at most
$O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}L_{\operatorname*{\mathbf{F}}}\leq O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{i=1}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}$
$\leq O((m_{\alpha}+\Delta\lambda_{\alpha})\log n).\qed$
5.3 Analysis of Construct-U-Fans (Proof of Lemma 4.1)
The algorithm begins by constructing a separable collection of $\lambda$ u-components $\operatorname*{\mathcal{U}}$. The following lemma shows that this can be done efficiently.
Lemma 5.14.
Given a graph $G$ and a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges, we can construct a separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$ in $O(m)$ time.
Proof.
The algorithm first initializes an empty separable collection $\operatorname*{\mathcal{U}}=\varnothing$. It then scans through the edges of the graph and retrieves the $\lambda$ edges $e_{1},\dots,e_{\lambda}$ that are uncolored by $\chi$. The algorithm then scans through each of these uncolored edges and, for each $e_{i}=(u_{i},v_{i})$, picks a missing color $\alpha_{i}\in\mathsf{miss}_{\chi}(u_{i})\setminus C_{\operatorname*{\mathcal{U}}}(u_{i})$ (see 3.5) and adds the u-edge $\operatorname*{\boldsymbol{e}}_{i}:=(u_{i},v_{i},\alpha_{i})$ to $\operatorname*{\mathcal{U}}$. It’s easy to verify that the resulting collection $\operatorname*{\mathcal{U}}$ is separable and contains $\lambda$ u-edges. Using the data structures outlined in Section 3.2, finding such a color and adding a u-edge to $\operatorname*{\mathcal{U}}$ can be done in $O(1)$ time. Thus, this entire process can be implemented in $O(m)$ time. ∎
Let $m_{\alpha}$ denote the number of edges that have color $\alpha$ w.r.t. the initial coloring $\chi$ when we first call Construct-U-Fans. The following claim shows that we cannot create too many more edges with the color $\alpha$ throughout this sequence of calls to the subroutines Prune-Vizing-Fans and Reduce-U-Edges for each color $\alpha\in[\Delta+1]$.
Claim 5.15.
For each color $\alpha\in[\Delta+1]$, we have that at most $m_{\alpha}+O(\lambda)$ edges have color $\alpha$ throughout the entire run of the algorithm Construct-U-Fans.
Proof.
This follows from the fact that the number of edges with color $\alpha$ can only increase when we extend the coloring $\chi$ to some uncolored edge. Furthermore, it can only increase by at most $2$ every time this happens. Thus, for any $\alpha\in[\Delta+1]$, the number of edges with color $\alpha$ can increase by at most $\lambda$ throughout the entire run of the algorithm Construct-U-Fans. ∎
Let $\lambda_{\alpha}$ denote the number of $\alpha$-primed u-edges in the initial collection $\operatorname*{\mathcal{U}}$. Note that the number of such u-edges can only decrease throughout the run of this algorithm. It follows from Lemmas 5.4 and 5.13 that the total running time of the algorithm Construct-U-Fans across all of these calls to the subroutines Prune-Vizing-Fans and Reduce-U-Edges is at most
$\sum_{\alpha\in[\Delta+1]}\left(O(\lambda_{\alpha}\Delta)+O((m_{\alpha}+\lambda+\lambda_{\alpha}\Delta)\log n)\right)\leq O((m+\Delta\lambda)\log n).$
Lemma 5.16.
The algorithm Construct-U-Fans either extends the coloring $\chi$ to at least $\lambda/9$ more edges or returns a separable collection of at least $\lambda/9$ u-fans $\operatorname*{\mathcal{U}}$.
Proof.
To see why this lemma holds, consider the following three quantities and how they evolve over time throughout the execution of the algorithm Construct-U-Fans: The number of edges that the algorithm has extended the coloring to so far, $\Psi_{\texttt{c}}$, the number of u-edges in $\operatorname*{\mathcal{U}}$, $\Psi_{\texttt{e}}:=|\mathcal{E}|$, and the number of u-fans in $\operatorname*{\mathcal{U}}$, $\Psi_{\texttt{f}}:=|\mathcal{U}\setminus\mathcal{E}|$. Immediately after constructing the separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$, it holds that $\Psi_{\texttt{c}}=0$, $\Psi_{\texttt{e}}=\lambda$ and $\Psi_{\texttt{f}}=0$. We next argue that by the time the execution of the algorithm has finished, either $\Psi_{\texttt{c}}\geq\lambda/9$ or $\Psi_{\texttt{f}}\geq\lambda/9$ must hold. To this end, we employ the following potential function argument.
Claim 5.17.
The quantity $3(\Psi_{\textnormal{{f}}}+2\Psi_{\textnormal{{c}}})+\Psi_{\textnormal{{e}}}$ is non-decreasing throughout the algorithm’s execution.
Proof.
Consider first a call made by algorithm Construct-U-Fans to the Prune-Vizing-Fans subroutine. By Lemma 5.3, the subroutine repeatedly removes at most $2$ u-edges from $\operatorname*{\mathcal{U}}$ and either (1) adds a u-fan to $\operatorname*{\mathcal{U}}$, or (2) extends the coloring to another edge. This may decrease $\Psi_{\texttt{e}}$ by at most $2$ while increasing one of $\Psi_{\texttt{f}}$ or $\Psi_{\texttt{c}}$ by at least $1$, so $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}$ increases in this case.
Next, consider a call to the Reduce-U-Edges subroutine. By Lemma 5.9, during any iteration, the subroutine removes at most $3$ u-edges from $\operatorname*{\mathcal{U}}$ and either (1) adds a u-fan to $\operatorname*{\mathcal{U}}$, or (2) extends the coloring to another edge while removing at most one u-fan from $\operatorname*{\mathcal{U}}$. This may decrease $\Psi_{\texttt{e}}$ by at most $3$ while increasing $\Psi_{\texttt{f}}+2\Psi_{\texttt{c}}$ by at least $1$, so $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}$ cannot decrease in value. ∎
Immediately after constructing the separable collection of $\lambda$ u-edges in $\operatorname*{\mathcal{U}}$, we have $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}=\lambda$. 5.17 implies that $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}\geq\lambda$ holds at all times afterwards. Since the algorithm iterates until $\mathcal{E}=\varnothing$, we know that $\Psi_{\texttt{e}}=0$ when the algorithm’s execution terminates, so at that moment we have $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})\geq\lambda$. Hence, either $\Psi_{\texttt{f}}\geq\lambda/9$ or $\Psi_{\texttt{c}}\geq\lambda/9$ must hold. ∎
6 The Algorithm Color-U-Fans (Proof of Lemma 4.2)
As input, the algorithm Color-U-Fans is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, and a collection of separable u-fans $\mathcal{U}$ of size $\lambda$. It then uses two further subroutines, Prime-U-Fans and Activate-U-Fans, to prime u-fans with the same colors and then activate them. More specifically, the algorithm first identifies the two least common colors $\alpha,\beta\in[\Delta+1]$. It calls the subroutine Prime-U-Fans which proceeds to prime $\Omega(\lambda/\Delta)$ of the u-fans in $\operatorname*{\mathcal{U}}$ with the colors $\alpha$ and $\beta$ (i.e. to modify the coloring $\chi$ and these u-fans so that they are $\{\alpha,\beta\}$-primed). It then calls Activate-U-Fans which extends the coloring $\chi$ to $\Omega(\lambda/\Delta)$ of the edges in these u-fans. The algorithm repeats this process until it has colored $\Omega(\lambda)$ edges. Algorithm 5 gives the pseudocode for Color-U-Fans.
1 for $\Delta/2$ iterations do
2       Let $\alpha,\beta\in[\Delta+1]$ be the two least common colors in $\chi$
[…]
5
[…]
Organization of Section 6: We begin by describing and analyzing the subroutines Prime-U-Fans and Activate-U-Fans used by Color-U-Fans before proving Lemma 4.2 in Section 6.3.
6.1 The Subroutine Prime-U-Fans
As input, this subroutine is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, a collection of separable u-fans $\mathcal{U}$ of size $\lambda$, and the two least common colors $\alpha,\beta\in[\Delta+1]$ (see 6.1). It then proceeds in iterations, where in each iteration it samples a u-fan $\operatorname*{\boldsymbol{f}}$ from $\mathcal{U}$ uniformly at random and attempts to prime $\operatorname*{\boldsymbol{f}}$ with the colors $\alpha$ and $\beta$. The subroutine maintains a subset $\Phi\subseteq\operatorname*{\mathcal{U}}$ of $\{\alpha,\beta\}$-primed u-fans. The subroutine performs iterations until $|\Phi|=\Omega(\lambda/\Delta)$, after which we proceed to call the subroutine Activate-U-Fans to extend the coloring $\chi$ to edges in the u-fans in $\Phi$. The pseudocode in Algorithm 6 gives a formal description of the subroutine Prime-U-Fans.
1 $\Phi\leftarrow\varnothing$ and $\lambda\leftarrow|\mathcal{U}|$
2 while $|\Phi|<\lambda/(48\Delta)$ do
3       Sample $\operatorname*{\boldsymbol{f}}=(u,v,w,\gamma,\delta)\sim U$ independently and u.a.r.
4       $(\alpha^{\prime},\beta^{\prime})\leftarrow(\alpha,\beta)$     // $\alpha^{\prime}$ and $\beta^{\prime}$ ensure we never flip an $\{\alpha,\beta\}$-alt. path
5       if $\gamma=\beta$ or $\delta=\alpha$ then
6             $(\alpha^{\prime},\beta^{\prime})\leftarrow(\beta,\alpha)$
7
8      Let $P_{u}$ be the $\{\alpha^{\prime},\gamma\}$-alternating path starting at $u$
9       Let $P_{v}$ and $P_{w}$ be the $\{\beta^{\prime},\delta\}$-alternating paths starting at $v$ and $w$ respectively
10       $\textnormal{{cost}}(\operatorname*{\boldsymbol{f}}):=|P_{u}|+|P_{v}|+|P_{w}|$
11       if $\textnormal{{cost}}(\operatorname*{\boldsymbol{f}})\leq 128m/\lambda$ then
12             Let $S$ denote the set of endpoints of $P_{u}$, $P_{v}$ and $P_{w}$
13             if the vertices in $S\cup\{u,v,w\}$ are not in any u-fan in $\Phi$ then
14                   Flip the alternating paths $P_{u}$, $P_{v}$ and $P_{w}$
15                   Remove $\operatorname*{\boldsymbol{f}}$ and any damaged u-fans from $\mathcal{U}$
16                   Add the u-fan $(u,v,w,\alpha^{\prime},\beta^{\prime})$ to $\operatorname*{\mathcal{U}}$ and $\Phi$
17
18return $(\alpha,\beta)$
[…]
Analysis of Prime-U-Fans
We say that an iteration of (the while loop in) Prime-U-Fans is successful if it adds a u-fan to the set $\Phi$. Otherwise, we say that the iteration fails. Note that the algorithm repeatedly performs iterations until it has performed at least $\lambda/(48\Delta)$ successful iterations. The following lemmas summarize the main properties of the subroutine Prime-U-Fans.
Claim 6.1.
The total number of edges with color $\alpha$ or $\beta$ is $O(m/\Delta)$ at any point throughout the run of Prime-U-Fans.
Proof.
Since $\alpha$ and $\beta$ are initially the two least common colors, there are initially $O(m/\Delta)$ such edges. Since each successful iteration increases the number of such edges by $O(1)$ and there are at most $O(\lambda/\Delta)\leq O(m/\Delta)$ such iterations, the claim follows. ∎
Claim 6.2.
$|\mathcal{U}|\geq(1-1/(2\Delta))\lambda$ at any point throughout the run of Prime-U-Fans.
Proof.
Consider a successful iteration of the algorithm where we sample a u-fan $\operatorname*{\boldsymbol{f}}\in\mathcal{U}$. The algorithm removes $\operatorname*{\boldsymbol{f}}$ from $\mathcal{U}$, along with any other damaged u-fans in $\mathcal{U}$. It follows from Lemma 3.4 that flipping the colors of the alternating paths $P_{u}$, $P_{v}$ and $P_{w}$ damages at most $6$ u-fans in $\operatorname*{\mathcal{U}}$. Since our algorithm runs for at most $\lambda/(48\Delta)$ iterations, it follows that $|\mathcal{U}|\geq\lambda-7\cdot\lambda/(48\Delta)$. ∎
Lemma 6.3.
The u-fans in $\Phi$ are all vertex-disjoint and $\{\alpha,\beta\}$-primed at any point throughout the run of Prime-U-Fans.
Proof.
If the algorithm samples a u-fan $\operatorname*{\boldsymbol{f}}$ that shares a vertex $x$ with a u-fan $\operatorname*{\boldsymbol{f}}^{\prime}\in\Phi$, then the iteration fails. Thus, the u-fans in $\Phi$ are all vertex-disjoint.
Now, note that whenever we add a u-fan $(u,v,w,\alpha^{\prime},\beta^{\prime})$ to $\Phi$, we do this immediately after flipping the paths $P_{u}$, $P_{v}$, and $P_{w}$. Thus, we have that $\alpha^{\prime}\in\mathsf{miss}_{\chi}(u)$ and $\beta^{\prime}\in\mathsf{miss}_{\chi}(v)\cap\mathsf{miss}_{\chi}(w)$, so the u-fan is $\{\alpha,\beta\}$-primed when we add it to $\Phi$. Since we only flip the paths in Algorithm 6 if their endpoints do not touch any u-fans in $\Phi$, this operation cannot change what colors are available at vertices in u-fans in $\Phi$, and hence cannot change whether or not any u-fan in $\Phi$ is $\{\alpha,\beta\}$-primed. ∎
The following standard claim bounds the total length of all maximal $\{c,\cdot\}$-alternating paths.
Claim 6.4.
For any color $c\in[\Delta+1]$, the total length of all maximal $\{c,\cdot\}$-alternating paths in $\chi$ is at most $4m$.
Proof.
Let $\mathcal{P}_{c}$ denote the set of all such alternating paths. First note that the total length of all alternating paths in $\mathcal{P}_{c}$ with length $1$ is at most $m$. Now, let $c^{\prime}\in[\Delta+1]\setminus\{c\}$ and let $P$ be a maximal $\{c,c^{\prime}\}$-alternating path with $|P|\geq 2$. We can observe that at least a third of the edges in $P$ have color $c^{\prime}$, and that each edge with color $c^{\prime}$ only appears in one path in $\mathcal{P}_{c}$. Thus, we have that
$\sum_{P\in\mathcal{P}_{c}}|P|\leq m+3\cdot|\{e\in E\mid\chi(e)\neq c\}|\leq 4m.\qed$
Lemma 6.5.
Each iteration of Prime-U-Fans succeeds with probability at least $1/4$.
Proof.
Consider the state of the algorithm at the start of some iteration. For each u-fan $\operatorname*{\boldsymbol{f}}^{\prime}\in\operatorname*{\mathcal{U}}$, let $P_{x}(\operatorname*{\boldsymbol{f}}^{\prime})$ denote the alternating path starting at $x\in\operatorname*{\boldsymbol{f}}^{\prime}$ that is considered by the algorithm if it samples $\operatorname*{\boldsymbol{f}}^{\prime}\in\operatorname*{\mathcal{U}}$. We define the cost of the u-fan $\operatorname*{\boldsymbol{f}}^{\prime}$ to be
$\textnormal{{cost}}(\operatorname*{\boldsymbol{f}}{\!{}^{\prime}}):=\sum_{x\in\operatorname*{\boldsymbol{f}}^{\prime}}|P_{x}(\operatorname*{\boldsymbol{f}}{\!{}^{\prime}})|.$
Let $\mathcal{U}^{\star}\subseteq\mathcal{U}$ denote the subset of u-fans $\operatorname*{\boldsymbol{f}}^{\prime}\in\operatorname*{\mathcal{U}}$ such that none of the alternating paths $\{P_{x}(\operatorname*{\boldsymbol{f}}^{\prime})\}_{x\in\operatorname*{\boldsymbol{f}}^{\prime}}$ have endpoints at some u-fan in $\Phi$. We can see that the iteration is successful if and only if the u-fan $\operatorname*{\boldsymbol{f}}\in\operatorname*{\mathcal{U}}$ sampled during the iteration satisfies $\textnormal{{cost}}(\operatorname*{\boldsymbol{f}})\leq 128m/\lambda$ (see Algorithm 6) and $\operatorname*{\boldsymbol{f}}\in\operatorname*{\mathcal{U}}^{\star}$ (see Algorithm 6). Thus, we now show that this happens with probability at least $1/4$.
We first begin with the following claim.
Claim 6.6.
Let $P$ be an $\{\alpha,\cdot\}$- or $\{\beta,\cdot\}$-alternating path in $\chi$. Then there are at most $2$ u-fans $\operatorname*{\boldsymbol{f}}^{\prime}\in\mathcal{U}$ that, if sampled during the iteration, might cause the algorithm to flip $P$.
Proof.
Suppose that the path $P$ is a $\{c,c^{\prime}\}$-alternating path for some colors $c\in\{\alpha,\beta\}$ and $c^{\prime}\notin\{\alpha,\beta\}$ with endpoints $x$ and $y$ [^10] Note that the algorithm never flips $\{\alpha,\beta\}$-alternating paths. Since the collection of u-fans $\mathcal{U}$ is separable, we know that at most one u-fan in $\mathcal{U}$ containing $x$ (resp. $y$) has $c^{\prime}$ assigned as the color missing at $x$ (resp. $y$). Thus, we must sample one of these $2$ u-fans for the algorithm to flip the path $P$. ∎
Let $\mathcal{P}$ denote the set of all maximal $\{\alpha,\cdot\}$- and $\{\beta,\cdot\}$-alternating paths in $\chi$. We can see that the expected cost of the u-fan $\operatorname*{\boldsymbol{f}}$ sampled during the iteration is
$\mathbb{E}[\textnormal{{cost}}(\operatorname*{\boldsymbol{f}})]=\frac{1}{|\mathcal{U}|}\cdot\sum_{\operatorname*{\boldsymbol{f}}\in\operatorname*{\mathcal{U}}}\sum_{x\in\operatorname*{\boldsymbol{f}}}|P_{x}(\operatorname*{\boldsymbol{f}})|\leq\frac{2}{|\mathcal{U}|}\cdot\sum_{P\in\mathcal{P}}|P|\leq\frac{32m}{\lambda},$
where we are using the facts that (1) each path in $P\in\mathcal{P}$ appears at most twice while summing over the paths $|P_{x}(\operatorname*{\boldsymbol{f}})|$ by 6.6, (2) that $|\mathcal{U}|\geq\lambda/2$ by 6.2, and (3) that that the total length of all paths in $\mathcal{P}$ is $8m$ by 6.4. Applying Markov’s inequality, it follows that
$\Pr\!\left[\textnormal{{cost}}(\operatorname*{\boldsymbol{f}})\geq\frac{128m}{\lambda}\right]\leq\frac{1}{4}.$
Since each u-fan in $\Phi$ has 3 vertices and at most $2\Delta$$\{\alpha,\cdot\}$- or $\{\beta,\cdot\}$-alternating paths end at each of these vertices, we know that there are at most $6\Delta|\Phi|$ alternating paths that could be considered during some iteration that end at a u-fan in $\Phi$. By 6.6, there are at most $12\Delta|\Phi|$ u-fans in $\mathcal{U}$ that could cause the algorithm to consider one of these paths. Thus, at least $|\mathcal{U}|-12\Delta|\Phi|$ of the u-fans in $\mathcal{U}$ are contained in $\mathcal{U}^{\star}$. It follows that
$|\mathcal{U}^{\star}|\geq|\mathcal{U}|-12\Delta|\Phi|\geq|\mathcal{U}|-\frac{\lambda}{4}\geq\frac{|\mathcal{U}|}{2},$
where we are using the facts that $|\Phi|\leq\lambda/(48\Delta)$ and $|\mathcal{U}|\geq\lambda/2$ from 6.2. Since $\operatorname*{\boldsymbol{f}}$ is sampled uniformly at random from $\operatorname*{\mathcal{U}}$, it follows that $\Pr[\operatorname*{\boldsymbol{f}}\in\mathcal{U}^{\star}]\geq 1/2$. The lemma follows by applying a union bound. ∎
Lemma 6.7.
Each iteration of Prime-U-Fans takes time $O(m/\lambda)$.
Proof.
Using standard data structures, each iteration can be implemented in time proportional to the length of the alternating paths $P_{u}$, $P_{v}$ and $P_{w}$ considered by the algorithm during the iteration. We can check if $|P_{u}|+|P_{v}|+|P_{w}|\leq 128m/\lambda$ in $O(m/\lambda)$ time by traversing these paths and aborting if we notice that their total length exceeds $128m/\lambda$. If their total length is at most $128m/\lambda$, then we can flip these paths and update the collection $\operatorname*{\mathcal{U}}$ in $O(m/\lambda)$ time. ∎
Lemma 6.8.
The subroutine Prime-U-Fans runs in time $O(m\log n/\Delta)$ with high probability.
Proof.
Since each iteration of Prime-U-Fans succeeds with probability at least $1/4$ by Lemma 6.5 and the subroutine performs iterations until it succeeds $\lambda/(48\Delta)$ times, it follows that it performs at most $O(\lambda\log n/\Delta)$ iterations with high probability. Since each iteration takes $O(m/\lambda)$ time by Lemma 6.7, it follows that the total running time is $O(m\log n/\Delta)$ with high probability. ∎
6.2 The Subroutine Activate-U-Fans
As input, this subroutine is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, and a subset $\Phi\subseteq\operatorname*{\mathcal{U}}$ of $\mu$ vertex-disjoint $\{\alpha,\beta\}$-primed u-fans such that at most $O(m/\Delta)$ edges have color $\alpha$ or $\beta$. The subroutine repeatedly picks any $\operatorname*{\boldsymbol{f}}\in\Phi$ and proceeds to activate the u-fan. It repeats this process until $\Phi=\varnothing$. The pseudocode in Algorithm 7 gives a formal description of the subroutine.
1 while $\Phi\neq\varnothing$ do
2       Let $\operatorname*{\boldsymbol{f}}\in\Phi$
3       Let $P$ and $P^{\prime}$ be the $\{\alpha,\beta\}$-alternating paths starting at the leaves of $\operatorname*{\boldsymbol{f}}$
4       Activate the u-fan $\operatorname*{\boldsymbol{f}}$ by flipping the path $P$ or $P^{\prime}$
5       Remove $\operatorname*{\boldsymbol{f}}$ and any damaged u-fans from $\operatorname*{\mathcal{U}}$ and $\Phi$
6
[…]
Analysis of Activate-U-Fans
The following lemmas summarise the key properties of the subroutine Activate-U-Fans.
Lemma 6.9.
The subroutine Activate-U-Fans extends the coloring $\chi$ to at least $\mu/2$ more edges.
Proof.
During each iteration of the while loop, we activate a u-fan $\operatorname*{\boldsymbol{f}}$ and extend the coloring to an uncolored edge in $\operatorname*{\boldsymbol{f}}$. By Lemma 3.4, this process damages at most $2$ u-fans in $\operatorname*{\mathcal{U}}$ (including $\operatorname*{\boldsymbol{f}}$). Thus, in each iteration $|\Phi|$ decreases by at most $2$, and hence we extend the coloring to at least $\mu/2$ more edges across all the iterations performed by the subroutine. ∎
Lemma 6.10.
The subroutine Activate-U-Fans has a running time of $O(m/\Delta)$.
Proof.
Let $\mathcal{P}$ denote the collection of maximal $\{\alpha,\beta\}$-alternating paths in the coloring $\chi$ when we first call Activate-U-Fans. Since at most $O(m/\Delta)$ edges have color $\alpha$ and $\beta$ in this coloring, we have that the total length of the paths in $\mathcal{P}$ is $O(m/\Delta)$ since the paths in $\mathcal{P}$ are all vertex-disjoint.
Throughout the run of Activate-U-Fans, we only modify $\chi$ by flipping the colors of $\{\alpha,\beta\}$-alternating paths (see Algorithm 7). Thus, the structure of the paths in $\mathcal{P}$ does not change (but their colors might be flipped). Furthermore, each time the subroutine flips an $\{\alpha,\beta\}$-alternating path, it removes any u-fan from $\Phi$ that contains an endpoint of this alternating path (see Algorithm 7). Thus, each path in $\mathcal{P}$ is only flipped at most once.
Using standard data structures, each iteration of the while loop can be implemented in time proportional to the length of the alternating path that is flipped during the iteration (see Algorithm 7). It follows that the total running time of the subroutine across all iterations is at most $\sum_{P\in\mathcal{P}}O(|P|)\leq O(m/\Delta)$. ∎
Claim 6.11.
$|\mathcal{U}|$ decreases by at most $2\mu$ throughout the run of Activate-U-Fans.
Proof.
During each iteration, we flip an alternating path and remove the damaged u-fans from $\operatorname*{\mathcal{U}}$ and $\Phi$. It follows from Lemma 3.4 that each iteration removes at most $2$ u-fans from these sets. Since we perform at most $\mu$ iterations, the claim follows. ∎
6.3 Analysis of Color-U-Fans (Proof of Lemma 4.2)
Given a separable collection of $\lambda$ u-fans $\mathcal{U}$, the algorithm Color-U-Fans repeatedly calls Prime-U-Fans and Activate-U-Fans with the set $\mathcal{U}$ as described in Algorithm 5. It repeats this process for $\Delta/2$ iterations. It follows from Claims 6.2 and 6.11 that $|\mathcal{U}|$ decreases by at most a $(1-1/\Delta)$ factor during each iteration. Thus, by Bernoulli’s inequality, we get that
$|\mathcal{U}|\geq\lambda\cdot\left(1-\frac{1}{\Delta}\right)^{\Delta/2}\geq\frac{\lambda}{2}$
throughout the entire run of the algorithm. In each iteration of Algorithm 5, we extend the coloring to $\Omega(\lambda/\Delta)$ edges in $O(m\log n/\Delta)$ time w.h.p. Thus, in total, we extend the coloring to $\Omega(\lambda)$ edges in $O(m\log n)$ time w.h.p.
7 Implementation and Data Structures
In this section, we describe the key data structures that we use to implement an edge coloring $\chi$ and a separable collection $\operatorname*{\mathcal{U}}$, allowing us to efficiently implement the operations performed by our algorithms. We first describe the data structures and then show how they can be used to efficiently implement the queries described in Section 3.2.
Implementing an Edge Coloring: Let $G=(V,E)$ and a graph of maximum degree $\Delta$ and let $\mathcal{C}:=[\Delta+1]\cup\{\bot\}$. We implement an edge coloring $\chi:E\longrightarrow\mathcal{C}$ of $G$ using the following:
  • The map $\phi:E\longrightarrow\mathcal{C}$ where $\phi(e):=\chi(e)$ for all $e\in E$.
  • The map $\phi^{\prime}:V\times\mathcal{C}\longrightarrow E$ where $\phi_{u}^{\prime}(c):=\{e\ni u\mid\chi(e)=c\}$.
  • The set $\chi^{-1}(c):=\{e\in E\mid\chi(e)=c\}$, for all $c\in\mathcal{C}$.
  • The set $\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$, for all $u\in V$ [^11] We take this intersection with $[\deg_{G}(u)+1]$ instead of maintaining $\mathsf{miss}_{\chi}(u)$ directly to ensure that the space complexity and initialization time of the data structures are $\tilde{O}(m)$ and not $\Omega(\Delta n)$.
We implement all of the maps and sets using hashmaps, allowing us to perform membership queries, insertions and deletions in $O(1)$ time (see Proposition 3.6). [^12] By using balanced search trees instead of hashmaps, we can make the data structures deterministic while increasing the time taken to perform these operations to $O(\log n)$. The map $\phi^{\prime}$ allows us to check if a color $c\in[\Delta+1]$ is available at a vertex $u\in V$ in $O(1)$ time, and if it is not, to find the edge $e\ni u$ with $\chi(e)=c$. The sets $\{\chi^{-1}(c)\}_{c\in\mathcal{C}}$ allow us to easily return all edges with a specific color (including $\bot$). Furthermore, we can determine which color classes are the least common in $O(1)$ time. [^13] For example, we can then maintain a list of colors $c\in\mathcal{C}$ sorted by the values of $|\chi^{-1}(c)|$. Each time an edge $e$ changes color under $\chi$, we can easily update all of these data structures in $O(1)$ time. Furthermore, given $O(1)$ time query access to an edge coloring $\chi$, we can initialize these data structures in $O(m)$ time. We note that $\chi$ is a proper edge coloring if and only if $|\phi^{\prime}_{u}(c)|\leq 1$ for all $u\in V,c\in[\Delta+1]$.
Since the hashmap used to implement $\phi$ stores $m$ elements, it follows that it can be implemented with $O(m)$ space. Similarly, the map $\phi^{\prime}$ stores $2m$ elements (if $\{e\ni u\mid\chi(e)=c\}=\varnothing$, then we do not store anything for $\phi^{\prime}_{u}(c)$) and thus can be implemented with $O(m)$ space since each element has size $O(1)$ (recall that $|\phi^{\prime}_{u}(c)|\leq 1$ since the coloring is proper). since each set $\chi^{-1}(c)$ can be stored in space $O(|\chi^{-1}(c)|)$ and $\sum_{c}|\chi^{-1}(c)|=m$, these sets can be implemented in space $O(m)$. Similarly, since $\sum_{u}|\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]|=O(m)$, the sets $\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$ can also be implemented in $O(m)$ space.
Implementing a Separable Collection: We implement a separable collection $\operatorname*{\mathcal{U}}$ in a similar manner using the following:
  • The map $\psi:V\times[\Delta+1]\longrightarrow\operatorname*{\mathcal{U}}$ where $\psi_{u}(c):=\{\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}\mid u\in\operatorname*{\boldsymbol{g}},c_{\operatorname*{\boldsymbol{g}}}(u)=c\}$.
  • The set $C_{\operatorname*{\mathcal{U}}}(u):=\{c_{\operatorname*{\boldsymbol{g}}}(u)\mid\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}},u\in\operatorname*{\boldsymbol{g}}\}$, for all $u\in V$.
  • The set $\overline{C}_{\operatorname*{\mathcal{U}}}(u):=\left(\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]\right)\setminus C_{\operatorname*{\mathcal{U}}}(u)$, for all $u\in V$.
We again implement all of the maps and sets using hashmaps, allowing us to access and change entries in $O(1)$ time. We note that, since $\operatorname*{\mathcal{U}}$ is separable, $|\psi_{u}(c)|\leq 1$ for all $u\in V,c\in[\Delta+1]$. Thus, we can determine the size of $\operatorname*{\mathcal{U}}$ and also sample from $\operatorname*{\mathcal{U}}$ uniformly at random in $O(1)$ time. [^14] For example, we can sample a number $r\sim[|\mathcal{U}|]$ u.a.r. and then return the $r^{th}$ element in the hashmap that implements $\psi$. Each time we remove a color $c\in[\Delta+1]$ from the palette $\mathsf{miss}_{\chi}(u)$ of a vertex $u\in V$, we can update $\overline{C}_{\operatorname*{\mathcal{U}}}(u)$ in $O(1)$ time and check $\psi_{u}(c)$ in $O(1)$ time to find any u-component that has been damaged. Each time we add or remove a u-component from $\operatorname*{\mathcal{U}}$, we can update all of these data structures in $O(1)$ time. Furthermore, we can initialize these data structures for an empty collection in $O(m)$ time by creating an empty map $\psi$, empty sets $C_{\operatorname*{\mathcal{U}}}(u)$ for each $u\in V$ and copying the sets $\overline{C}_{\operatorname*{\mathcal{U}}}(u)=\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$ for each $u\in V$ which are maintained by the data structures for the edge coloring $\chi$. Since $\operatorname*{\mathcal{U}}$ is separable, we can see that $\overline{C}_{\operatorname*{\mathcal{U}}}(u)\neq\varnothing$. Thus, whenever we want a color from the set $\mathsf{miss}_{\chi}(u)\setminus C_{\operatorname*{\mathcal{U}}}(u)$, it suffices to take an arbitrary color from $\overline{C}_{\operatorname*{\mathcal{U}}}(u)$.
For each $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$, we can see that $\operatorname*{\boldsymbol{g}}$ is contained at most $3$ times in $\psi$. Thus, the total space required to store the hashmap that implements $\psi$ is $O(|\operatorname*{\mathcal{U}}|)$. Since $\operatorname*{\mathcal{U}}$ is separable, the u-components in $\operatorname*{\mathcal{U}}$ are edge-disjoint, and thus $|\mathcal{U}|\leq m$. It follows that map $\psi$ can be stored with $O(m)$ space. For each $u\in V$, we can observe that $|C_{\operatorname*{\mathcal{U}}}(u)|\leq\deg_{G}(u)$ since at most $\deg_{G}(u)$ many u-components in $\operatorname*{\mathcal{U}}$ contain the vertex $u$, and $|\overline{C}_{\operatorname*{\mathcal{U}}}(u)|\leq\deg_{G}(u)+1$ since $\overline{C}_{\operatorname*{\mathcal{U}}}(u)\subseteq[\deg_{G}(u)+1]$. Thus, the total space required to store the sets $\{C_{\operatorname*{\mathcal{U}}}(u)\}_{u\in V}$ and $\{\overline{C}_{\operatorname*{\mathcal{U}}}(u)\}_{u\in V}$ is $O(m)$.
7.1 Implementing the Operations from Section 3.2
We now describe how to implement each of the operations from Section 3.2.
Implementing $\textsc{Initialize}(G,\chi)$: Suppose that we are given the graph $G$ and $O(1)$ time query access to an edge coloring $\chi$ of $G$. We can initialize the data structures used to maintain the maps $\phi$ and $\phi^{\prime}$ in $O(m)$ time. We can then scan through the edges $e\in E$ and initialize the sets $\chi^{-1}(c)$ in $O(m)$ time. Finally, we can scan through the vertices $u\in V$ and initialize the sets $\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$ in $O(m)$ time. We can then initialize the data structures for an empty separable collection in $O(m)$ time by creating an empty map $\psi$ and, for each $u\in V$, initializing the sets $C_{\operatorname*{\mathcal{U}}}(u)\leftarrow\varnothing$ and $\overline{C}_{\operatorname*{\mathcal{U}}}(u)\leftarrow\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$.
Implementing $\textsc{Insert}_{\operatorname*{\mathcal{U}}}(\operatorname*{\boldsymbol{g}})$: By performing at most $3$ queries to the map $\psi$, we can check if $\operatorname*{\mathcal{U}}\cup\{\operatorname*{\boldsymbol{g}}\}$ is separable. If so we can update $\psi$ and the sets $C_{\operatorname*{\mathcal{U}}}(x)$ and $\overline{C}_{\operatorname*{\mathcal{U}}}(x)$ for $x\in\operatorname*{\boldsymbol{g}}$ in $O(1)$ time in order to insert $\operatorname*{\boldsymbol{g}}$ into $\operatorname*{\mathcal{U}}$. Otherwise, we return fail.
Implementing $\textsc{Delete}_{\operatorname*{\mathcal{U}}}(\operatorname*{\boldsymbol{g}})$: We can first make a query to $\psi$ to ensure that $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$. If so, we can update $\psi$ and the sets $C_{\operatorname*{\mathcal{U}}}(x)$ and $\overline{C}_{\operatorname*{\mathcal{U}}}(x)$ for $x\in\operatorname*{\boldsymbol{g}}$ in $O(1)$ time to remove $\operatorname*{\boldsymbol{g}}$ from $\operatorname*{\mathcal{U}}$.
[…]
Acknowledgements
Part of this work was conducted while the first and second named authors were visiting the Simons Institute for the Theory of Computing as part of the Sublinear Algorithms program.
Appendix A Logarithmic Factors and Small Values of $\Delta$
The main contribution and the primary message of our work is presenting the first near-linear time algorithm for $(\Delta+1)$-edge coloring and polynomial time factor improvements over prior work. As such, in the presentation of our algorithm and its analysis, we made no attempt to optimize the logarithmic factors, which led to an algorithm with $O(m\log^{3}{n})$ time for finding a $(\Delta+1)$-edge coloring with high probability.
Nevertheless, we can indeed optimize our algorithm to achieve a runtime of only $O(m\log{\Delta})$ time (with high probability), which matches the classical bound of [CH82, COS01] for $\Delta$-coloring bipartite graphs (although that algorithm is deterministic). Our algorithm is also related to a recent line of work in [BD23, BD24, Dha24] that focuses on the $\Delta=n^{o(1)}$ case; our result improves prior $O(m\Delta^{4}\log\Delta)$ bounds of [BD24] exponentially in terms of dependence on $\Delta$.
Theorem A.1.
There is a randomized algorithm that, given any simple undirected graph $G$ on $n$ vertices and $m$ edges with maximum degree $\Delta$, finds a $(\Delta+1)$-edge coloring of $G$ in $O(m\log{\Delta})$ time with high probability.
Let us first start by listing where the $\log(n)$-terms come from in the proof of Theorem 1.1:
  1. 1.
    In Lemma 4.1, we are losing an $O(\log{n})$ factor in the subroutine Reduce-U-Edges while turning u-edges into u-fans or coloring them. Specifically, in the proof of Lemma 5.13, we lose an additional $O(\log{n})$ factor when handling all u-edges (see the Harmonic series in the proof).
  2. 2.
    In Lemma 4.2, when coloring u-fans, we are losing an $O(\log{n})$ factor to ensure the probabilistic guarantees of the algorithm hold with high probability in each step (see Lemma 6.8).
  3. 3.
    Our application of Lemmas 4.1 and 4.2 can color a constant fraction of remaining uncolored edges, hence, we need to run them $O(\log{n})$ times to color all uncolored edges (see Lemma 4.3).
  4. 4.
    And finally, the entire framework of reducing $(\Delta+1)$ coloring to extending the coloring to $O(m/\Delta)$ uncolored edges using Eulerian partition technique (in the proof of Theorem 1.1) leads to a recursion depth of $O(\log{n})$ leading to another $O(\log{n})$ overhead in the runtime.
All in all, these factors lead to our $O(m\log^{3}{n})$ time algorithm (note that the losses in Steps 1 and 2 do not compound as we run these algorithms sequentially and thus pay an additive factor of $O(\log{n})$ between the two of them, not a multiplicative one).
We now show how these log factors can be removed or replaced by $O(\log{\Delta})$ in the following. Throughout, we continue our assumption in Section 3.2 that the high probability events of the hash maps used in the algorithm happen and we ignore their randomness (e.g., if an algorithm is deterministic modulo use of hash maps, we simply refer to it as deterministic here).
Step 1: The following lemma shows the $O(\log{n})$-term of Lemma 4.1 can be easily circumvented.
Lemma A.2 (A slight modification of Lemma 4.1).
There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$, and $\lambda$ uncolored edges, does one of the following in $O(m+\lambda\cdot\Delta)$ time deterministically:
  1. 1.
    Extends the coloring to $\Omega(\lambda)$ uncolored edges.
  2. 2.
    Modifies $\chi$ to obtain a separable collection of $\Omega(\lambda)$ u-fans.
Proof Sketch.
The $O(\log{n})$ term of the runtime in the proof of Lemma 4.1 comes from the Harmonic series calculation in the proof of Lemma 5.13. Specifically, for any color $\alpha$, assuming we start with $\lambda_{\alpha}$ $\alpha$-primed u-edges, by 5.12, the runtime needed for “handling” the $i^{th}$ such edge is $O((m_{\alpha}+\lambda_{\alpha})/i)$ time (by handling, we mean either coloring this edge or turning it into a u-fan). In the proof of Lemma 4.1, using 5.11, this results in a total runtime of
$O(\Delta\cdot\lambda_{\alpha})+O(1)\cdot\sum_{i=1}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}=O(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\log\lambda_{\alpha})=O(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\log{n}).$
However, a simple observation is that we do not need to handle all the $\lambda_{\alpha}$ $\alpha$-primed u-edges given that our final goal is to extend the coloring or create u-fans out of only a constant fraction of the initial $\lambda$ uncolored edges. Specifically, suppose our goal is to reduce the number of $\alpha$-primed u-edges to some $\gamma\cdot\lambda_{\alpha}$ for any arbitrary constant $\gamma\in(0,1)$; then, the runtime will be
$O(\Delta\cdot\lambda_{\alpha})+O(1)\cdot\sum_{i=\gamma\cdot\lambda_{\alpha}}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}=O\left(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\cdot\log{\left(\frac{\lambda_{\alpha}}{\gamma\cdot\lambda_{\alpha}}\right)}\right)=O(\Delta\cdot\lambda_{\alpha}+m_{\alpha}),$
given that $\gamma$ is a constant. As such, the time it takes to handle any arbitrary large constant fraction of u-edges will be only $O(m)$ (by summing over choices of $\alpha\in[\Delta+1]$). The rest of the proof follows exactly as in Lemma 4.1 by simply removing the small constant fraction of u-edges not handled from the graph for the rest of the proof. ∎
Step 2: We remove the extra $O(\log{n})$-term of Lemma 4.2 in a similar way although by focusing on the expected runtime of the algorithm (plus a crucial tail inequality). In the last step, we show how to recover the final result even from this weaker guarantee.
Lemma A.3 (A slight modification of Lemma 4.2).
There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ and a separable collection of $\lambda$ u-fans, extends $\chi$ to $\Omega(\lambda)$ in $T$ (randomized) time such that for some $T_{0}=O(m)$, we have,
$\mathbb{E}\left[T\right]\leq T_{0}\quad\text{and for all $\delta>0$}\quad\Pr\left[T\geq(1+\delta)\cdot 2T_{0}\right]\leq\exp\left(-\frac{\delta\cdot\lambda}{100}\right).$
Proof Sketch.
The runtime for coloring each single edge in Lemma 4.2 is $O(m/\lambda)$ by Lemma 6.7 assuming the coloring succeeds, which happens with probability at least $1/4$ by Lemma 6.5. We emphasize that this is in an average sense: the algorithm Color-U-Fans first uses Prime-U-Fans to prime $\Omega(\lambda/\Delta)$ u-fans in expected $O(m/\lambda)$ time per u-fan and thus $O((\lambda/\Delta)\cdot(m/\lambda))=O(m/\Delta)$ expected total time, and then deterministically colors them in $O(m/\Delta)$ time using Activate-U-Fans. It then repeats this process $\Delta/2$ times to color $\Omega(\lambda)$ edges, implying that $\mathbb{E}\left[T\right]=O(m)$ time.
We now prove the desired tail inequality on $T$ as well. Let us number all iterations of the while-loop in Prime-U-Fans from $1$ to $t$, across all $\Delta/2$ times Prime-U-Fans is called in Color-U-Fans: for the $i^{th}$ iteration, define an indicator random variable $X_{i}$ which is $1$ iff this iteration of the while-loop is successful in priming a u-fan (i.e., increasing the size of $\Phi$). Additionally, for every such $i$, define $S_{i}:=\sum_{j=1}^{i}X_{j}$ which denotes the number of successful iterations after the algorithm has done $i$ iterations in total. Thus, $t$ is the smallest integer where $S_{t}=\gamma\cdot\lambda$ for some integer $\gamma\in(0,1)$ which is the fraction of u-fans the algorithm colors (precisely, $\gamma=1/48\cdot 1/2\geq 1/100$).
For any $i\geq 1$, $S_{i}$ stochastically dominates the binomial distribution with parameters $i$ and $p=1/4$ (by Lemma 6.5 each iteration is successful with probability at least $1/4$). Hence, for $\delta>0$, by concentration results for the binomial distribution,
$\Pr\left[t\geq(1+\delta)\cdot 8\gamma\cdot\lambda\right]\leq\exp\left(-\frac{(2(1+\delta))^{2}}{2+2(1+\delta)}\cdot 4\gamma\cdot\lambda\right)\leq\exp\left(-\frac{\delta\cdot\lambda}{100}\right),$
using a loose upper bound in the last step. Given that the runtime of the algorithm is $t\cdot O(m/\lambda)$, the tail inequality follows. ∎
We note Lemma A.3 guarantees that as long as $\lambda=\omega(\log{n})$, the $O(m)$ runtime of the algorithm also holds with high probability. For smaller values of $\lambda$ (which happens only as a corner case in the algorithm [^15] The only case in our algorithm where this tail bound cannot be replaced with a high probability bound is when $m/\Delta=o(\log{n})$ and at the same time $\Delta\log^{2}\Delta=\omega(m\log{\Delta})$, which means $m=o(n\log{n})$ and yet $\Delta=\omega(n/\log{n})$.), we need the specific tail inequality proven in the lemma instead.
Step 3: In the next step, we provide a more fine-grained version of Lemma 4.3 that will be used in the last step of the argument to (somewhat) bypass the $O(\log{n})$ factor loss of the original lemma.
Lemma A.4 (A slight modification of Lemma 4.3).
There exists an absolute constant $b>0$ such that the following holds. There is an algorithm given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges, and an integer $\lambda_{0}\leq\lambda$, extends $\chi$ to all but $\lambda_{0}$ uncolored edges in $T$ (randomized) time such that for some $T_{0}=O(m\cdot\log{(\lambda/\lambda_{0})}+\Delta\lambda)$, we have,
$\mathbb{E}\left[T\right]\leq T_{0}\quad\text{and for all $\delta>0$}\quad\Pr\left[T\geq(1+\delta)\cdot 2T_{0}\right]\leq\exp\left(-\frac{\delta\cdot\lambda_{0}}{200}\right).$
Proof.
The algorithm is verbatim as in Lemma 4.3 by replacing Lemma 4.1 with Lemma A.2 and Lemma 4.2 with Lemma A.3: repeatedly color a constant fraction of remaining edges as long as $\lambda_{0}$ uncolored edges remain. Since we start with $\lambda$ uncolored edges and finish with $\lambda_{0}$ many, and reduce uncolored edges by a constant factor each time, we apply Lemmas A.2 and A.3 for a total of $O(\log{(\lambda/\lambda_{0})})$ times. Moreover, the $\Delta\lambda$ terms in the runtime of Lemma A.2 form a geometric series and thus their contribution to the runtime is $O(\Delta\lambda)$ in total. This proves the expected bound on the runtime, i.e., $\mathbb{E}\left[T\right]\leq T_{0}=O(m\cdot\log{(\lambda/\lambda_{0})}+\Delta\lambda)$.
As for the tail bound, for every application of Lemma A.3 with an intermediate value $\lambda^{\prime}\in[\lambda_{0},\lambda]$, and $T^{\prime}_{0}=O(m)$ being the $T_{0}$-parameter of Lemma A.3, we have,
[…]
Thus, by union bound,
Pr[total $\Pr\left[\text{total runtime}\geq k\cdot T_{0}\right]$ $\leq\sum_{\lambda^{\prime}}\exp\left(-\frac{\delta\cdot\lambda^{\prime}}{100}\right)$
$\leq\sum_{i=0}^{\infty}\cdot\exp\left(-\frac{\delta\cdot\lambda_{0}\cdot\gamma^{i}}{100}\right)$ (as number of uncolored edges drops by some constant factor, say, $\gamma=1+\Theta(1)$ each time)
$\leq\exp\left(-\frac{\delta\cdot\lambda_{0}}{200}\right),$
by (loosely) upper bounding the sum of the geometric series using its first term. ∎
Step 4: Finally, we show how to put these lemmas together to $(\Delta+1)$ edge color the graph in $O(m\log{\Delta})$ randomized time, and conclude the proof of Theorem A.1.
Before getting to the main argument, we need the following specialized concentration inequality that we prove here (although, its formulation appears general enough to be of use elsewhere also). The proof follows a standard moment generating function argument.
Proposition A.5.
Let $\{X_{i}\}_{i=1}^{n}$ be $n$ independent non-negative random variables associated with parameters $\{\alpha_{i}\}_{i=1}^{n}$ and $\{\beta_{i}\}_{i=1}^{n}$ such that for each $i\in[n]$, $\alpha_{i},\beta_{i}\geq 1$, and for every $\delta>0$,
$\Pr\left[X_{i}\geq(1+\delta)\cdot\alpha_{i}\right]\leq\exp\left(-\delta\cdot\beta_{i}\right);$
then, for every $t\geq 0$,
$\Pr\left[\sum_{i=1}^{n}X_{i}\geq\sum_{i=1}^{n}\alpha_{i}+t\right]\leq\exp\left(-\left(\min_{i=1}^{n}\frac{\beta_{i}}{2\alpha_{i}}\right)\cdot\left(t-2\sum_{i=1}^{n}\frac{\alpha_{i}}{\beta_{i}}\right)\right).$
Proof.
Let $s:=1/2\cdot\min_{i}\beta_{i}/\alpha_{i}$ and for $i\in[n]$, $Y_{i}:=\max(X_{i}-\alpha_{i},0)$. Letting $Y:=\sum_{i=1}^{n}Y_{i}$, leads
$\Pr\left[\sum_{i=1}^{n}X_{i}\geq\sum_{i=1}^{n}\alpha_{i}+t\right]\leq\Pr\left[Y\geq t\right],$
for all $t\geq 0$. We further have,
$\Pr\left[Y\geq t\right]=\Pr\left[e^{s\cdot Y}\geq e^{s\cdot t}\right]\leq e^{-s\cdot t}\cdot\mathbb{E}\left[e^{s\cdot Y}\right]=e^{-s\cdot t}\cdot\mathbb{E}\left[\prod_{i=1}^{n}e^{s\cdot Y_{i}}\right]=e^{-s\cdot t}\cdot\prod_{i=1}^{n}\mathbb{E}\left[e^{s\cdot Y_{i}}\right]$ (2)
where the inequality is by Markov inequality and the final equality is by the independence of the variables. Now, for every $i\in[n]$,
[…]
$=\frac{1}{\beta_{i}\cdot(1-s\cdot({\alpha_{i}}/{\beta_{i}}))}$ (since the integral of $e^{-Kx}=-e^{-Kx}/K$ and $s<\beta_{i}/\alpha_{i}$ thus the value of this function in $\infty$ is zero)
$\leq\frac{1}{1-s\cdot(\alpha_{i}/\beta_{i})}$ (as $\beta_{i}\geq 1$)
$\leq\exp\left(2s\cdot\alpha_{i}/\beta_{i}\right).$ (as $s\cdot\alpha_{i}/\beta_{i}<1/2$ and $1-x\geq e^{-2x}$ for $x<1/2$)
Plugging this bounds in Equation 2, implies that
$\Pr\left[Y\geq t\right]\leq\exp\left(-s\cdot t+\sum_{i=1}^{n}2s\cdot\frac{\alpha_{i}}{\beta_{i}}\right)=\exp\left(-\left(\min_{i=1}^{n}\frac{\beta_{i}}{2\alpha_{i}}\right)\cdot\left(t-2\sum_{i=1}^{n}\frac{\alpha_{i}}{\beta_{i}}\right)\right),$
by the choice of $s$, concluding the proof. ∎
We now get to the main parts of the argument. The first part is based on a modification to the Eulerian partition approach used in the proof of Theorem 1.1.
Lemma A.6.
There is an algorithm that given a graph $G$, finds a $(\Delta+1)$ edge coloring of all but $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability.
Proof.
Consider the following recursive algorithm. Find an Eulerian tour of $G$ and partition $G$ into two edge-disjoint subgraphs $G_{1}$ and $G_{2}$ on the same vertex set such that $\Delta(G_{1}),\Delta(G_{2})\leq\lceil\Delta/2\rceil$ and $m(G_{1}),m(G_{2})\leq\lceil m/2\rceil$. For $i\in\{1,2\}$, recursively, find a $(\Delta(G_{i})+1)$-edge coloring $\chi_{i}$ of $G_{i}$ that leaves $m(G_{i})/\Delta(G_{i})=O(m/\Delta)$ edges uncolored. Combining $\chi_{1}$ and $\chi_{2}$ and uncoloring the two smallest color classes, gives a $(\Delta+1)$ coloring $\chi$ of all but $\lambda:=O(m/\Delta)$ edges. We then run Lemma A.4 to reduce the number of uncolored edges to $\lambda_{0}:=m/\Delta$ edges.
The correctness of the algorithm follows immediately. For the runtime, we have $k:=O(\Delta)$ sub-problems in total, with $2^{d}$ subproblems on $m/2^{d}$ edges and maximum degree $\Delta/2^{d}$ for $d\leq\left\lceil\log{\Delta}\right\rceil$. Let $X_{1},\ldots,X_{k}$ denote the runtime of these subproblems and thus $X:=\sum_{i=1}^{k}X_{i}$ is the total runtime. We have,
$\mathbb{E}\left[X\right]=\sum_{i}\mathbb{E}\left[X_{i}\right]=\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot O\left(\frac{m}{2^{d}}\log{(\lambda/\lambda_{0})}+\frac{\Delta}{2^{d}}\cdot\lambda\right)=O(m\log{\Delta})$
by Lemma A.4, since $\lambda=O(m/\Delta)$, $\lambda_{0}=m/\Delta$, and thus $\log{(\lambda/\lambda_{0})}=O(1)$ and $\Delta\cdot\lambda=O(m)$. Moreover, by Lemma A.4, for the sub-problem corresponding to $X_{i}$ at some level $d$ of the recursion, $T^{d}_{0}=O(m/2^{d})$, and every $\delta\geq 0$,
$\Pr\left[X_{i}\geq(1+\delta)\cdot 2T^{d}_{0}\right]\leq\exp\left(-\delta\cdot\frac{\lambda_{0}}{200}\right)=\exp\left(-\delta\cdot\frac{m}{200\Delta}\right).$
For this $X_{i}$, define $\alpha_{i}:=2T^{d}_{0}$ and $\beta_{i}:=m/200\Delta$. We can now apply Proposition A.5 and for $t:=1000\sum_{i=1}^{k}\alpha_{i}=1000\cdot\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot 2T^{d}_{0}=O(m\log{\Delta})$, have,
$\Pr\left[X\geq t\right]$ $\leq\exp\left(-\left(\min_{i=1}^{n}\frac{\beta_{i}}{2\alpha_{i}}\right)\cdot\left(t-2\sum_{i=1}^{n}\frac{\alpha_{i}}{\beta_{i}}\right)\right)$
$=\exp\left(-\frac{\Theta(1)}{\Delta}\cdot\left(t-2\cdot\frac{t\cdot 200\Delta}{1000\cdot m}\right)\right)$
$\leq\exp\left(-\frac{\Theta(1)}{\Delta}\cdot(3t/5)\right)$ (as $\Delta\leq m$ always trivially)
$=\exp\left(-\Theta(1)\cdot\frac{m\log{\Delta}}{\Delta}\right),$
where we can additionally make the constant in $\Theta(1)$ arbitrarily large without changing the asymptotic runtime of the algorithm. This probability is always at most $1/\operatorname*{poly}(n)$ because either $\Delta<\sqrt{m}$, in which case, the probability is at most $1/2^{\Theta(\sqrt{m})}$, or $\Delta>\sqrt{m}$, and thus $\log{\Delta}=\Theta(\log{n})$ and the probability is $1/\operatorname*{poly}(n)$ at the very least. This concludes the proof. ∎
In the next part, we design an algorithm that color the remaining $m/\Delta$ uncolored edges left by Lemma A.6. This can be easily done in $O(m\log{n})$ time using Lemma A.4, but doing this in $O(m\log{\Delta})$ time instead requires one more ingredient.
For this step, we use the following result from [BD23] (itself based on [Ber22, Chr23]) that allows for coloring a random uncolored edge in essentially (but not exactly) $\operatorname*{poly}(\Delta)$ time. We note that a recent work of [BD24] presents a further improvement to these bounds (in terms of the exponent of $\Delta$) but as our dependence on $\Delta$ is only logarithmic, this improvement is inconsequential and thus we stick with the simpler work of [BD23].
Proposition A.7 ([BD23]).
There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$, and $\lambda$ uncolored edges, extends the coloring to a randomly chosen uncolored edge in expected $\operatorname*{poly}(\Delta)\cdot\log{(n/\lambda)}$ time.
We use this result to obtain the following lemma for our purpose.
Lemma A.8.
Let $c$ be the $\operatorname*{poly}(\Delta)$-exponent in Proposition A.7, and $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$. There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ and $\lambda=\Theta(\lambda_{0})$ uncolored edges, extends the coloring to all uncolored edges in $O(m\log\Delta)$ time with high probability.
Proof.
Suppose first that the max-term of $\lambda_{0}$ is the second one, i.e., $\lambda_{0}=m\log{n}/n$. This implies that $\Delta\geq(n/\log{n})^{1/2c}$ and since $c$ is a constant, we have $\log(\Delta)=\Theta(\log{n})$ in this case. We can use the standard Vizing’s fan and Vizing’s chain approach in Lemma 2.2 to color the remaining edges deterministically in $O(\lambda_{0}\cdot n)=O(m\log{n}/n\cdot n)=O(m\log{\Delta})$ time in this case.
Now consider the case when $\lambda_{0}=m/\Delta^{2c}$. We run the algorithm of Proposition A.7 with the following modification: when coloring a random edge, if the runtime becomes two times more than the expected time guarantee of Proposition A.7, we terminate the coloring and repeat from the beginning. Thus, each step takes $O(\Delta^{c}\cdot\log{(n/i)})$ time deterministically and succeeds in extending the coloring by one edge with probability at least half.
For $i\in[\lambda]$, define $\alpha_{i}:=4\log{(n/i)}$ and $X_{i}$ as the random variable denoting $\ell_{i}\cdot\alpha_{i}$ where $\ell_{i}$ is the number of times the algorithm attempts to color the next random edge until it succeeds when there are only $i$ edges left uncolored. We first prove that
$\sum_{i=1}^{\lambda}\alpha_{i}=o(m/\Delta^{c}),$ (3)
and then show that
$\Pr\!\left[\sum_{i=1}^{\lambda}X_{i}\geq\frac{2m}{\Delta^{c}}\right]\ll 1/\operatorname*{poly}(m),$ (4)
which concludes the proof in this case as well since the runtime of coloring the remaining edges is $O(\Delta^{c_{0}}\cdot\sum_{i=1}^{\lambda}X_{i})$ as stated earlier.
Proof of Equation 3. We have,
$\sum_{i=1}^{\lambda}\alpha_{i}$ $=\sum_{i=1}^{\lambda}4\log{\!\left(\frac{n}{i}\right)}=4\log\!\left(\prod_{i=1}^{\lambda}\frac{n}{i}\right)=4\log\!\left(\frac{n^{\lambda}}{\lambda!}\right)\leq\log\!\left(\left(\frac{e\cdot n}{\lambda}\right)^{\lambda}\right)$ (using the inequality $x!\geq(x/e)^{x}$ for all $x\geq 1$)
$=4\cdot\frac{m}{\Delta^{2c}}\cdot\log\left(\frac{e\cdot n\cdot\Delta^{2c}}{m}\right)=o(m/\Delta^{c}).$ (as $\lambda=m/\Delta^{2c}$ and $m\geq n$)
Proof of Equation 4. Note that for any $i\in[\lambda]$ and $\delta>0$,
$\Pr\left[X_{i}\geq(1+\delta)\cdot\alpha_{i}\right]=\Pr\left[\ell_{i}\geq(1+\delta)4\right]\leq 2^{-(3+4\delta)}\leq\exp\left(-\delta\right),$
where the first inequality is because the probability that $\ell_{i}$ increases by one each time is at most $1/2$ by Markov’s inequality.
While $X_{i}$’s are technically not independent (the choice of some edge being colored may change the structure of the graph for coloring next uncolored edge), they are stochastically dominated by independent random variables with above tail bounds (given the guarantee of Proposition A.7 for each uncolored edge). Thus, we can apply Proposition A.5 with parameters $\{\alpha_{i}\}_{i=1}^{\lambda}$ as above and $\beta_{i}=1$ for all $i\in[\lambda]$ to obtain that for $t:=m/\Delta^{c}$,
$\Pr\!\left[\sum_{i=1}^{\lambda}X_{i}\geq\frac{2m}{\Delta^{c}}\right]\leq\Pr\left[\sum_{i=1}^{\lambda}X_{i}\geq\sum_{i=1}^{\lambda}\alpha_{i}+t\right]$ $\leq\exp\left(-\frac{1}{4\log{n}}\cdot\frac{m}{2\Delta^{c}}\right)\ll 1/\operatorname*{poly}(m),$
where the first two inequalities use Equation (3) for $\sum_{i}\alpha_{i}$ and the third uses $\Delta^{c}\leq(n/\log{n})^{c/2c}<\sqrt{n}$ in this case (when $\lambda_{0}$ is the first-term of its maximum). This concludes the proof. ∎
Finally, we present an algorithm for taking care of the gap between the $m/\Delta$ uncolored edges left by Lemma A.6 and the $\Theta(1)\cdot\max\{m/\Delta^{2c},m\log{n}/n\}$ edges that Lemma A.8 can handle.
Lemma A.9.
Let $c=\Theta(1)$ and $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$ be as in Lemma A.8. There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ with $\lambda=m/\Delta$ uncolored edges, extends the coloring to all but $\Theta(\lambda_{0})$ uncolored edges in $O(m\log{\Delta})$ time with high probability.
Proof.
We simply run our Lemma A.4 with the given parameters $\lambda$ and $\lambda_{0}$. By Lemma A.4, the expected runtime will be some $T_{0}=O(m\log{(\lambda/\lambda_{0})}+\Delta\lambda)=O(m\log{\Delta})$ since $\lambda_{0}\geq m/\Delta^{2c}$ for some absolute constant $c\geq 1$. Moreover, by the same lemma,
Pr[runtime of the $\Pr\left[\text{runtime of the algorithm}\geq 2000\cdot T_{0}\right]\leq\exp\left(-10\lambda_{0}\right)\leq\exp\left(-\Theta(m\log{n})/n\right)=1/\operatorname*{poly}(n),$
since $\lambda_{0}\geq m\log{n}/n$. Thus, the algorithm finishes in $O(m\log{\Delta})$ time with high probability. ∎
Putting our above lemmas together concludes the proof of Theorem A.1 as follows:
Proof of Theorem A.1.
The proof follows immediately from the lemmas in Step 4:
  • By Lemma A.6, we can color all but $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability;
  • Then, by Lemma A.9, we can extend the coloring to all but $\Theta(\lambda_{0})$ edges of the graph in $O(m\log{\Delta})$ time with high probability, where $\lambda_{0}$ is as in Lemma A.8;
  • Finally, by Lemma A.8, we can color all the remaining $\Theta(\lambda_{0})$ uncolored edges in $O(m\log{\Delta})$ time with high probability.
This concludes the proof. ∎

### v3

@@
 [^1] Throughout, we use $\tilde{O}(f):=O(f\operatorname*{polylog}{(n)})$ to suppress log-factors in the number of vertices of the graph.
-Very recently, independently and concurrently, using randomization, this runtime bound was further improved to $\tilde{O}(n^{2})$ by [Assadi, 2024] and $\tilde{O}(mn^{1/3})$ by [Bhattacharya, Carmon, Costa, Solomon and Zhang, 2024] (and subsequently to $\tilde{O}(mn^{1/4})$ time by [Bhattacharya, Costa, Solomon and Zhang, 2024]).
-We present an algorithm that computes a $(\Delta+1)$-edge coloring in $\tilde{O}(m)$ time—in fact, even $O(m\log{\Delta})$ time—with high probability, giving a near-optimal algorithm for this fundamental problem.
+Very recently, independently and concurrently, using randomization, this runtime bound was further improved to $\tilde{O}(n^{2})$ by [Assadi, 2024] and $\tilde{O}(mn^{1/3})$ by [Bhattacharya, Carmon, Costa, Solomon and Zhang, 2024] (and subsequently to $\tilde{O}(mn^{1/4})$ by [Bhattacharya, Costa, Solomon and Zhang, 2024]).
+In this paper, we present a randomized algorithm that computes a $(\Delta+1)$-edge coloring in near-linear time—in fact, only $O(m\log{\Delta})$ time—with high probability, giving a near-optimal algorithm for this fundamental problem.
+Contents
 1 Introduction
@@
 Moreover, it was proven by [Hol81] that it is NP-complete to distinguish whether the edge chromatic number of a given graph is $\Delta$ or $\Delta+1$, and therefore $\Delta+1$ is the best bound we can hope for with polynomial time algorithms.
-Vizing’s original proof easily extends to an $O(mn)$ time algorithm, which was improved to $\tilde{O}(m\sqrt{n})$ in the 1980s by Arjomandi [Arj82] and Gabow et al.
-[GNK+85] independently.
-More recently, Sinnamon [Sin19] simplified the algorithms of [Arj82, GNK+85] by using randomization, while shaving off extra logarithmic factors from their runtime complexities, achieving a clean $O(m\sqrt{n})$ runtime bound.
-Very recently, this longstanding $O(m\sqrt{n})$ time barrier was bypassed in two recent concurrent works by Assadi [Ass24] and Bhattacharya et al.
-[BCC+24] which improve the runtime bound to two incomparable bounds of $\tilde{O}(n^{2})$ and $\tilde{O}(mn^{1/3})$, respectively.
-In a follow-up work, Bhattacharya et al.
-[BCSZ25] improved the $\tilde{O}(mn^{1/3})$ time bound of [BCC+24] to $\tilde{O}(mn^{1/4})$.
-In this work, we resolve the time complexity of randomized $(\Delta+1)$-edge coloring (up to small logarithmic factors) by presenting a near-linear time algorithm for this problem.
+Vizing’s original proof easily extends to an $O(mn)$ time algorithm, which was improved to $\tilde{O}(m\sqrt{n})$ in the 1980s by [Arj82] and [GNK+85] independently.
+More recently, the algorithms of [Arj82, GNK+85] were simplified in [Sin19], while shaving off extra logarithmic factors from their runtime complexities, achieving a clean $O(m\sqrt{n})$ runtime bound.
+Very recently, this longstanding $O(m\sqrt{n})$ time barrier was bypassed in two concurrent works [Ass25] and [BCC+24] which improved the runtime bound to two incomparable bounds of $\tilde{O}(n^{2})$ and $\tilde{O}(mn^{1/3})$, respectively.
+In a follow-up work, the $\tilde{O}(mn^{1/3})$ runtime bound of [BCC+24] to was further improved to $\tilde{O}(mn^{1/4})$ in [BCSZ25].
+In this work, we resolve the time complexity of randomized $(\Delta+1)$-edge coloring up to at most a log factor by presenting a near-linear time algorithm for this problem.
 Theorem 1.1.
-There is a randomized algorithm that, given any simple undirected graph $G=(V,E)$ on $n$ vertices and $m$ edges with maximum degree $\Delta$, finds a $(\Delta+1)$-edge coloring of $G$ in $\tilde{O}(m)$ time with high probability.
-Remarks:
-  1. 1.
-    Our algorithm can be implemented to run in $O(m\log{\Delta})$ time with high probability.
-Nonetheless, our main focus is on polynomial runtime improvements in general, and achieving a near-linear time algorithm in particular.
-For this reason as well as for the sake of simplicity and transparency of our techniques, the main body of the paper presents and analyzes an algorithm that runs in $O(m\log^{3}n)$ time with high probability, which showcases our most novel ideas.
-A more careful implementation of the same exact algorithm, with proper adaptations, achieves $O(m\log\Delta)$ time; we defer the details of this part to Appendix A.
-  2. 2.
-    The $O(m\log\Delta)$ time bound that our final algorithm achieves matches the longstanding time bound for $\Delta$-edge coloring bipartite graphs [COS01] (which, to the best of our knowledge, is also the best known runtime for $(\Delta+1)$-edge coloring bipartite graphs).
-Our algorithm is also related to a recent line of work in [BD23, BD24, Dha24] that focuses on the $\Delta=n^{o(1)}$ case, which in particular gives a randomized $(\Delta+1)$-coloring that runs in $O(m\Delta^{4}\log\Delta)$ time with high probability.
+There is a randomized algorithm that, given any simple undirected graph $G=(V,E)$ on $n$ vertices and $m$ edges with maximum degree $\Delta$, finds a $(\Delta+1)$-edge coloring of $G$ in $O(m\log{n})$ time with high probability.
+Remarks: Several remarks on our Theorem 1.1 are in order:
+  • Our algorithm in Theorem 1.1 does not rely on any of the recent developments in [Ass25, BCC+24, BCSZ25] and takes an entirely different path.
+We present an overview of our approach, as well as a comparison to these recent developments in Section 1.2.
+  • Our main contribution in this work is to improve the time-complexity of $(\Delta+1)$-edge coloring by polynomial factors all the way to near-linear.
+For this reason, as well as for the sake of transparency of our techniques, we focus primarily on presenting an $O(m\log^{3}{n})$ time randomized algorithm (Theorem 6.1), which showcases our most novel ideas.
+Later in Section 10, we show that a more careful implementation of the same algorithm achieves a clean $O(m\log n)$ runtime.
+  • We can additionally use a result of [BD23] to further replace the $\log{n}$ term in Theorem 1.1 with a $\log{\Delta}$ term, leading to an algorithm for $(\Delta+1)$-edge coloring in $O(m\log{\Delta})$ time with high probability (Corollary C.1).
+This matches the longstanding time bound for $\Delta$-edge coloring bipartite graphs [COS01] (which, to the best of our knowledge, is also the best known runtime for $(\Delta+1)$-edge coloring bipartite graphs).
+This bound is also related to a recent line of work in [BD23, BD24, Dha24b] that focused on the $\Delta=n^{o(1)}$ case and gave a randomized $(\Delta+1)$-coloring algorithm that runs in $O(m\Delta^{4}\log\Delta)$ time with high probability [BD24].
+  • Vizing’s theorem generalizes for (loop-less) multigraphs, asserting that any multigraph with edge multiplicity at most $\mu$ can be $(\Delta+\mu)$-edge colored [Viz64, Viz65].
+A related result is Shannon’s theorem [Sha49] that asserts that any multigraph can be $\left\lfloor 3\Delta/2\right\rfloor$ edge colored independent of $\mu$; both these bounds are tight: see the so-called Shannon multigraphs [Viz65].
+We show that our techniques extend to these theorems as well, giving $O(m\log{\Delta})$ time algorithms for both problems (see Theorems A.1 and B.1 and Corollary C.1).
 1.1 Related Work
-There has been a long line of work on fast algorithms for edge coloring that use more than $\Delta+1$ colors.
+In addition to algorithms for Vizing’s theorem, there has also been a long line of work on fast algorithms for edge coloring that use more than $\Delta+1$ colors.
 It was first shown in [KS87] that an edge coloring can be computed in $\tilde{O}(m)$ time when we have $\Delta+\tilde{O}(\sqrt{\Delta})$ different colors.
-In addition, there are algorithms which run in linear or near-linear time for $(1+\epsilon)\Delta$-edge coloring [DHZ19, BCPS24c, EK24, BD24, Dha24] when $\epsilon\in(0,1)$ is a constant.
-Most recently, it was shown in [Ass24] that even a $(\Delta+O(\log n))$-edge coloring can be computed in near-linear time.
+In addition, there are algorithms which run in linear or near-linear time for $(1+\epsilon)\Delta$-edge coloring [DHZ19, BCPS24c, EK24, BD24, Dha24b] when $\epsilon\in(0,1)$ is a constant.
+Most recently, it was shown in [Ass25] that even a $(\Delta+O(\log n))$-edge coloring can be computed in $O(m\log{\Delta})$ expected time.
 There are other studies on restricted graph classes.
-In bipartite graphs, a $\Delta$-edge coloring can be computed in $\tilde{O}(m)$ time [CH82, COS01].
+In bipartite graphs, a $\Delta$-edge coloring can be computed in $\tilde{O}(m)$ time [CH82, COS01, Alo03, GKK10].
 In bounded degree graphs, one can compute a $(\Delta+1)$-edge coloring in $\tilde{O}(m\Delta)$ time [GNK+85], and it was generalized recently for bounded arboricity graphs [BCPS24b]; see also [BCPS24a, CRV24, Kow24] for further recent results on edge coloring in bounded arboricity graphs.
 Subfamilies of bounded arboricity graphs, including planar graphs, bounded tree-width graphs and bounded genus graphs, were studied in [CY89, CN90, CK08].
-Beside the literature in the classical sequential model, considerable effort has been devoted to the study of edge coloring in various computational models in the past few years, including dynamic [BM17, BCHN18, DHZ19, Chr23, BCPS24c, Chr24], online [CPW19, BGW21, SW21, KLS+22, BSVW24, DGS24], distributed [PR01, EPS14, FGK17, GKMU18, BBKO22, CHL+20, Ber22, Chr23, Dav23], and streaming [BDH+19, SB24, CMZ24, GS24] models, among others.
+Beside the literature on classical algorithms, considerable effort has been devoted to the study of edge coloring in various computational models in the past few years, including dynamic [BM17, BCHN18, DHZ19, Chr23, BCPS24c, Chr24], online [CPW19, BGW21, SW21, KLS+22, BSVW24, BSVW25, DGS25], distributed [PR01, EPS14, FGK17, GKMU18, GP20, BBKO22, CHL+20, Ber22, Chr23, Dav23], and streaming [BDH+19, SB24, CMZ24, GS24] models, among others.
 1.2 Technical Overview
 We now present an overview of prior approaches to $(\Delta+1)$-edge coloring and describe our techniques at a high level.
-For the following discussions, we will assume basic familiarity with Vizing’s proof and its underlying algorithm; see Section 2.2 for more details on Vizing’s algorithm.
-Prior Approaches: A generic approach for $(\Delta+1)$-edge coloring, dating back to Vizing’s proof itself, is to extend a partial $(\Delta+1)$-edge coloring of the graph one edge at a time, possibly by recoloring some subset of edges, until the entire graph becomes colored.
+For the following discussions, we will assume basic familiarity with Vizing’s proof and its underlying algorithm; see Section 4 for more details on Vizing’s algorithm.
+Prior Approaches: A generic approach for $(\Delta+1)$-edge coloring, dating back to Vizing’s proof itself, is to extend a partial $(\Delta+1)$-edge coloring of the graph one edge at a time, possibly by recoloring some edges, until the entire graph becomes colored.
 As expected, the main bottleneck in the runtime of this approach comes from extending the coloring to the last few uncolored edges.
-For instance, given a graph $G=(V,E)$ with maximum degree $\Delta$, we can apply Eulerian partition to divide $G$ into two edge-disjoint subgraphs $G=G_{1}\cup G_{2}$ with maximum degrees at most $\left\lceil\Delta/2\right\rceil$.
+For instance, given a graph $G=(V,E)$ with maximum degree $\Delta$, we can apply Eulerian partitions to divide $G$ into two edge-disjoint subgraphs $G=G_{1}\cup G_{2}$ with maximum degrees at most $\left\lceil\Delta/2\right\rceil$.
 We then find $(\left\lceil\Delta/2\right\rceil+1)$-edge colorings of the subgraphs $G_{1}$ and $G_{2}$ recursively.
@@
 Since all the Vizing chains of the same type are vertex-disjoint, they can be flipped simultaneously and their total length is only $O(n)$.
-This means that the runtime of coloring all edges of a single type can be bounded by $O(n)$ time as well.
+This means that the runtime of coloring all edges of a single type can be bounded by $O(n)$ as well.
 This leads to an $O(n\Delta^{2})$ time algorithm for handling all $O(\Delta^{2})$ types; a more careful analysis can bound this even with $O(m\Delta)$ time.
@@
 There has been some very recent progress that broke through this classical barrier of $O(m\sqrt{n})$ time in [GNK+85, Arj82].
-In [BCC+24], the authors proved a new useful condition for extending the coloring to uncolored edges: if the uncolored edges admit a small vertex cover, then there is also a faster algorithm for extending the coloring to them.
-They then show how to precondition the problem so that uncolored edges do indeed admit a small vertex cover, leading to an algorithm with $\tilde{O}(mn^{1/3})$ time.
-In [Ass24], the author avoided the need for Eulerian partition and recursion altogether by instead designing a new near-linear time algorithm for $(\Delta+O(\log n))$-edge coloring.
+In [BCC+24], the authors speed up the extension of the coloring to uncolored edges when these edges admit a small vertex cover.
+They then show how to precondition the problem so that uncolored edges admit a small vertex cover, leading to a $\tilde{O}(mn^{1/3})$ time algorithm.
+In [Ass25], the author avoided the need for Eulerian partition and recursion altogether by instead designing a new near-linear time algorithm for $(\Delta+O(\log n))$-edge coloring.
 This algorithm borrows insights from sublinear matching algorithms in regular bipartite graphs by [GKK10] and is thus completely different from other edge coloring algorithms mentioned above.
@@
 Applying Vizing’s procedure for these uncolored edges takes additional $\tilde{O}(mn/\Delta)=\tilde{O}(n^{2})$ time, leading to an $\tilde{O}(n^{2})$ time algorithm for $(\Delta+1)$-edge coloring.
-Finally, in [BCSZ25], the authors showed that a $(\Delta+1)$-coloring can be computed in $\tilde{O}(mn^{1/4})$ time, by using the algorithm of [Ass24] for initial coloring of the graph and then presenting an improved color extension subroutine with a runtime of $\tilde{O}(\Delta^{2}+\sqrt{\Delta n})$ for coloring each remaining edge; the best previous color extension time bounds were either the trivial $O(n)$ bound or the bound $\tilde{O}(\Delta^{6})$ by [Ber22].
+Finally, in [BCSZ25], the authors showed that a $(\Delta+1)$-coloring can be computed in $\tilde{O}(mn^{1/4})$ time, by using the algorithm of [Ass25] for initial coloring of the graph and then presenting an improved color extension subroutine with a runtime of $\tilde{O}(\Delta^{2}+\sqrt{\Delta n})$ for coloring each remaining edge; the best previous color extension time bounds were either the trivial $O(n)$ bound or the bound $\tilde{O}(\Delta^{4})$ by [Ber22, BD24].
 Our Approach: A Near-Linear Time Color Extension Algorithm
-Our algorithm takes an entirely different path compared to previous work on this problem; specifically, we no longer attempt to design a faster color extension for a single edge, and instead color them in large batches, which allows for a much better amortization of runtime in coloring multiple edges.
-This ultimately leads to our main contribution — a new randomized algorithm for solving the aforementioned color extension problem for the last $O(m/\Delta)$ edges in $\tilde{O}(m)$ time.
-With this algorithm at hand, we can simply follow the Eulerian partition approach mentioned earlier and obtain an algorithm for $(\Delta+1)$-edge coloring whose runtime $T(m)$ follows the recursion $T(m)\leq 2T(m/2)+\tilde{O}(m)$; this implies that $T(m)=\tilde{O}(m)$, hence, giving us a near-linear time algorithm for $(\Delta+1)$-edge coloring.
-This way, we will not even need to rely on the $(\Delta+O(\log{n}))$-edge coloring algorithm of [Ass24] to color the earlier parts of the graph (although one can use that algorithm instead of Eulerian partition approach to the same effect).
+We will no longer attempt to design a faster color extension for a single edge, and instead color them in large batches like in [GNK+85], which allows for a much better amortization of runtime in coloring multiple edges.
+This ultimately leads to our main technical contribution: a new randomized algorithm for solving the aforementioned color extension problem for the last $O(m/\Delta)$ edges in $\tilde{O}(m)$ time.
+With this algorithm at hand, we can follow the aforementioned Eulerian partition approach and obtain a $(\Delta+1)$-edge coloring algorithm whose runtime $T(m)$ follows the recursion $T(m)\leq 2T(m/2)+\tilde{O}(m)$ with high probability; this implies that $T(m)=\tilde{O}(m)$, hence, giving us a near-linear time randomized algorithm for $(\Delta+1)$-edge coloring.
+This way, we will not even need to rely on the $(\Delta+O(\log{n}))$-edge coloring algorithm of [Ass25] to color the earlier parts of the graph (although one can use that algorithm instead of Eulerian partition approach to the same effect).
 We now discuss the main ideas behind our color extension algorithm.
@@
 Color Type Reduction: Recall that the runtime of $O(n\Delta^{2})$ for the color extension algorithm of [GNK+85] is due to the fact that there are generally $O(\Delta^{2})$ types of alternating paths in the graph and that the total length of the paths of each color type is bounded by $O(n)$ edges.
-However, if it so happens that the existing partial coloring of the graph only involves $O(\Delta)$ color types instead, then the same algorithm will only take $O(n\Delta)=O(m)$ time (by the near-regularity assumption) and we will be done.
-As such, the underlying idea behind our algorithm is to modify the current partial edge coloring (without decreasing the number of uncolored edges) so that the total number of color types reduces from $O(\Delta^{2})$ to $O(\Delta)$ only.
-To explore this direction, let us assume for now the input graph is bipartite, which greatly simplifies the structure of Vizing fans and Vizing chains, thus allowing us to convey the key ideas more clearly; later we highlight some of the key challenges that arise when dealing with general graphs.
+However, if it so happens that the existing partial coloring only involves $O(\Delta)$ color types instead, then the same algorithm will only take $O(n\Delta)=O(m)$ time (by the near-regularity assumption).
+The underlying idea behind our algorithm is to modify the current partial edge coloring (without decreasing the number of uncolored edges) so that the number of color types reduces from $O(\Delta^{2})$ to $O(\Delta)$ only.
+To explore this direction, let us assume for now the input graph is bipartite, which greatly simplifies the structure of Vizing fans and Vizing chains, thus allowing us to convey the key ideas more clearly (see Section 3 for a more detailed exposition); later we highlight some of the key challenges that arise when dealing with general graphs.
 We shall note that it has been known since the 80s that one can $\Delta$-edge color bipartite graphs in $\tilde{O}(m)$ time [CH82, COS01].
@@
 More specifically, say that we have popularized a set $\Phi$ of uncolored edges.
-However, when popularizing the next uncolored edge $(u,v)$, it could be the case that the $\{\alpha,c_{u}\}$-alternating path $P_{u}$ is ending at a vertex $u^{\prime}$ for some edge $(u^{\prime},v^{\prime})\in\Phi$.
+When popularizing the next uncolored edge $(u,v)$, it could be the case that the $\{\alpha,c_{u}\}$-alternating path $P_{u}$ is ending at a vertex $u^{\prime}$ for some edge $(u^{\prime},v^{\prime})\in\Phi$.
 If we flip the path $P_{u}$, then $(u^{\prime},v^{\prime})$ would no longer be of $\{\alpha,\beta\}$-type as $\alpha$ would not be missing at $u^{\prime}$ anymore.
+See Figure 1 for an illustration.
+[…]
+Figure 1: In this picture, we attempt to popularize edge $(u,v)$ by flipping the $\{\alpha,\gamma\}$-alternating path from $u$ and the $\{\beta,\lambda\}$-alternating path from $v$.
+However, flipping the $\{\alpha,\gamma\}$-alternating path from $u$ makes a previously popular edge $(u^{\prime},v^{\prime})$ unpopular as $u^{\prime}$ will not miss color $\alpha$ anymore.
 Our key observation is that when $|\Phi|$ is relatively small, most choices for an alternating path $P_{u}$ cannot be ending at edges in $\Phi$.
@@
 As such, ranging over all possible choices of $\gamma\in[\Delta+1]$, there are at most $O(|\Phi|\Delta)$ uncolored edges $(u,v)$ whose alternating paths could damage the set $\Phi$.
-Therefore, as long as the size of $\Phi$ is a $o(1/\Delta)$ fraction (or more precisely, a $O(1/\Delta)$ fraction, for a sufficiently small constant hiding in the $O$-notation) of all uncolored edges, the following property holds: a constant fraction of uncolored edges $e\notin\Phi$ can be popularized using the above method without damaging $\Phi$.
+Therefore, as long as the size of $\Phi$ is a $o(1/\Delta)$ fraction (or more precisely, an $O(1/\Delta)$ fraction, for a sufficiently small constant hiding in the $O$-notation) of all uncolored edges, the following property holds: a constant fraction of uncolored edges $e\notin\Phi$ can be popularized using the above method without damaging $\Phi$.
+See Figure 2 for an illustration.
 This property resolves both technical issues raised earlier simultaneously.
@@
 As a direct corollary, we can color all uncolored edges in $O(m\log{n})$ time, hence solving the color extension problem, in (near-regular) bipartite graphs, in near-linear time.
-The above argument will be detailed in the proof of Lemma 4.2 in Section 6.
+The above argument will be detailed in the proof of Lemma 3.2 in Section 3 (see also Lemma 6.3 of Section 8 for the analogous but not identical argument in general graphs).
+$u^{\prime}$$v^{\prime}$$v_{1}$$u_{1}$$v_{2}$$u_{2}$$v_{3}$$u_{3}$$v_{4}$$u_{4}$$v_{5}$$u_{5}$$\beta$$\beta$$\beta$$\beta$$\beta$$\beta$$\beta$$\beta$$\beta$$\beta$
+Figure 2: In this picture, $(u^{\prime},v^{\prime})\in\Phi$ with $c_{u^{\prime}}=\alpha,c_{v^{\prime}}=\beta$.
+For each uncolored edge $(u_{i},v_{i})$, flipping the $\{c_{i},\beta\}$-alternating path from $v_{i}$ would damage the property that $c_{v^{\prime}}=\beta$.
+Fortunately, there are at most $\Delta$ many different such $(u_{i},v_{i})$ because each of them is at the end of an $\{\beta,\cdot\}$-alternating path starting at $v^{\prime}$.
 Collecting U-Fans in General Graphs: We now discuss the generalization of our scheme above to non-bipartite graphs.
 Existence of odd cycles in non-bipartite graphs implies that we can no longer assign a color type $\{c_{u},c_{v}\}$ to an uncolored edge $(u,v)$ for $c_{u}\in\mathsf{miss}_{\chi}(u),c_{v}\in\mathsf{miss}_{\chi}(v)$, and hope that flipping the $\{c_{u},c_{v}\}$-alternating path from $u$ allows us to color the edge $(u,v)$ with $c_{v}$ (because the path may end at $v$, and thus after flipping it $c_{v}$ will no longer be missing from $v$).
-This is where Vizing fans and Vizing chains come into play: in non-bipartite graphs, a color type of an uncolored edge $(u,v)$ is $\{c_{u},\gamma_{u,v}\}$ where $c_{u}$ is an arbitrary color in $\in\mathsf{miss}_{\chi}(u)$ but $\gamma_{u,v}$ is determined by the Vizing fan around $u$ and the alternating path that we take for coloring $(u,v)$ (or the Vizing chain of $(u,v)$).
-Thus, while as before we can switch $c_{u}$ with some fixed color $\alpha$, it is unclear how to flip alternating paths to change $\gamma_{u,v}$ to $\beta$ also, in order to popularize the edge $(u,v)$ to be of some designated type $\{\alpha,\beta\}$ —without damaging another popularized edge as a result.
+This is where Vizing fans and Vizing chains come into play: in non-bipartite graphs, a color type of an uncolored edge $(u,v)$ is $\{c_{u},\gamma_{u,v}\}$ where $c_{u}$ is an arbitrary color in $\mathsf{miss}_{\chi}(u)$ but $\gamma_{u,v}$ is determined by the Vizing fan around $u$ and the alternating path that we take for coloring $(u,v)$ (or the Vizing chain of $(u,v)$).
+Thus, while as before we can switch $c_{u}$ with some fixed color $\alpha$, it is unclear how to flip alternating paths to change $\gamma_{u,v}$ to $\beta$ also, in order to popularize the edge $(u,v)$ to be of some designated type $\{\alpha,\beta\}$, without damaging another popularized edge as a result.
 To address this challenge, we rely on the notion of a u-fan, introduced by [GNK+85], which is the non-bipartite graph analogue of an uncolored edge in bipartite graphs.
@@
 This naturally leads to a win-win situation: if the Vizing chain is short, then $(u,v)$ can be colored efficiently using Vizing’s procedure; otherwise if most Vizing chains are long, then there must be a pair of Vizing chains meeting together after a few steps, so we can shift two uncolored edges to form a u-fan efficiently.
+[…]
+Figure 3: In this picture, we have two different uncolored edges $(u_{1},v_{1}),(u_{2},v_{2})$ such that $\alpha\in\mathsf{miss}_{\chi}(u_{1})\cap\mathsf{miss}_{\chi}(u_{2})$, and their Vizing chains first intersect at edge $(x,y)$ which currently has color $\alpha$ under $\chi$.
+Then we can rotate their Vizing fans and flip part of their Vizing chains to shift $(u_{1},v_{1}),(u_{2},v_{2})$ to $(w_{1},x),(w_{2},x)$ respectively to form a u-fan; note that $\alpha\in\mathsf{miss}_{\chi}(w_{1})\cap\mathsf{miss}_{\chi}(w_{2})$ after this shifting procedure.
 Let us make this a bit more precise.
@@
 Also, assume there are $m_{\alpha}$ edges colored $\alpha$ under $\chi$.
-If most $\{\alpha,\cdot\}$-Vizing chains have length larger than $\Omega(m_{\alpha}/|U_{\alpha}|)$, then on average, two Vizing chains should meet within the first $O(m_{\alpha}/|U_{\alpha}|)$ steps; in this case, we can repeatedly pick two intersecting Vizing chains and create a u-fan by shifting their initial uncolored edges to the intersection of these Vizing chains.
+If most $\{\alpha,\cdot\}$-Vizing chains have length larger than $\Omega(m_{\alpha}/|U_{\alpha}|)$, then on average, two Vizing chains should meet within the first $O(m_{\alpha}/|U_{\alpha}|)$ steps; in this case, we can repeatedly pick two intersecting Vizing chains and create a u-fan by shifting their initial uncolored edges to the intersection of these Vizing chains; see Figure 3 for illustration.
 Given the length of the chains, this takes $O(m_{\alpha}/|U_{\alpha}|)$ time.
@@
 Additionally, we have ignored all algorithmic aspects of (efficiently) finding pairs of intersecting Vizing chains, as well as the corner cases of Vizing fan intersections and fan-chain intersections.
-We defer the discussions on these details to the main body of the paper.
-2 Preliminaries
-We begin by describing the basic notation used throughout the paper in Section 2.1 and then give the classic definitions of Vizing fans and Vizing chains in Section 2.2.
-In Section 3, we introduce the new definitions that provide the basic building blocks for our algorithm.
-2.1 Notation
-Let $G=(V,E)$ be graph on $n$ vertices with $m$ edges and maximum degree $\Delta$ and let $\chi:E\longrightarrow[\Delta+1]\cup\{\bot\}$ be a (partial) $(\Delta+1)$-edge coloring of $G$.
+We defer the discussions on these details to the actual proofs in subsequent sections.
+2 Basic Notation
+Let $G=(V,E)$ be graph on $n$ vertices with $m$ edges and maximum degree $\Delta$ and let $\chi:E\rightarrow[\Delta+1]\cup\{\bot\}$ be a (partial) $(\Delta+1)$-edge coloring of $G$.
 We refer to edges $e\in E$ with $\chi(e)=\bot$ as uncolored.
@@
 We define the length $i$ prefix of the path $P$ to be the path $P_{\leq i}:=e_{1},\dots,e_{i}$.
-Given an uncolored edge $e$ under $\chi$, by extending $\chi$ to $e$ we mean modifying $\chi$ so that $\chi(e)\neq\bot$ without creating new uncolored edges.
-Our algorithm will always work by modifying a partial coloring $\chi$; unless explicitly specified otherwise, every new concept we define (such as u-fans and separable collection) will be defined with respect to this particular partial coloring $\chi$.
-2.2 Vizing fans and chains
-Vizing Fans: We now define the notion of Vizing fans, which has been used extensively in the edge coloring literature [Viz64, GNK+85, Sin19].
+Consider a set $U\subseteq E$ of edges that are uncolored under $\chi$, i.e., $\chi(e)=\bot$ for all $e\in U$.
+We use the phrase “extending $\chi$ to $U$ to mean the following: Modify $\chi$ so as to ensure that $\chi(e)\neq\bot$ for all $e\in U$, without creating any new uncolored edges.
+When the set $U$ consists of a single edge $e$ (i.e., when $U=\{e\}$), we use the phrase “extending $\chi$ to the edge $e$ instead of “extending $\chi$ to $U$.
+Our algorithms will always work by modifying a partial coloring $\chi$; unless explicitly specified otherwise, every new concept we define (such as u-fans and separable collection in Section 5) will be defined with respect to this particular partial coloring $\chi$.
+3 Showcase: Our Algorithm Instantiated on Bipartite Graphs
+In this section, we instantiate our algorithm on bipartite graphs to showcase some of our key insights, and outline a proof of Theorem 3.1 below.
+Theorem 3.1.
+There is a randomized algorithm that, given a bipartite graph $G=(V,E)$ with maximum degree $\Delta$, returns a $\Delta$-edge coloring of $G$ in $\tilde{O}(m)$ time with high probability.
+As explained in Section 1.2, focusing on bipartite graphs allows us to ignore the technical issues that arise while dealing with Vizing fans.
+At the same time, this does not trivialize the main conceptual ideas underpinning our algorithm.
+In particular, we prove Theorem 3.1 via Lemma 3.2 below, which gives a specific color-extension algorithm on bipartite graphs.
+Although near-linear time $\Delta$-edge coloring algorithms on bipartite graphs existed since the 1980s [CH82, COS01], to the best of our knowledge there was no known algorithm for Lemma 3.2 prior to our work.
+Lemma 3.2.
+Let $\chi:E\rightarrow[\Delta]\cup\{\bot\}$ be a partial $\Delta$-edge coloring in a bipartite graph $G=(V,E)$, and let $U\subseteq E$ be a matching of size $\lambda$ such that every edge $e\in U$ is uncolored in $\chi$.
+Furthermore, suppose that we have access to an “auxiliary data structure”, which allows us to detect in $\tilde{O}(1)$ time the two least common colors $\alpha,\beta\in[\Delta]$ in $\chi$.
+[^2] Specifically, for every $\gamma\in[\Delta]\setminus\{\alpha,\beta\}$ and $\gamma^{\prime}\in\{\alpha,\beta\}$, we have $\left|\{e\in E\mid\chi(e)=\gamma\right\}|\geq\left|\{e\in E\mid\chi(e)=\gamma^{\prime}\}\right|$.
+Then there is a randomized algorithm that can extend $\chi$ to $\Omega(\lambda/\Delta)$ many edges of $U$ in $\tilde{O}(m/\Delta)$ time with high probability.
+Let us start by showing that Lemma 3.2 implies Theorem 3.1 easily.
+Proof of Theorem 3.1.
+We follow the strategy outlined in Section 1.2.
+Given a bipartite graph $G$, we first find an Eulerian partition of the graph to partition the edges of $G$ into two subgraphs of maximum degree $\left\lceil\Delta/2\right\rceil$ each, and color them recursively using different colors.
+This leads to a $2\cdot\left\lceil\Delta/2\right\rceil\leq(\Delta+2)$ edge coloring of $G$.
+We then form a partial $\Delta$ edge coloring $\chi$ by uncoloring the two color classes of this $(\Delta+2)$-edge coloring with the fewest edges assigned to them, which leaves us with two edge-disjoint matchings $U_{1}$ and $U_{2}$ to color.
+This is our previously mentioned color extension problem.
+To solve this problem, we apply Lemma 3.2 to $U_{1}$ first to extend $\chi$ to $\Omega(1/\Delta)$ fraction of it, and keep applying this lemma to the remaining uncolored edges of $U_{1}$ until they are all colored.
+We then move to $U_{2}$ in the same exact way and extend $\chi$ to its edges as well, obtaining a $\Delta$-coloring of the entire $G$ as a result.
+The correctness of the algorithm follows immediately by induction.
+The runtime can also be analyzed as follows.
+When coloring $U_{1}$ (or $U_{2}$), each invocation of Lemma 3.2 reduces the number of uncolored edges in $U_{1}$ (or $U_{2}$) by a $(1-\Omega(1/\Delta))$ factor and thus we apply this lemma a total of $O(\Delta\cdot\log{n})$ time.
+Moreover, each application of Lemma 3.2 takes $\tilde{O}(m/\Delta)$ time with high probability.
+Thus, with high probability, it only takes $\tilde{O}(m)$ time to extend the coloring $\chi$ to $U_{1}$ and $U_{2}$ in the color extension problem.
+Hence, the runtime of the algorithm, with high probability, follows the recurrence $T(m,\Delta)\leq 2T(m/2,\Delta/2)+\tilde{O}(m)$, and thus itself is $\tilde{O}(m)$ time.
+Finally, note that with $O(m)$ preprocessing time, we can maintain access to the “auxiliary data structure” throughout the repeated invocations of Lemma 3.2 above: All we need to do is to maintain a counter for each color $\gamma\in[\Delta]$, which keeps track of how many edges in $G$ are currently assigned the color $\gamma$ in $\chi$.
+We maintain these counters in a balanced search tree, and update the relevant counter whenever we change the color assignment of an edge in $G$.
+∎
+The rest of this section is dedicated to the proof of Lemma 3.2.
+3.1 Our Bipartite Color Extension Algorithm in Lemma 3.2
+At a high level, our algorithm for Lemma 3.2 consists of the following three steps.
+  1. 1.
+    Pick the two least common colors $\alpha,\beta\in[\Delta]$ in $\Delta$. This implies that there are at most $O(m/\Delta)$ edges in $G$ that are colored with $\alpha$ or $\beta$ in $\chi$.
+  2. 2.
+    Modify the coloring $\chi$ so that $\Omega(\lambda/\Delta)$ of the edges $(u,v)\in U$ either receive a color under $\chi$, or have $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)$.
+While implementing this step, we ensure that the total number of edges with colors $\alpha$ or $\beta$ remains at most $O(m/\Delta)$.
+  3. 3.
+    Let $\Phi$ denote the set of edges $(u,v)\in U$ with $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)$.
+We call these edges popular.
+We extend $\chi$ to a constant fraction of the edges in $\Phi$, by flipping a set of maximal $\{\alpha,\beta\}$-alternating paths.
+We now formalize the algorithm; the pseudocode is provided in Algorithm 1.
+As input, we are given a bipartite graph $G=(V,E)$, a partial $\Delta$-edge coloring $\chi$ of $G$, and a matching $U\subseteq E$ of uncolored edges of size $\lambda$.
+The algorithm starts by fixing the two least common colors $\alpha,\beta\in[\Delta]$ in $\chi$.
+The main part is the while loop in Algorithm 1, which runs in iterations.
+In each iteration of the while loop, the algorithm samples an edge $e=(u,v)$ from $U$ u.a.r.
+and attempts to either (1) directly extend the coloring $\chi$ to $e$ (see Algorithm 1), which adds $e$ to a set $C\subseteq U$ of colored edges in $U$ or (2) modify $\chi$ so that $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)$—we refer to this as making the edge $(u,v)$ popular—, which adds $e$ to a set $\Phi\subseteq U$ of popular edges (see Algorithm 1).
+The attempt to modifying $\chi$ is done by essentially finding a maximal $\{c_{u},\alpha\}$-alternating path $P_{u}$ starting at $u$ and a $\{c_{v},\beta\}$-alternating path $P_{v}$ starting at $v$ for $c_{u}\in\mathsf{miss}_{\chi}(u)$ and $c_{v}\in\mathsf{miss}_{\chi}(v)$ (see Algorithm 1 and Algorithms 1 and 1).
+The modification itself is done only if $P_{u}$ and $P_{v}$ do not intersect any other popular edge already in $\Phi$.
+We say that the concerned iteration of the while loop FAILS if it chooses an already colored edge in $C$ (Algorithm 1), or modifying the color leads to an already popular edge in $\Phi$ to no longer remain popular (Algorithm 1); otherwise, we say the iteration succeeds.
+As stated earlier, the algorithm maintains a subset $\Phi\subseteq U$ of popular edges, and a subset of edges $C\subseteq U$ that got colored since the start of the while loop.
+Initially, we have $C=\Phi=\varnothing$.
+Thus, the quantity $|\Phi|+|C|$ denotes the number of successful iterations of the while loop that have been performed so far.
+The algorithm performs iterations until $|\Phi|+|C|=\Omega(\lambda/\Delta)$, and then it proceeds to extend the coloring $\chi$ to at least a constant fraction of the edges in $\Phi$ by finding $\{\alpha,\beta\}$-alternating paths for edges in $\Phi$ that admits such paths (see the for loop in Algorithm 1).
+3.2 Analysis of the Bipartite Color Extension Algorithm: Proof of Lemma 3.2
+We start by summarizing a few key properties of Algorithm 1.
+Claim 3.3.
+Throughout the while loop in Algorithm 1, there are at most $O(m/\Delta)$ edges in $G$ that receive either the color $\alpha$ or the color $\beta$, under $\chi$.
+Proof.
+We start with $O(m/\Delta)$ such edges in $G$ and each successful iteration of the while loop increases the number of such edges by $O(1)$, and there are $O(\lambda/\Delta)=O(m/\Delta)$ such iterations.
+∎
+1Let $\alpha,\beta\in[\Delta]$ be the two least common colors in $\chi$ // We have $\alpha\neq\beta$
+2 Initialize $\Phi\leftarrow\varnothing$, $C\leftarrow\varnothing$, and set $\lambda\leftarrow|U|$
+3 while $|\Phi|+|C|<\lambda/(10\Delta)$ do
+4   Sample an edge $e=(u,v)\sim U$ independently and u.a.r.
+5 if $(u,v)\in\Phi\cup C$ then
+6      The iteration FAILS
+7    go to Algorithm 1
+8
+9  Identify (arbitrarily) two colors $c_{u}\in\mathsf{miss}_{\chi}(u)$ and $c_{v}\in\mathsf{miss}_{\chi}(v)$
+10 if $c_{u}=c_{v}$ then
+11      Set $\chi(u,v)\leftarrow c_{u}$ and $C\leftarrow C\cup\{(u,v)\}$
+12    go to Algorithm 1
+13
+14 if $c_{u}=\beta$ or $c_{v}=\alpha$ then
+15      Set $(u,v)\leftarrow(v,u)$ // Now $c_{u}\neq c_{v}$, $c_{u}\neq\beta$, $c_{v}\neq\alpha$ (see Algorithms 1 and 1)
+16
+17  Let $P_{u}$ be the maximal $\{c_{u},\alpha\}$-alternating path starting at $u$ ($P_{u}=\varnothing$ if $\alpha\in\mathsf{miss}_{\chi}(u)$)
+18   Let $P_{v}$ be the maximal $\{c_{v},\beta\}$-alternating path starting at $v$ ($P_{v}=\varnothing$ if $\beta\in\mathsf{miss}_{\chi}(v)$)
+19 if either $P_{u}$ or $P_{v}$ ends at a node that is incident on some edge in $\Phi\setminus\{e\}$ then
+20      The iteration FAILS
+21
+22 else
+23      Modify $\chi$ by flipping the alternating paths $P_{u}$ and $P_{v}$
+24      Set $\Phi\leftarrow\Phi\cup\{(u,v)\}$ // Now $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)$
+25
+[…]
+27 for each edge $e=(u,v)\in\Phi^{\prime}$ do
+28 $\Phi^{\prime}\leftarrow\Phi^{\prime}\setminus\{e\}$
+29   W.l.o.g., suppose that $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)$
+30 if there exists a color $c\in\{\alpha,\beta\}$ such that $c\in\mathsf{miss}_{\chi}(u)\cap\mathsf{miss}_{\chi}(v)$ then
+31    $\chi(u,v)\leftarrow c$
+32
+33 else
+34      Let $P^{\star}_{u}$ be the maximal $\{\alpha,\beta\}$-alternating path starting at $u$
+35      (Since $G$ is bipartite, $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)$, $P^{\star}_{u}$ does not end at $v$)
+36      Modify $\chi$ by flipping the alternating path $P^{\star}_{u}$, and set $\chi(u,v)\leftarrow\beta$
+37    if the path $P^{\star}_{u}$ ends at a node that is incident on some edge in $e^{\prime}\in\Phi^{\prime}\setminus\{e\}$ then
+38       $\Phi^{\prime}\leftarrow\Phi^{\prime}\setminus\{e^{\prime}\}$
+39
+[…]
+Lemma 3.4.
+Throughout the execution of the while loop in Algorithm 1, the following conditions hold: (i) the set $C$ consists of all the edges in $U$ that are colored under $\chi$; (ii) for every edge $(u,v)\in\Phi$, we have $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)$.
+Proof.
+Part (i) of the lemma follows from Algorithm 1 and Algorithm 1.
+For part (ii), consider an edge $e=(u,v)$ that gets added to $\Phi$.
+This happens only after flipping the paths $P_{u}$ and $P_{v}$ in Algorithm 1.
+Just before we execute Algorithm 1, the following conditions hold:
+  • $\alpha\neq\beta$ (see Algorithm 1).
+  • $c_{u}\in\mathsf{miss}_{\chi}(u)$ and $c_{v}\in\mathsf{miss}_{\chi}(v)$ (see Algorithm 1).
+  • $c_{u}\neq c_{v}$, $c_{u}\neq\beta$ and $c_{v}\neq\alpha$ (see Algorithm 1).
+  • If $\alpha\in\mathsf{miss}_{\chi}(u)$ then $P_{u}=\varnothing$ (see Algorithm 1), and if $\beta\in\mathsf{miss}_{\chi}(v)$ then $P_{v}=\varnothing$ (see Algorithm 1).
+  • The path $P_{u}$ (resp. $P_{v}$) does not end at that a vertex that is incident on some edge in $\Phi\setminus\{e\}$ (see Algorithms 1 and 1), although it might possibly end at $v$ (resp. $u$).
+From these conditions, it follows that the paths $P_{u}$ and $P_{v}$ are edge-disjoint, and after we flip them in Algorithm 1, we have $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)$ in Algorithm 1.
+In subsequent iterations of the while loop, the only places where we change the coloring $\chi$ are Algorithms 1 and 1.
+Since the edges in $U$ form a matching, changing the coloring in Algorithm 1 cannot affect whether or not the edge $(u,v)\in\Phi$ remains popular (i.e., has $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)$).
+Finally, during a subsequent iteration of the while loop where we sample an edge $(u^{\prime},v^{\prime})\sim U$, we flip the paths $P_{u^{\prime}},P_{v^{\prime}}$ in Algorithm 1 only if their endpoints are not incident on any edges in $\Phi\setminus\{(u^{\prime},v^{\prime})\}$ (see Algorithm 1), and in particular, on $u$ or $v$.
+Thus, this operation cannot change what colors are available at $u$ and $v$, and so cannot change whether or not the edge $(u,v)\in\Phi$ remains popular.
+∎
+We use the following lemma to bound the number of iterations of the while loop in Algorithm 1.
+Lemma 3.5.
+Each iteration of the while loop in Algorithm 1 increases the value of $|\Phi|+|C|$ by an additive one, with probability at least $1/2$ and otherwise keep it unchanged.
+Proof.
+Fix any given iteration of the while loop.
+At the start of this iteration, we sample an edge from $U$ u.a.r.
+We say that an edge $e\in U$ is bad if the iteration FAILS when we sample $e$ (see Algorithm 1 and Algorithm 1), and good otherwise.
+Note that if we sample a good edge $e\in U$, then the iteration either adds one edge to the set $\Phi$ (see Algorithm 1), or adds one edge to the set $C$ (see Algorithm 1).
+In other words, if we sample a good (resp.
+bad) edge $e\in U$ at the start of the iteration, then this increases in the value of $|\Phi|+|C|$ by one (resp.
+keeps the value of $|\Phi|+|C|$ unchanged).
+We will show that at most $\lambda/2$ edges in $U$ are bad.
+Since $|U|=\lambda$, this will imply the lemma.
+To see why this claimed upper bound on the number of bad edges holds, first note that there are $(|\Phi|+|C|)$ many bad edges that cause the iteration to FAIL in Algorithm 1.
+It now remains to bound the number of bad edges which cause the iteration to FAIL in Algorithm 1.
+Towards this end, note that for each edge $(u^{\prime},v^{\prime})\in\Phi$, there are at most $4\Delta$ many maximal $\{\alpha,\cdot\}$- or $\{\beta,\cdot\}$-alternating paths that end at either $u^{\prime}$ or $v^{\prime}$.
+Furthermore, each such alternating path has its other endpoint incident on at most one edge in $U$ since the edges in $U$ form a matching.
+Thus, for each edge $(u^{\prime},v^{\prime})\in\Phi$, there are at most $4\Delta$ many edges $f_{(u^{\prime},v^{\prime})}\in U$ that satisfy the following condition: Some alternating path constructed by the algorithm after sampling $f_{(u^{\prime},v^{\prime})}$ ends at either $u^{\prime}$ or $v^{\prime}$ (see Algorithm 1 and Algorithm 1).
+Each such edge $f_{(u^{\prime},v^{\prime})}$ is a bad edge which causes the iteration to FAIL in Algorithm 1, and moreover, only such edges are causing the iteration to FAIL in Algorithm 1.
+Thus, the number of such bad edges is at most $|\Phi|\cdot 4\Delta$.
+To summarize, the total number of bad edges is at most $(|\Phi|+|C|)+|\Phi|\cdot 4\Delta<\lambda/2$, where the last inequality holds since $|\Phi|+|C|<\lambda/(10\Delta)$ (see Algorithm 1).
+This concludes the proof.
+∎
+Similarly, we can bound the expected runtime of each iteration of the while loop in Algorithm 1.
+Lemma 3.6.
+Each iteration of the while loop in Algorithm 1 takes $\tilde{O}(m/\lambda)$ time in expectation, regardless of the outcome of previous iterations.
+Proof.
+Alternating path flips can be done in time proportional to the path lengths using standard data structures, so we only need to analyze the path lengths.
+Fix any given iteration of the while loop.
+At the start of this iteration, we can classify the edges in $U$ into one of the following three categories: An edge $e\in U$ is of “Type I” if the iteration ends at Algorithm 1 when we sample $e$, is of “Type II” if the iteration ends at Algorithm 1 when we sample $e$, and is of “Type III” otherwise.
+Let $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$ respectively denote the total number of Type I, Type II and Type III edges, with $\lambda_{1}+\lambda_{2}+\lambda_{3}=\lambda$.
+For every Type III edge $e=(u,v)\in U$, we refer to the alternating paths $P_{u}$ and $P_{v}$ (see Algorithm 1 and Algorithm 1) as the “characteristic alternating paths” for $e$.
+Let $\mathcal{P}_{3}$ denote the collection of characteristic alternating paths of all Type III edges.
+Since the set of Type III edges is a subset of $U$, they form a matching, and hence different paths in $\mathcal{P}_{3}$ have different starting points.
+Furthermore, every path in $\mathcal{P}_{3}$ is either a maximal $\{\alpha,\cdot\}$-alternating path or a maximal $\{\beta,\cdot\}$-alternating path.
+Accordingly, 3.3 implies that the total length of all the paths in $\mathcal{P}_{3}$ is at most $O((m/\Delta)\cdot\Delta)=O(m)$.
+Now, if at the start of the iteration, we happen to sample either a Type I or a Type II edge $e\in U$, then the concerned iteration takes $O(1)$ time.
+In the paragraph below, we condition on the event that the edge $e=(u,v)\in U$ sampled at the start of the iteration is of Type III.
+Using appropriate data structures, the time taken to implement the concerned iteration is proportional (up to $\tilde{O}(1)$ factors) to the lengths of the alternating paths $P_{u}$ and $P_{v}$ (see Algorithm 1 and Algorithm 1).
+The key observation is that for each $x\in\{u,v\}$, the path $P_{x}$ is sampled almost uniformly (i.e., with probability $\Theta(1/\lambda_{3})$) from the collection $\mathcal{P}_{3}$.
+Since the total length of all the paths in $\mathcal{P}_{3}$ is $O(m)$, it follows that the expected length of each of the paths $P_{u},P_{v}$ is $O(m/\lambda_{3})$.
+To summarize, with probability $\lambda_{3}/\lambda$, we sample a Type III edge at the start of the concerned iteration of the while loop, and conditioned on this event the expected time spent on that iteration is $\tilde{O}(m/\lambda_{3})$.
+In contrast, if we sample a Type I or a Type II edge at the start of the concerned iteration, then the time spent on that iteration is $O(1)$.
+This implies that we spend at most $\tilde{O}(m/\lambda_{3})\cdot(\lambda_{3}/\lambda)+O(1)=\tilde{O}(m/\lambda)$ expected time per iteration of the while loop.
+∎
+Finally, we show that in the very last step of the algorithm, the for loop in Algorithm 1, the algorithm succeeds in coloring a constant fraction of popular edges.
+Lemma 3.7.
+The for loop in Algorithm 1 extends the coloring $\chi$ to at least half of the edges in $\Phi$.
+Proof.
+Consider any given iteration of the for loop where we pick an edge $e=(u,v)\in\Phi^{\prime}$ in Algorithm 1, where w.l.o.g.
+$\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)$.
+It is easy to verify that during this iteration, we successfully extend the coloring $\chi$ to $e$, either in Algorithm 1 or in Algorithm 1.
+In the latter case, we crucially rely on the fact that the graph $G$ is bipartite (see Algorithm 1), and hence the maximal $\{\alpha,\beta\}$-alternating path $P^{\star}_{u}$ starting at $u$ cannot end at $v$; in fact, this is the only place where we rely on the biparteness of $G$.
+Algorithms 1 and 1 ensure that the following invariant is satisfied: For every edge $e^{\prime}=(u^{\prime},v^{\prime})\in\Phi^{\prime}$, we have $\alpha\in\mathsf{miss}_{\chi}(u^{\prime})$ and $\beta\in\mathsf{miss}_{\chi}(v^{\prime})$, i.e., the edge $e^{\prime}$ is popular; indeed, Lemma 3.4 implies that this invariant holds just before the for loop starts (see Algorithm 1), and any edge $e^{\prime}\in\Phi^{\prime}$ that may violate this invariant at a later stage, which may only occur due to flipping an alternating path that ends at a node incident on $e^{\prime}$ (in Algorithm 1), is removed from $\Phi^{\prime}$ (in Algorithm 1).
+Now, the lemma follows from the observation that each time we successfully extend the coloring to one edge $e$ in $\Phi^{\prime}$, we remove at most one other edge $e^{\prime}\neq e$ from $\Phi^{\prime}$ (see Algorithm 1), due to the vertex-disjointness of the edges in $U\supseteq\Phi\supseteq\Phi^{\prime}$.
+∎
+We are now ready to conclude the running time analysis of Algorithm 1 and establish the required lower bound on the number of newly colored edges in $U$ under $\chi$ by this algorithm.
+Lemma 3.8.
+Algorithm 1 takes $\tilde{O}(m/\Delta)$ time in expectation and extends $\chi$ to $\Omega(\lambda/\Delta)$ new edges.
+Proof.
+We start with the runtime analysis:
+  • Algorithm 1 can be implemented in $\tilde{O}(1)$ time using the auxiliary data structure, and Algorithms 1 and 1 take constant time.
+  • Next, we bound the running time of the while loop (Algorithm 1).
+For any integer $k\geq 0$, let $T(k)$ denote the expected runtime of while loop if we start the loop under the condition that $|\Phi|+|C|=k$.
+We are interested in $T(0)$ and we know that $T(\lambda/10\Delta)=O(1)$ by the termination condition of the loop.
+By Lemma 3.5 and Lemma 3.6, for any $0<k<\lambda/10\Delta$, we have,
+    $T(k)\leq\tilde{O}(m/\lambda)+\frac{1}{2}\cdot T(k)+\frac{1}{2}\cdot T(k+1),$
+    where we additionally used the monotonicity of $T(\cdot)$, as well as the fact that each while-loop of Algorithm 1 has expected runtime $\tilde{O}(m/\lambda)$ regardless of previous iterations, according to Lemma 3.6.
+Thus, $T(k)\leq T(k+1)+\tilde{O}(m/\lambda)$ and hence $T(0)\leq\lambda/10\Delta\cdot\tilde{O}(m/\lambda)=\tilde{O}(m/\Delta)$.
+  • Finally, since the total number of edges with colors $\alpha$ and $\beta$ just before Algorithm 1 is $O(m/\Delta)$ (see 3.3), the for loop can be implemented in $\tilde{O}(m/\Delta)$ time deterministically in a straightforward manner.
+Thus, the total runtime is $\tilde{O}(m/\Delta)$ in expectation.
+We now establish the bound on the number of newly colored edges.
+When the while loop terminates, we have $|\Phi|+|C|\geq\lambda/(10\Delta)$ (see Algorithm 1), and all the edges in $C$ are colored under $\chi$ (see Lemma 3.4).
+Next, by Lemma 3.7, the for loop in Algorithm 1 further extends the coloring $\chi$ to a constant fraction of the edges in $\Phi$, by only flipping $\{\alpha,\beta\}$-alternating paths.
+Consequently, we get at least $\Omega(\lambda/\Delta)$ newly colored edges in $U$ under $\chi$.
+This concludes the proof.
+∎
+We can now conclude the proof of Lemma 3.2.
+To achieve the algorithm in this lemma, we simply run Algorithm 1 in parallel $\Theta(\log{n})$ times and use the coloring of whichever one finishes first (and terminate the rest at that point).
+This ensures the high probability guarantee of Lemma 3.2 still in $\tilde{O}(m/\Delta)$ runtime.
+This concludes the entire proof.
+3.3 Extension to General Graphs: Roadmap for the Rest of the Paper
+In our Lemma 3.2, we crucially need the graph $G$ to be bipartite while executing Algorithms 1, 1 and 1 in Algorithm 1.
+Otherwise, if $G$ contains odd cycles, then the maximal $\{\alpha,\beta\}$-alternating path $P^{\star}_{u}$ starting from $u$ can end at $v$.
+In that case, the color $\beta$ will no longer be missing at $v$ once we flip the path $P^{\star}_{v}$, and so we will not be able to extend the coloring $\chi$ to the edge $(u,v)$ via $\chi(u,v)\leftarrow\beta$.
+We shall emphasize that this is not a minor technicality, but rather the key reason general graphs are not necessarily $\Delta$ edge colorable and rather require $(\Delta+1)$ colors.
+The standard machinery to deal with this issue is the Vizing fan (see Section 4).
+However, if we try to use Vizing fans inside the framework of Algorithm 1 in a naive manner, then we lose control over one of the colors in the alternating path being flipped while extending the coloring to an edge, leading to a weaker averaging argument and a running time of $\tilde{O}(\Delta m)$ instead of $\tilde{O}(m)$.
+To address this bottleneck, one of our key conceptual contributions is to focus on Vizing fans with respect to an object called a separable collection of u-components (see Section 5).
+Using this concept, in Section 6 we present our algorithmic framework in general graphs.
+Our main result (see Theorem 6.1) relies upon two fundamental subroutines.
+The second subroutine (see Lemma 6.3) generalizes Algorithm 1 presented in this section.
+In contrast, the first subroutine (see Lemma 6.2) either efficiently extends the current coloring to a constant fraction of the uncolored edges, or changes the colors of some edges in the input graph so as to create a situation whereby we can invoke Lemma 6.3.
+We devote Sections 7 and 8 towards proving Lemmas 6.2 and 6.3, respectively.
+For clarity of presentation, we defer the details on supporting data structures to Section 9, such as the use of hash tables or binary search tree data structures for basic operations like picking a missing color at a vertex.
+Finally, Section 10 shows how to optimize our algorithmic framework further and achieve an $O(m\log{n})$ time with high probability.
+4 Preliminaries: Vizing Fans and Vizing Chains
+We now define the notion of Vizing fans, which has been used extensively in the edge coloring literature [Viz64, GNK+85, Sin19].
 Definition 4.1 (Vizing fan).
@@
   3. 3.
-    $\chi(u,v_{1})=\bot$ and $\chi(u,v_{i})=c_{i}$ for all $i>1$.
+    $\chi(u,v_{1})=\bot$ and $\chi(u,v_{i})=c_{i-1}$ for all $i>1$.
   4. 4.
     Either $c_{k}\in\mathsf{miss}_{\chi}(u)$ or $c_{k}\in\{c_{1},\dots,c_{k-1}\}$.
-We say that the Vizing fan $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{0},c_{0}),\dots,(v_{k},c_{k})$ is $\alpha$-primed, has center $u$ and leaves $v_{1},\dots,v_{k}$.
+We say that the Vizing fan $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ is $\alpha$-primed, has center $u$ and leaves $v_{1},\dots,v_{k}$.
 We refer to $c_{i}$ as the color of $v_{i}$ within $\operatorname*{\mathbf{F}}$.
@@
 As input, it takes a vertex $u$ and a color $\alpha\in\mathsf{miss}_{\chi}(u)$, and returns an $\alpha$-primed Vizing fan with center $u$.
-1 For each $x\in V$, let $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)$
+1For each $x\in V$, let $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)$
 2 $k\leftarrow 1$ and $v_{1}\leftarrow v$
@@
 return $(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$
-[…]
+Algorithm 2 $\textnormal{{Vizing-Fan}}(u,v,\alpha)$
 The Algorithm Vizing: We now describe the algorithm Vizing that, given a Vizing fan $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ as input, extends the coloring $\chi$ to the edge $(u,v_{1})$ by building a Vizing chain.
 Algorithm 3 gives a formal description of this procedure.
-1 if $\operatorname*{\mathbf{F}}$ is trivial then
+1if $\operatorname*{\mathbf{F}}$ is trivial then
 2 $\chi(u,v_{i})\leftarrow c_{i}$ for all $i\in[k]$
@@
 6 Extend $\chi$ to $(u,v_{1})$ by flipping the path $P$ and rotating colors in $\operatorname*{\mathbf{F}}$ (details in Lemma 4.2)
-Algorithm 2 Vizing$(\operatorname*{\mathbf{F}})$
+[…]
 Thus, running $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ extends the coloring $\chi$ to the uncolored edge $(u,v_{1})$ by rotating colors in the Vizing fan $\operatorname*{\mathbf{F}}$ and flipping the colors of the alternating path $P$.
@@
 Using standard data structures (see Section 9), we can implement this algorithm to run in time proportional to the number of edges that change their colors, which is $O(\Delta+|P|)$. ∎
-[…]
+Given a Vizing fan $\operatorname*{\mathbf{F}}$, we denote the path $P$ considered by Algorithm 3 by $\texttt{Vizing-Path}(\operatorname*{\mathbf{F}})$.
+If the Vizing fan $\operatorname*{\mathbf{F}}$ is trivial, then $\texttt{Vizing-Path}(\operatorname*{\mathbf{F}})$ denotes an empty path $\varnothing$.
 5 Basic Building Blocks
@@
 For notational convenience, we also say that the u-edge $\operatorname*{\boldsymbol{e}}$ is $\alpha$-primed and define $c_{\operatorname*{\boldsymbol{e}}}(u):=\alpha$ and $c_{\operatorname*{\boldsymbol{e}}}(v)=\bot$.
-[^3] Whenever we refer to an ‘uncolored edge $e$’, we are referring to an edge $e\in E$ such that $\chi(e)=\bot$, whereas a ‘u-edge $\operatorname*{\boldsymbol{e}}$’ always refers to the object from LABEL:{def:u-edge} and is denoted in bold.
+[^4] Whenever we refer to an “uncolored edge ee”, we are referring to an edge $e\in E$ such that $\chi(e)=\bot$, whereas a ‘u-edge $\operatorname*{\boldsymbol{e}}$’ always refers to the object from LABEL:{def:u-edge} and is denoted in bold.
 Collections of U-Components: While working with both u-fans and u-edges simultaneously, we sometimes refer to some $\operatorname*{\boldsymbol{g}}$ that is either a u-fan or a u-edge as a u-component.
@@
     For each $x\in V$, the colors in the multi-set $C_{\mathcal{U}}(x):=\{c_{\operatorname*{\boldsymbol{g}}}(x)\mid\operatorname*{\boldsymbol{g}}\in\mathcal{U},\,x\in\operatorname*{\boldsymbol{g}}\}$ are distinct.
+We remark that the second property of this definition is rather important because we need to ensure that different u-components are not interfering with each other when they share common vertices.
+Check Figure 4 for an illustration.
+[…]
+Figure 4: This picture shows two u-fans $(u_{1},x_{1},x,*,\beta_{1})$ and $(u_{2},x_{2},x,*,\beta_{2})$ sharing a common vertex $x$.
+The separable condition requires that $\beta_{1}\neq\beta_{2}$; for instance $\beta_{1},\beta_{2}$ could be magenta and cyan as shown here.
 Damaged U-Components: Suppose we have a partial $(\Delta+1)$-edge coloring $\chi$ and a separable collection $\operatorname*{\mathcal{U}}$ w.r.t.
@@
 In Section 9, we describe the data structures that we use to implement our algorithm.
-On top of the standard data structures used to maintain the edge coloring $\chi$, which allows us to implement Algorithms 1 and 2 efficiently, we also use data structures that allow us to efficiently maintain a separable collection $\operatorname*{\mathcal{U}}$.
+On top of the standard data structures used to maintain the edge coloring $\chi$, which allows us to implement Algorithms 2 and 3 efficiently, we also use data structures that allow us to efficiently maintain a separable collection $\operatorname*{\mathcal{U}}$.
 More specifically, the data structures that we use to implement a separable collection $\operatorname*{\mathcal{U}}$ support the following queries.
@@
 In response, the data structure removes $\operatorname*{\boldsymbol{g}}$ from $\operatorname*{\mathcal{U}}$ if $\operatorname*{\boldsymbol{g}}\in\operatorname*{\mathcal{U}}$ and outputs fail otherwise.
-[…]
-[…]
+  • $\textsc{Find-Component}_{\operatorname*{\mathcal{U}}}(x,c)$: The input to this query is a vertex $x\in V$ and a color $c\in[\Delta+1]$.
+In response, the data structure returns the u-component $g\in\operatorname*{\mathcal{U}}$ with $c_{\operatorname*{\boldsymbol{g}}}(x)=c$ if such a u-component exists and outputs fail otherwise.
+  • $\textsc{Missing-Color}_{\operatorname*{\mathcal{U}}}(x)$: The input to this query is a vertex $x\in V$.
+In response, the data structure returns an arbitrary color from the set $\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)$.
+[^5] Note that, since $|C_{\operatorname*{\mathcal{U}}}(x)|<|\mathsf{miss}_{\chi}(x)|$, such a color always exists.
 The following claim shows that it is always possible to answer a Missing-Color query.
@@
 These queries provide the ‘interface’ via which our algorithm will interact with the u-components.
-A Remark on Randomization: In order to get good space complexity and running time simultaneously, we implement the standard data structures (used for Algorithms 1 and 2) and the data structure supporting the preceding queries using hashmaps (see Section 7).
+A Remark on Randomization: In order to get good space complexity and running time simultaneously, we implement the standard data structures (used for Algorithms 2 and 3) and the data structure supporting the preceding queries using hashmaps (see Section 9).
 The following proposition describes the hashmaps used by our data structures.
@@
 Let $\operatorname*{\mathcal{U}}$ be a separable collection and $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},c_{1}),\dots,(v_{k},c_{k})$ be a Vizing fan.
-We say that the Vizing fan $\operatorname*{\mathbf{F}}$ if $\operatorname*{\mathcal{U}}$-avoiding if $c_{i}\in\mathsf{miss}_{\chi}(v_{i})\setminus C_{\operatorname*{\mathcal{U}}}(v_{i})$ for each leaf $v_{i}\in\operatorname*{\mathbf{F}}$.
+We say that the Vizing fan $\operatorname*{\mathbf{F}}$ is $\operatorname*{\mathcal{U}}$-avoiding if $c_{i}\in\mathsf{miss}_{\chi}(v_{i})\setminus C_{\operatorname*{\mathcal{U}}}(v_{i})$ for each leaf $v_{i}\in\operatorname*{\mathbf{F}}$.
 We say that a Vizing fan $\operatorname*{\mathbf{F}}$ is a Vizing fan of the u-edge $\operatorname*{\boldsymbol{e}}=(u,v,\alpha)$ if $\operatorname*{\mathbf{F}}$ is $\alpha$-primed, has center $u$ and its first leaf is $v$.
@@
 Proof.
-By 3.5, we can always find a collection of colors $\{\mathsf{clr}(x)\}_{x\in V}$ such that $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)$ for each $x\in V$.
+By 5.5, we can always find a collection of colors $\{\mathsf{clr}(x)\}_{x\in V}$ such that $\mathsf{clr}(x)\in\mathsf{miss}_{\chi}(x)\setminus C_{\operatorname*{\mathcal{U}}}(x)$ for each $x\in V$.
 If we construct a Vizing fan $\operatorname*{\mathbf{F}}$ by calling Vizing-Fan and using such a collection of colors within the algorithm, then it follows that the Vizing fan $\operatorname*{\mathbf{F}}$ is $\operatorname*{\mathcal{U}}$-avoiding.
@@
 6 The Main Algorithm
-Our final algorithm consists of two main components.
+We are now ready to present our main technical result, which is a slightly weaker version of Theorem 1.1, and focuses on achieving a near-linear time algorithm for $(\Delta+1)$ edge coloring (instead of the exact $O(m\log{n})$ time in Theorem 1.1; see the remark after that theorem).
+We will then use this theorem in Section 10 to conclude the proof of Theorem 1.1.
+Theorem 6.1.
+There is a randomized algorithm that, given any simple undirected graph $G=(V,E)$ on $n$ vertices and $m$ edges with maximum degree $\Delta$, finds a $(\Delta+1)$-edge coloring of $G$ in $\tilde{O}(m)$ time with high probability.
+Our main algorithm consists of two main components.
 The first component is an algorithm called Construct-U-Fans that takes a partial $(\Delta+1)$-edge coloring $\chi$ with $\lambda$ uncolored edges and either extends $\chi$ to $\Omega(\lambda)$ of these edges or modifies the coloring to construct a separable collection of $\Omega(\lambda)$ u-fans.
-Lemma 4.1 summarizes the behaviour of this algorithm.
+Lemma 6.2 summarizes the behavior of this algorithm.
 Lemma 6.2.
-Given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ and a set of $\lambda$ uncolored edges $U$, the algorithm Construct-U-Fans does one of the following in $O((m+\Delta\lambda)\log n)$ time:
+Given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ and a set of $\lambda$ uncolored edges $U$, the algorithm Construct-U-Fans does one of the following in $O(m+\Delta\lambda)$ time:
   1. 1.
@@
 The second component is an algorithm called Color-U-Fans that takes a collection of $\lambda$ u-fans and extends the coloring to $\Omega(\lambda)$ of the edges in the u-fans.
-Lemma 4.2 summarizes the behaviour of this algorithm.
+Lemma 6.3 summarizes the behavior of this algorithm.
+The reader may find it helpful to keep in mind that the algorithm for Lemma 6.3 is a generalization of algorithm for Lemma 3.2 in Section 3.
 Lemma 6.3.
 Given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ and a separable collection of $\lambda$ u-fans $\mathcal{U}$, the algorithm Color-U-Fans extends $\chi$ to $\Omega(\lambda)$ edges in $O(m\log n)$ time w.h.p.
-In Sections 5 and 6, we describe these algorithms and prove Lemmas 4.1 and 4.2 respectively.
-Using these lemmas, we now show how to efficiently extend an edge coloring $\chi$ to the remaining uncolored edges.
+In Sections 7 and 8, we prove Lemmas 6.2 and 6.3 respectively. Using these lemmas, we now show how to efficiently extend an edge coloring $\chi$ to the remaining uncolored edges.
 Lemma 6.4.
@@
 Let $U$ denote the set of edges that are uncolored by $\chi$.
-We can then apply Lemma 4.1 to either extend $\chi$ to a constant fraction of the edges in $U$ or construct a separable collection of $\Omega(\lambda)$ u-fans $\mathcal{U}$ in $O((m+\Delta\lambda)\log n)$ time.
+We can then apply Lemma 6.2 to either extend $\chi$ to a constant fraction of the edges in $U$ or construct a separable collection of $\Omega(\lambda)$ u-fans $\mathcal{U}$ in $O(m+\Delta\lambda)$ time.
 In the second case, we can then apply Lemma 6.3 to color $\Omega(\lambda)$ of the edges contained in these u-fans in $O(m\log n)$ time w.h.p.
-Thus, we can extend $\chi$ to a constant fraction of the uncolored edges in $O((m+\Delta\lambda)\log n)$ time w.h.p.
+Thus, we can extend $\chi$ to a constant fraction of the uncolored edges in $O(m\log n+\Delta\lambda)$ time w.h.p.
 After repeating this process $O(\log\lambda)\leq O(\log n)$ many times, no edges remain uncolored.
@@
 ∎
-5 The Algorithm Construct-U-Fans (Proof of Lemma 4.1)
+7 The Algorithm Construct-U-Fans: Proof of Lemma 6.2
 As input, the algorithm Construct-U-Fans is given a graph $G$ and a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges.
@@
 Algorithm 4 gives the pseudocode for this algorithm.
-1 Construct a separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$
+1Construct a separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$
 2 for each $\alpha\in[\Delta+1]$ do
 3 $\mathcal{F}\leftarrow\textnormal{{Prune-Vizing-Fans}}(\operatorname*{\mathcal{U}},\alpha)$
-[…]
+4 $\textnormal{{Reduce-U-Edges}}(\operatorname*{\mathcal{U}},\alpha,\mathcal{F})$
 5
@@
 Algorithm Description: The subroutine Prune-Vizing-Fans scans through the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and constructs $\operatorname*{\mathcal{U}}$-avoiding Vizing fans at these u-edges (see Lemma 5.8), maintaining a vertex-disjoint subset $\mathcal{F}$ of these Vizing fans throughout the process.
-After constructing a Vizing fan $\operatorname*{\mathbf{F}}$ for a u-edge $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$, the subroutine checks if $F$ intersects some other fan in $\mathcal{F}$.
+After constructing a Vizing fan $\operatorname*{\mathbf{F}}$ for a u-edge $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$, the subroutine checks if $\operatorname*{\mathbf{F}}$ intersects some other fan in $\mathcal{F}$.
 If not, it adds $\operatorname*{\mathbf{F}}$ to $\mathcal{F}$.
 Otherwise, it uses these intersecting Vizing fans to either extend the coloring $\chi$ or construct a u-fan, removing the corresponding Vizing fans and u-edges from $\mathcal{F}$ and $\operatorname*{\mathcal{U}}$ respectively.
+See Figure 5 for an illustration.
+[…]
+Figure 5: In this picture, we look at the Vizing fan around an uncolored edge $(u_{1},v_{1})$, and it intersects with the Vizing fan of another edge $(u_{2},v_{2})$ currently residing in $\mathcal{F}$ at vertex $w$.
+Then, we can rotate both Vizing fans and pair these two uncolored edges as a u-fan.
 More formally, let $\operatorname*{\boldsymbol{e}}_{1},\dots,\operatorname*{\boldsymbol{e}}_{\ell}$ denote the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ at the time when the subroutine is called and let $\operatorname*{\boldsymbol{e}}_{i}=(u_{i},v_{i},\alpha)$.
@@
 This is true trivially for $i=1$.
-Now, suppose that this is true for $i-1$.
+Now, suppose that this is true for i1i-1.
 If the $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}_{i}$ of $\operatorname*{\boldsymbol{e}}_{i}$ does not intersect any of the Vizing fans in $\mathcal{F}$, then we add $\operatorname*{\mathbf{F}}_{i}$ to $\mathcal{F}$ and are done.
@@
 This is true trivially for $i=1$, since the subroutine does not change $\chi$ or $\operatorname*{\mathcal{U}}$ while scanning this u-edge.
-Now, suppose that this is true for $i-1$.
+Now, suppose that this is true for i1i-1.
 It follows from the inductive argument in Lemma 7.1 that the Vizing fans currently in $\mathcal{F}$ are all $\operatorname*{\mathcal{U}}$-avoiding.
@@
 The subroutine then either extends the coloring to the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ or uses them to construct u-fans.
-[…]
+Algorithm Description: The subroutine Reduce-U-Edges begins by activating all of the trivial Vizing fans in $\mathcal{F}$ and then removing them from $\mathcal{F}$.
+It then constructs Vizing chains at each of the remaining Vizing fans in $\mathcal{F}$ and explores the alternating paths $\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})$ ‘in parallel’ (meaning, one edge at a time across all these paths).
+The subroutine continues this process until it either (1) reaches the end of one of these paths, or (2) two of these paths intersect.
+In the first case, it identifies the Vizing fan $\operatorname*{\mathbf{F}}\in\mathcal{F}$ whose Vizing chain is fully explored and calls $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ before removing $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$.
+In the second case, it identifies the Vizing fans $\operatorname*{\mathbf{F}},\operatorname*{\mathbf{F}}^{\prime}\in\mathcal{F}$ corresponding to the intersecting Vizing chains and either extends the coloring $\chi$ to one more uncolored edge or creates a u-fan by shifting uncolored edges down these Vizing chains, before removing $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$.
+It repeats this process until at least half of the Vizing fans have been removed from $\mathcal{F}$.
 More formally, the subroutine begins by scanning through all of the trivial Vizing fans $\operatorname*{\mathbf{F}}\in\mathcal{F}$ and calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$, removing the $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ from $\operatorname*{\mathcal{U}}$, and removing any other u-component damaged by calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ from $\operatorname*{\mathcal{U}}$ (see Lemma 5.9).
@@
 The subroutine then proceeds in rounds which consist of updating each of the paths $P_{\operatorname*{\mathbf{F}}}$ while maintaining the following invariant.
-Invariant 5.5.
+Invariant 7.5.
 After updating a path $P_{\operatorname*{\mathbf{F}}}$, the following hold:
   1. 1.
-[…]
+    For each $\operatorname*{\mathbf{F}}\in\mathcal{F}$, $P_{\operatorname*{\mathbf{F}}}$ is the length $|P_{\operatorname*{\mathbf{F}}}|$ prefix of the alternating path $\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})$, which we denote by $\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})_{\leq|P_{\operatorname*{\mathbf{F}}}|}$.
   2. 2.
@@
 The prefix paths $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$ maintained by the algorithm also satisfy the following invariant.
-Invariant 5.6.
+Invariant 7.6.
 At the start of each round, we have that $|P_{\operatorname*{\mathbf{F}}}|=L$ for all $\operatorname*{\mathbf{F}}\in\mathcal{F}$.
+Immediately after starting a round, the subroutine increases the value of $L$ by $1$, invalidating 7.6.
+To restore this invariant, we need to update each of the prefix paths $P_{\operatorname*{\mathbf{F}}}$, which in turn may require us to update the other objects maintained by the subroutine to ensure that the other conditions of 7.5 are satisfied.
+The subroutine does this by scanning through each $\operatorname*{\mathbf{F}}\in\mathcal{F}$ and calling $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$, which we describe formally in Algorithm 5.
+The subroutine performs these rounds until $|\mathcal{F}|\leq\lambda_{\alpha}/2$, at which point it terminates.
 […]
-1 $P_{\operatorname*{\mathbf{F}}}\leftarrow\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})_{\leq|P_{\operatorname*{\mathbf{F}}}|+1}$
-2 Let $(x,y)$ be the $(L+1)^{th}$ edge in $P_{\operatorname*{\mathbf{F}}}$ and $y$ be an endpoint of $P_{\operatorname*{\mathbf{F}}}$
+2 Let $(x,y)$ be the $L^{th}$ edge in $P_{\operatorname*{\mathbf{F}}}$ and $y$ be an endpoint of $P_{\operatorname*{\mathbf{F}}}$
 3 if $(x,y)\in S$ then
@@
 33
-[…]
+Algorithm 5 $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$
 Analysis of Reduce-U-Edges
@@
 Lemma 7.7.
-The subroutine Reduce-U-Edges satisfies 5.5.
+The subroutine Reduce-U-Edges satisfies 7.5.
 Proof.
-[…]
+We show that, as long as 7.5 is satisfied, calling $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ for some $\operatorname*{\mathbf{F}}\in\mathcal{F}$ maintains the invariant.
+We first note that 7.5 is trivially satisfied immediately after initializing $L$, $S$ and $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$.
 After activating and removing all of the trivial Vizing fans in $\mathcal{F}$, Conditions 1-3 are clearly still satisfied (since the paths are all empty).
@@
 This establishes the base case of the induction.
-[…]
+Now, assume that 7.5 is satisfied and suppose that we call $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ for some $\operatorname*{\mathbf{F}}\in\mathcal{F}$.
+We now argue that each condition of 7.5 is satisfied after handling this call.
 Conditions 2 and 3: After extending $P_{\operatorname*{\mathbf{F}}}$ by one more edge, we check if there is some other path $P_{\operatorname*{\mathbf{F}}^{\prime}}$ that intersects the updated path $P_{\operatorname*{\mathbf{F}}}$ at this new edge.
@@
 We now consider the following three cases.
-(1) If we enter the if statement on Algorithm 4, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ and $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}}^{\prime})$ after uncoloring edges on their Vizing chains does not damage any u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$, which are both removed from $\operatorname*{\mathcal{U}}$.
+  1. (1)
+    If we enter the if statement on Algorithm 5, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ and $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}}^{\prime})$ after uncoloring edges on their Vizing chains does not damage any u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$, which are both removed from $\operatorname*{\mathcal{U}}$.
 This is because shifting colors around the Vizing fans $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ does not damage any u-components apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ (see Lemma 5.9) and truncating an alternating path before flipping it ensures that we do not remove any colors from the palettes of the vertices on that path.
 Finally, note that the u-fan $\operatorname*{\boldsymbol{f}}$ that we add to $\operatorname*{\mathcal{U}}$ only uses colors that were previously unavailable.
-(2) If we enter the else statement on Algorithm 4, then the argument is completely analogous to the previous case, except that we do not add any u-components to $\operatorname*{\mathcal{U}}$ (see Lemma 3.9).
-(3) If we enter the if statement on Algorithm 4, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ damages at most one u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ (see Lemma 3.9).
+See the example Figure 6 for an illustration.
+[…]
+    Figure 6: In this picture, $\alpha$ is blue, and two uncolored edges $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}=(u_{1},v_{1}),\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}=(u_{2},v_{2})$ generate Vizing fans $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$, and the two Vizing chains are joining at $(x,y)$ for the first time and in the same direction.
+Then, we can rotate both Vizing fans and partially flip $P_{\operatorname*{\mathbf{F}}}$ and $P_{\operatorname*{\mathbf{F}}^{\prime}}$ to make a u-fan
+  2. (2)
+    If we enter the else statement on Algorithm 5, then the argument is completely analogous to the previous case, except that we do not add any u-components to $\operatorname*{\mathcal{U}}$ (see Lemma 5.9).
+See Figure 7 for an illustration.
+[…]
+    Figure 7: In this picture, $\alpha$ is blue, and two uncolored edges $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}=(u_{1},v_{1}),\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}=(u_{2},v_{2})$ generate Vizing fans $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$, and the two Vizing chains are joining at $(x,y)$ for the first time and in the opposite direction.
+Then, we can uncolor $(x,y)$ and color both $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}},\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$.
+  3. (3)
+    If we enter the if statement on Algorithm 5, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ damages at most one u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ (see Lemma 5.9).
 In particular, if there is such a u-component $\operatorname*{\boldsymbol{g}}$, it will contain the vertex $y$ and have $c_{\operatorname*{\boldsymbol{g}}}(x)=\chi(x,y)$.
@@
 Thus, if $x$ is contained in the Vizing fan $\operatorname*{\mathbf{F}}^{\prime}$, its color within $\operatorname*{\mathbf{F}}^{\prime}$ is not $c_{\operatorname*{\boldsymbol{f}}}(x)$.
-[…]
+Condition 1: Let $\operatorname*{\mathbf{F}}^{\prime}$ be some Vizing fan that remains in $\mathcal{F}$ after handling this call.
+We now show that the first $|P_{\operatorname*{\mathbf{F}}^{\prime}}|$ edges of $\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}}^{\prime})$ do not change throughout this call.
+It follows from the argument for Condition 5 that neither the colors of the edges nor the palettes of the vertices in $\operatorname*{\mathbf{F}}^{\prime}$ change during this call.
+It remains to show that none of the edges in $P_{\operatorname*{\mathbf{F}}^{\prime}}$ change color during this call.
+We can see that $P_{\operatorname*{\mathbf{F}}^{\prime}}$ and $P_{\operatorname*{\mathbf{F}}}$ are edge-disjoint since otherwise $\operatorname*{\mathbf{F}}^{\prime}$ would have been removed from $\mathcal{F}$.
+Consequently, $P_{\operatorname*{\mathbf{F}}^{\prime}}$ is edge-disjoint from any Vizing chain that is activated during the call, and hence none of the edges in $P_{\operatorname*{\mathbf{F}}^{\prime}}$ change colors.
+∎
 Lemma 7.8.
-The subroutine Reduce-U-Edges satisfies 5.6.
+The subroutine Reduce-U-Edges satisfies 7.6.
 Proof.
-[…]
+This follows from the fact that initially $|P_{\operatorname*{\mathbf{F}}}|=0$ for all $\operatorname*{\mathbf{F}}\in\mathcal{F}$ and that the subroutine calls $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ for each $\operatorname*{\mathbf{F}}\in\mathcal{F}$ during each round, which increases $|P_{\operatorname*{\mathbf{F}}}|$ by $1$.
+∎
 Lemma 7.9.
@@
 Proof.
-[…]
+While activating some trivial Vizing fan $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}\in\mathcal{F}$ at the start, we remove $\operatorname*{\boldsymbol{e}}$ from $\operatorname*{\mathcal{U}}$ (and thus also from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$) along with at most one other u-component in $\operatorname*{\mathcal{U}}$ that is damaged by this operation (see Lemma 5.9).
+Now, suppose that the subroutine calls $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ for some $\operatorname*{\mathbf{F}}\in\mathcal{F}$.
+By considering each case, we can verify that the subroutine either makes no changes to $\chi$ and $\operatorname*{\mathcal{U}}$ or it removes at most $2$ u-edges from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and either (1) extends the coloring $\chi$ to one more edge (see Lines 5-5 and Algorithm 5) and removes at most one other u-component from $\operatorname*{\mathcal{U}}$ (see Algorithm 5), or (2) adds a u-fan to $\operatorname*{\mathcal{U}}$ (see Algorithm 5).
+∎
 Running Time: We now show how to implement this subroutine and analyse its running time. We begin with the following claim which shows that Update-Path can be implemented efficiently.
 Claim 7.10.
-[…]
+Each call to $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ for some $\operatorname*{\mathbf{F}}\in\mathcal{F}$ that does not remove $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$ takes $O(1)$ time.
+Otherwise, it takes $O(\Delta+L)$ time.
 Proof.
-[…]
+By Invariants 7.5 and 7.6, we know that $P_{\operatorname*{\mathbf{F}}}=\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})_{\leq L}$ when the subroutine calls $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$.
+Thus, updating $P_{\operatorname*{\mathbf{F}}}$ to $\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})_{\leq L+1}$ only requires computing the next edge in the path, which can be done in $O(1)$ time using our data structures (see Section 9).
 If this call does not remove $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$, then we know that $P_{\operatorname*{\mathbf{F}}}$ is not maximal and is also edge-disjoint from the other prefix paths maintained by the subroutine.
@@
 ∎
-Let $\mathcal{F}^{\star}$ denote the set $\mathcal{F}$ when we first call the subroutine Reduce-U-Edges.
+Let $\mathcal{F}^{\star}$ denote the subset of Vizing fans that get removed from $\mathcal{F}$ by the subroutine Reduce-U-Edges before it terminates.
 For each $\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}$, let $L_{\operatorname*{\mathbf{F}}}$ denote the value of $L$ at the time that $\operatorname*{\mathbf{F}}$ is removed from $\mathcal{F}$.
@@
 Proof.
-[…]
+Let $\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}$.
+We can observe that the subroutine calls $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ at most $L_{\operatorname*{\mathbf{F}}}$ times.
+It follows from 7.10 that the last call takes $O(\Delta+L_{\operatorname*{\mathbf{F}}})$ time while the rest take $O(1)$ time.
+Thus, the total time spent handling calls to $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ is at most $O(\Delta+L_{\operatorname*{\mathbf{F}}})$.
+Summing over each $\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}$, we get that the total time spent handling calls to Update-Path is at most
 $\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}O(\Delta+L_{\operatorname*{\mathbf{F}}})\leq O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}L_{\operatorname*{\mathbf{F}}}.\qed$
-Now, assume w.l.o.g.
-that =$\mathcal{F}^{\star}=\{\operatorname*{\mathbf{F}}_{1},\dots,\operatorname*{\mathbf{F}}_{\lambda_{\alpha}}\}$ and $L_{\lambda_{\alpha}}\leq\dots\leq L_{1}$, where $L_{i}:=L_{\operatorname*{\mathbf{F}}_{i}}$.
 Recall that $m_{\alpha}$ denotes the number of edges with color $\alpha$ when we first call the subroutine.
 Claim 7.12.
-For each $i\in[\lambda_{\alpha}]$, we have that $L_{i}\leq 4(m_{\alpha}+2\lambda_{\alpha})/i$.
+For each $\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}$, we have that $L_{\operatorname*{\mathbf{F}}}\leq O(m_{\alpha}/\lambda_{\alpha})$.
 Proof.
-Let $i\in[\lambda_{\alpha}]$, $\tau=4(m_{\alpha}+2\lambda_{\alpha})/i$ and suppose that $\tau<L_{i}$.
-Then, at the start of iteration $\tau+1$, the Vizing fans $\operatorname*{\mathbf{F}}_{1},\dots\operatorname*{\mathbf{F}}_{i}$ are all contained in $\mathcal{F}$.
-Furthermore, it follows from Invariants 5.5 and 5.6 that, at the start of iteration $\tau+1$, the alternating paths $P_{\operatorname*{\mathbf{F}}_{1}},\dots,P_{\operatorname*{\mathbf{F}}_{i}}$ form a collection of $i$ edge-disjoint $\{\alpha,\cdot\}$-alternating paths of length $\tau$ in $G$.
+Let $L_{\max}$ denote the value of $L$ at the start of the final round performed by the subroutine.
+We now show that $L_{\max}\leq 8(m_{\alpha}+4)/\lambda_{\alpha}$, which implies the claim.
+At the start of the final round, we know that $|\mathcal{F}|>\lambda_{\alpha}/2$, otherwise the subroutine would terminate.
+Furthermore, it follows from Invariants 7.5 and 7.6 that, at the start of this round, the alternating paths $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$ form a collection of at least $\lambda_{\alpha}/2$ edge-disjoint $\{\alpha,\cdot\}$-alternating paths of length $L_{\max}$ in $G$.
 Let $T$ denote the total length of these paths.
@@
 Thus, it follows that
-$4(m_{\alpha}+2\lambda_{\alpha})=i\tau\leq T\leq 3m_{\alpha}^{\prime}\leq 3(m_{\alpha}+2\lambda_{\alpha}),$
-giving a contradiction. ∎
+$\frac{\lambda_{\alpha}}{2}\cdot L_{\max}\leq T\leq 3m_{\alpha}^{\prime}\leq 3(m_{\alpha}+2\lambda_{\alpha}),$
+and so $L_{\max}\leq 6m_{\alpha}/\lambda_{\alpha}+12$. ∎
 Lemma 7.13.
-We can implement Reduce-U-Edges to run in $O((m_{\alpha}+\Delta\lambda_{\alpha})\log n)$ time.
+We can implement Reduce-U-Edges to run in $O(m_{\alpha}+\Delta\lambda_{\alpha})$ time.
 Proof.
@@
 The running time of the rest of the subroutine is dominated by the time taken to handle the call to Update-Path.
-Combining Claims 5.11 and 5.12, it follows that the total time spent handling calls to Update-Path is at most
-$O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}L_{\operatorname*{\mathbf{F}}}\leq O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{i=1}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}$
-$\leq O((m_{\alpha}+\Delta\lambda_{\alpha})\log n).\qed$
-5.3 Analysis of Construct-U-Fans (Proof of Lemma 4.1)
+Combining Claims 7.11 and 7.12, it follows that the total time spent handling calls to Update-Path is at most
+$O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}L_{\operatorname*{\mathbf{F}}}\leq O(\Delta\lambda_{\alpha})+O(\lambda_{\alpha})\cdot O\!\left(\frac{m_{\alpha}}{\lambda_{\alpha}}\right)\leq O(m_{\alpha}+\Delta\lambda_{\alpha}).\qed$
+7.3 Analysis of Construct-U-Fans: Proof of Lemma 6.2
 The algorithm begins by constructing a separable collection of $\lambda$ u-components $\operatorname*{\mathcal{U}}$. The following lemma shows that this can be done efficiently.
@@
 It then scans through the edges of the graph and retrieves the $\lambda$ edges $e_{1},\dots,e_{\lambda}$ that are uncolored by $\chi$.
-The algorithm then scans through each of these uncolored edges and, for each $e_{i}=(u_{i},v_{i})$, picks a missing color $\alpha_{i}\in\mathsf{miss}_{\chi}(u_{i})\setminus C_{\operatorname*{\mathcal{U}}}(u_{i})$ (see 3.5) and adds the u-edge $\operatorname*{\boldsymbol{e}}_{i}:=(u_{i},v_{i},\alpha_{i})$ to $\operatorname*{\mathcal{U}}$.
+The algorithm then scans through each of these uncolored edges and, for each $e_{i}=(u_{i},v_{i})$, picks a missing color $\alpha_{i}\in\mathsf{miss}_{\chi}(u_{i})\setminus C_{\operatorname*{\mathcal{U}}}(u_{i})$ (see 5.5) and adds the u-edge $\operatorname*{\boldsymbol{e}}_{i}:=(u_{i},v_{i},\alpha_{i})$ to $\operatorname*{\mathcal{U}}$.
 It’s easy to verify that the resulting collection $\operatorname*{\mathcal{U}}$ is separable and contains $\lambda$ u-edges.
@@
 Note that the number of such u-edges can only decrease throughout the run of this algorithm.
-It follows from Lemmas 5.4 and 5.13 that the total running time of the algorithm Construct-U-Fans across all of these calls to the subroutines Prune-Vizing-Fans and Reduce-U-Edges is at most
-$\sum_{\alpha\in[\Delta+1]}\left(O(\lambda_{\alpha}\Delta)+O((m_{\alpha}+\lambda+\lambda_{\alpha}\Delta)\log n)\right)\leq O((m+\Delta\lambda)\log n).$
+It follows from Lemmas 7.4 and 7.13 that the total running time of the algorithm Construct-U-Fans across all of these calls to the subroutines Prune-Vizing-Fans and Reduce-U-Edges is at most
+$\sum_{\alpha\in[\Delta+1]}\left(O(\lambda_{\alpha}\Delta)+O(m_{\alpha}+\lambda+\lambda_{\alpha}\Delta)\right)\leq O(m+\Delta\lambda).$
 Lemma 7.16.
-The algorithm Construct-U-Fans either extends the coloring $\chi$ to at least $\lambda/9$ more edges or returns a separable collection of at least $\lambda/9$ u-fans $\operatorname*{\mathcal{U}}$.
+The algorithm Construct-U-Fans either extends the coloring $\chi$ to at least $\lambda/18$ more edges or returns a separable collection of at least $\lambda/18$ u-fans $\operatorname*{\mathcal{U}}$.
 Proof.
@@
 Immediately after constructing the separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$, it holds that $\Psi_{\texttt{c}}=0$, $\Psi_{\texttt{e}}=\lambda$ and $\Psi_{\texttt{f}}=0$.
-We next argue that by the time the execution of the algorithm has finished, either $\Psi_{\texttt{c}}\geq\lambda/9$ or $\Psi_{\texttt{f}}\geq\lambda/9$ must hold.
+We next argue that by the time the execution of the algorithm has finished, either $\Psi_{\texttt{c}}\geq\lambda/18$ or $\Psi_{\texttt{f}}\geq\lambda/18$ must hold.
 To this end, we employ the following potential function argument.
@@
 7.17 implies that $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}\geq\lambda$ holds at all times afterwards.
-Since the algorithm iterates until $\mathcal{E}=\varnothing$, we know that $\Psi_{\texttt{e}}=0$ when the algorithm’s execution terminates, so at that moment we have $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})\geq\lambda$.
-Hence, either $\Psi_{\texttt{f}}\geq\lambda/9$ or $\Psi_{\texttt{c}}\geq\lambda/9$ must hold.
+Since we have that $|\mathcal{E}|\leq\lambda/2$ when the algorithm terminates, we know that $\Psi_{\texttt{e}}\leq\lambda/2$ at this time, so at that moment we have $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})\geq\lambda/2$.
+Hence, either $\Psi_{\texttt{f}}\geq\lambda/18$ or $\Psi_{\texttt{c}}\geq\lambda/18$ must hold.
 ∎
-6 The Algorithm Color-U-Fans (Proof of Lemma 4.2)
+8 The Algorithm Color-U-Fans: Proof of Lemma 6.3
 As input, the algorithm Color-U-Fans is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, and a collection of separable u-fans $\mathcal{U}$ of size $\lambda$.
@@
 Algorithm 6 gives the pseudocode for Color-U-Fans.
-1 for $\Delta/2$ iterations do
+1for $\Delta/2$ iterations do
 2   Let $\alpha,\beta\in[\Delta+1]$ be the two least common colors in $\chi$
-[…]
-[…]
+3 $\textnormal{{Prime-U-Fans}}(\operatorname*{\mathcal{U}},\alpha,\beta)$
+4 $\textnormal{{Activate-U-Fans}}(\operatorname*{\mathcal{U}},\alpha,\beta)$
 5
-[…]
+Algorithm 6 $\textnormal{{Color-U-Fans}}(\mathcal{U})$
 Organization of Section 8: We begin by describing and analyzing the subroutines Prime-U-Fans and Activate-U-Fans used by Color-U-Fans before proving Lemma 6.3 in Section 8.3.
 8.1 The Subroutine Prime-U-Fans
-As input, this subroutine is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, a collection of separable u-fans $\mathcal{U}$ of size $\lambda$, and the two least common colors $\alpha,\beta\in[\Delta+1]$ (see 6.1).
+As input, this subroutine is given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$, a collection of separable u-fans $\mathcal{U}$ of size $\lambda$, and the two least common colors $\alpha,\beta\in[\Delta+1]$ (see 8.1).
 It then proceeds in iterations, where in each iteration it samples a u-fan $\operatorname*{\boldsymbol{f}}$ from $\mathcal{U}$ uniformly at random and attempts to prime $\operatorname*{\boldsymbol{f}}$ with the colors $\alpha$ and $\beta$.
@@
 The pseudocode in Algorithm 7 gives a formal description of the subroutine Prime-U-Fans.
-1 $\Phi\leftarrow\varnothing$ and $\lambda\leftarrow|\mathcal{U}|$
+[…]
 2 while $|\Phi|<\lambda/(48\Delta)$ do
@@
 18return $(\alpha,\beta)$
-[…]
+Algorithm 7 $\textnormal{{Prime-U-Fans}}(\mathcal{U})$
 Analysis of Prime-U-Fans
 We say that an iteration of (the while loop in) Prime-U-Fans is successful if it adds a u-fan to the set $\Phi$.
-Otherwise, we say that the iteration fails.
+Otherwise, we say that the iteration fails (see Figure 8 for an illustration).
 Note that the algorithm repeatedly performs iterations until it has performed at least $\lambda/(48\Delta)$ successful iterations.
 The following lemmas summarize the main properties of the subroutine Prime-U-Fans.
+[…]
+Figure 8: In this picture, $\alpha$ is red, $\beta$ is orange, $\gamma$ is blue, and $\delta$ is green.
+$(u^{\prime},v^{\prime},w^{\prime})\in\Phi$ is an existing $\{\alpha,\beta\}$-primed u-fan in $\Phi$.
+If we flip the $\{\alpha,\gamma\}$-alternating path at $u$ and the $\{\beta,\delta\}$-alternating paths at $v$ and $w$, then we would destroy the property that $\beta\in\mathsf{miss}_{\chi}(v^{\prime})$.
 Claim 8.1.
@@
 Proof.
-Suppose that the path $P$ is a $\{c,c^{\prime}\}$-alternating path for some colors $c\in\{\alpha,\beta\}$ and $c^{\prime}\notin\{\alpha,\beta\}$ with endpoints $x$ and $y$ [^10] Note that the algorithm never flips $\{\alpha,\beta\}$-alternating paths.
+Suppose that the path $P$ is a $\{c,c^{\prime}\}$-alternating path for some colors $c\in\{\alpha,\beta\}$ and $c^{\prime}\notin\{\alpha,\beta\}$ with endpoints $x$ and yy.
+[^11] Note that the algorithm never flips $\{\alpha,\beta\}$-alternating paths.
 Since the collection of u-fans $\mathcal{U}$ is separable, we know that at most one u-fan in $\mathcal{U}$ containing $x$ (resp.
@@
 $\mathbb{E}[\textnormal{{cost}}(\operatorname*{\boldsymbol{f}})]=\frac{1}{|\mathcal{U}|}\cdot\sum_{\operatorname*{\boldsymbol{f}}\in\operatorname*{\mathcal{U}}}\sum_{x\in\operatorname*{\boldsymbol{f}}}|P_{x}(\operatorname*{\boldsymbol{f}})|\leq\frac{2}{|\mathcal{U}|}\cdot\sum_{P\in\mathcal{P}}|P|\leq\frac{32m}{\lambda},$
-where we are using the facts that (1) each path in $P\in\mathcal{P}$ appears at most twice while summing over the paths $|P_{x}(\operatorname*{\boldsymbol{f}})|$ by 6.6, (2) that $|\mathcal{U}|\geq\lambda/2$ by 6.2, and (3) that that the total length of all paths in $\mathcal{P}$ is $8m$ by 6.4.
+where we are using the facts that (1) each path in $P\in\mathcal{P}$ appears at most twice while summing over the paths $|P_{x}(\operatorname*{\boldsymbol{f}})|$ by 8.6, (2) that $|\mathcal{U}|\geq\lambda/2$ by 8.2, and (3) that that the total length of all paths in $\mathcal{P}$ is $8m$ by 8.4.
 Applying Markov’s inequality, it follows that
@@
 Since each u-fan in $\Phi$ has 3 vertices and at most $2\Delta$ $\{\alpha,\cdot\}$- or $\{\beta,\cdot\}$-alternating paths end at each of these vertices, we know that there are at most $6\Delta|\Phi|$ alternating paths that could be considered during some iteration that end at a u-fan in $\Phi$.
-By 6.6, there are at most $12\Delta|\Phi|$ u-fans in $\mathcal{U}$ that could cause the algorithm to consider one of these paths.
+By 8.6, there are at most $12\Delta|\Phi|$ u-fans in $\mathcal{U}$ that could cause the algorithm to consider one of these paths (see Figure 9 for an illustration).
 Thus, at least $|\mathcal{U}|-12\Delta|\Phi|$ of the u-fans in $\mathcal{U}$ are contained in $\mathcal{U}^{\star}$.
@@
 $|\mathcal{U}^{\star}|\geq|\mathcal{U}|-12\Delta|\Phi|\geq|\mathcal{U}|-\frac{\lambda}{4}\geq\frac{|\mathcal{U}|}{2},$
-where we are using the facts that $|\Phi|\leq\lambda/(48\Delta)$ and $|\mathcal{U}|\geq\lambda/2$ from 6.2.
+where we are using the facts that $|\Phi|\leq\lambda/(48\Delta)$ and $|\mathcal{U}|\geq\lambda/2$ from 8.2.
 Since $\operatorname*{\boldsymbol{f}}$ is sampled uniformly at random from $\operatorname*{\mathcal{U}}$, it follows that $\Pr[\operatorname*{\boldsymbol{f}}\in\mathcal{U}^{\star}]\geq 1/2$.
 The lemma follows by applying a union bound.
-∎
+[…]
+Figure 9: In this picture, $\alpha$ is red, $\beta$ is orange, and $\operatorname*{\boldsymbol{f}}^{\prime}=(u^{\prime},v^{\prime},w^{\prime},\alpha,\beta)\in\Phi$ is a u-fan, and we have drawn 99 different $\{\alpha,\cdot\}$- or $\{\beta,\cdot\}$-alternating paths starting from $\operatorname*{\boldsymbol{f}}^{\prime}$.
 Lemma 8.7.
@@
 The pseudocode in Algorithm 8 gives a formal description of the subroutine.
-1 while $\Phi\neq\varnothing$ do
+1while $\Phi\neq\varnothing$ do
 2   Let $\operatorname*{\boldsymbol{f}}\in\Phi$
@@
 6
-[…]
+Algorithm 8 $\textnormal{{Activate-U-Fans}}(\mathcal{U},\alpha,\beta)$
 Analysis of Activate-U-Fans
@@
 ∎
-6.3 Analysis of Color-U-Fans (Proof of Lemma 4.2)
+8.3 Analysis of Color-U-Fans: Proof of Lemma 6.3
 Given a separable collection of $\lambda$ u-fans $\mathcal{U}$, the algorithm Color-U-Fans repeatedly calls Prime-U-Fans and Activate-U-Fans with the set $\mathcal{U}$ as described in Algorithm 6.
 It repeats this process for $\Delta/2$ iterations.
-It follows from Claims 6.2 and 6.11 that $|\mathcal{U}|$ decreases by at most a $(1-1/\Delta)$ factor during each iteration.
+It follows from Claims 8.2 and 8.11 that $|\mathcal{U}|$ decreases by at most a $(1-1/\Delta)$ factor during each iteration.
 Thus, by Bernoulli’s inequality, we get that
@@
 We first describe the data structures and then show how they can be used to efficiently implement the queries described in Section 5.2.
-Implementing an Edge Coloring: Let $G=(V,E)$ and a graph of maximum degree $\Delta$ and let $\mathcal{C}:=[\Delta+1]\cup\{\bot\}$.
+Implementing an Edge Coloring: Let $G=(V,E)$ be a graph of maximum degree $\Delta$ and let $\mathcal{C}:=[\Delta+1]\cup\{\bot\}$.
 We implement an edge coloring $\chi:E\longrightarrow\mathcal{C}$ of $G$ using the following:
@@
   • The set $\chi^{-1}(c):=\{e\in E\mid\chi(e)=c\}$, for all $c\in\mathcal{C}$.
-  • The set $\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$, for all $u\in V$ [^11] We take this intersection with $[\deg_{G}(u)+1]$ instead of maintaining $\mathsf{miss}_{\chi}(u)$ directly to ensure that the space complexity and initialization time of the data structures are $\tilde{O}(m)$ and not $\Omega(\Delta n)$.
+  • The set $\mathsf{miss}_{\chi}(u)\cap[\deg_{G}(u)+1]$, for all $u\in V$.
+[^12] We take this intersection with $\deg_{G}(u)+1]$ instead of maintaining $\mathsf{miss}_{\chi}(u)$ directly to ensure that the space complexity and initialization time of the data structures are $\tilde{O}(m)$ and not $\Omega(\Delta n)$.
 We implement all of the maps and sets using hashmaps, allowing us to perform membership queries, insertions and deletions in $O(1)$ time (see Proposition 5.6).
@@
 Since $\operatorname*{\mathcal{U}}$ is separable, the u-components in $\operatorname*{\mathcal{U}}$ are edge-disjoint, and thus $|\mathcal{U}|\leq m$.
-It follows that map $\psi$ can be stored with $O(m)$ space.
+It follows that the map $\psi$ can be stored with $O(m)$ space.
 For each $u\in V$, we can observe that $|C_{\operatorname*{\mathcal{U}}}(u)|\leq\deg_{G}(u)$ since at most $\deg_{G}(u)$ many u-components in $\operatorname*{\mathcal{U}}$ contain the vertex $u$, and $|\overline{C}_{\operatorname*{\mathcal{U}}}(u)|\leq\deg_{G}(u)+1$ since $\overline{C}_{\operatorname*{\mathcal{U}}}(u)\subseteq[\deg_{G}(u)+1]$.
@@
 If so, we can update $\psi$ and the sets $C_{\operatorname*{\mathcal{U}}}(x)$ and $\overline{C}_{\operatorname*{\mathcal{U}}}(x)$ for $x\in\operatorname*{\boldsymbol{g}}$ in $O(1)$ time to remove $\operatorname*{\boldsymbol{g}}$ from $\operatorname*{\mathcal{U}}$.
-[…]
-[…]
-Acknowledgements
-Part of this work was conducted while the first and second named authors were visiting the Simons Institute for the Theory of Computing as part of the Sublinear Algorithms program.
-Appendix A Logarithmic Factors and Small Values of $\Delta$
-The main contribution and the primary message of our work is presenting the first near-linear time algorithm for $(\Delta+1)$-edge coloring and polynomial time factor improvements over prior work.
-As such, in the presentation of our algorithm and its analysis, we made no attempt to optimize the logarithmic factors, which led to an algorithm with $O(m\log^{3}{n})$ time for finding a $(\Delta+1)$-edge coloring with high probability.
-Nevertheless, we can indeed optimize our algorithm to achieve a runtime of only $O(m\log{\Delta})$ time (with high probability), which matches the classical bound of [CH82, COS01] for $\Delta$-coloring bipartite graphs (although that algorithm is deterministic).
-Our algorithm is also related to a recent line of work in [BD23, BD24, Dha24] that focuses on the $\Delta=n^{o(1)}$ case; our result improves prior $O(m\Delta^{4}\log\Delta)$ bounds of [BD24] exponentially in terms of dependence on $\Delta$.
-Theorem A.1.
-There is a randomized algorithm that, given any simple undirected graph $G$ on $n$ vertices and $m$ edges with maximum degree $\Delta$, finds a $(\Delta+1)$-edge coloring of $G$ in $O(m\log{\Delta})$ time with high probability.
+Implementing $\textsc{Find-Component}_{\operatorname*{\mathcal{U}}}(x,c)$: We make a query to $\psi$ by checking if there is an element $\psi_{x}(c)$.
+If no such element is contained in $\psi$, then return fail.
+Otherwise, return the unique u-component in the set $\psi_{x}(c)$.
+Implementing $\textsc{Missing-Color}_{\operatorname*{\mathcal{U}}}(x)$: Return an arbitrary color from the set $\overline{C}_{\operatorname*{\mathcal{U}}}(x)$.
+10 The Final Algorithm: Proof of Theorem 1.1
+Up until this point in the presentation of our algorithm and its analysis, we made no attempt in optimizing the logarithmic runtime factors, which led to an algorithm with $O(m\log^{3}{n})$ time for finding a $(\Delta+1)$-edge coloring with high probability in Theorem 6.1.
+We now show that we can further optimize the algorithm and obtain an $O(m\log{n})$ time algorithm and conclude the proof of Theorem 1.1.
 Let us first start by listing where the $\log(n)$-terms come from in the proof of Theorem 6.1:
   1. 1.
-    In Lemma 4.1, we are losing an $O(\log{n})$ factor in the subroutine Reduce-U-Edges while turning u-edges into u-fans or coloring them.
-Specifically, in the proof of Lemma 5.13, we lose an additional $O(\log{n})$ factor when handling all u-edges (see the Harmonic series in the proof).
+    In Lemma 6.3, when coloring u-fans, we are losing an $O(\log{n})$ factor to ensure the probabilistic guarantees of the algorithm hold with high probability in each step (see Lemma 8.8).
   2. 2.
-    In Lemma 4.2, when coloring u-fans, we are losing an $O(\log{n})$ factor to ensure the probabilistic guarantees of the algorithm hold with high probability in each step (see Lemma 6.8).
+    Our application of Lemmas 6.2 and 6.3 can color a constant fraction of remaining uncolored edges, hence, we need to run them $O(\log{n})$ times to color all uncolored edges (see Lemma 6.4).
   3. 3.
-    Our application of Lemmas 4.1 and 4.2 can color a constant fraction of remaining uncolored edges, hence, we need to run them $O(\log{n})$ times to color all uncolored edges (see Lemma 4.3).
-  4. 4.
     And finally, the entire framework of reducing $(\Delta+1)$ coloring to extending the coloring to $O(m/\Delta)$ uncolored edges using Eulerian partition technique (in the proof of Theorem 6.1) leads to a recursion depth of $O(\log{n})$ leading to another $O(\log{n})$ overhead in the runtime.
-All in all, these factors lead to our $O(m\log^{3}{n})$ time algorithm (note that the losses in Steps 1 and 2 do not compound as we run these algorithms sequentially and thus pay an additive factor of $O(\log{n})$ between the two of them, not a multiplicative one).
-We now show how these log factors can be removed or replaced by $O(\log{\Delta})$ in the following.
-Throughout, we continue our assumption in Section 3.2 that the high probability events of the hash maps used in the algorithm happen and we ignore their randomness (e.g., if an algorithm is deterministic modulo use of hash maps, we simply refer to it as deterministic here).
-Step 1: The following lemma shows the $O(\log{n})$-term of Lemma 4.1 can be easily circumvented.
-Lemma A.2 (A slight modification of Lemma 4.1).
-There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$, and $\lambda$ uncolored edges, does one of the following in $O(m+\lambda\cdot\Delta)$ time deterministically:
-  1. 1.
-    Extends the coloring to $\Omega(\lambda)$ uncolored edges.
-  2. 2.
-    Modifies $\chi$ to obtain a separable collection of $\Omega(\lambda)$ u-fans.
-Proof Sketch.
-The $O(\log{n})$ term of the runtime in the proof of Lemma 4.1 comes from the Harmonic series calculation in the proof of Lemma 5.13.
-Specifically, for any color $\alpha$, assuming we start with $\lambda_{\alpha}$ $\alpha$-primed u-edges, by 5.12, the runtime needed for “handling” the $i^{th}$ such edge is $O((m_{\alpha}+\lambda_{\alpha})/i)$ time (by handling, we mean either coloring this edge or turning it into a u-fan).
-In the proof of Lemma 4.1, using 5.11, this results in a total runtime of
-$O(\Delta\cdot\lambda_{\alpha})+O(1)\cdot\sum_{i=1}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}=O(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\log\lambda_{\alpha})=O(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\log{n}).$
-However, a simple observation is that we do not need to handle all the $\lambda_{\alpha}$ $\alpha$-primed u-edges given that our final goal is to extend the coloring or create u-fans out of only a constant fraction of the initial $\lambda$ uncolored edges.
-Specifically, suppose our goal is to reduce the number of $\alpha$-primed u-edges to some $\gamma\cdot\lambda_{\alpha}$ for any arbitrary constant $\gamma\in(0,1)$; then, the runtime will be
-$O(\Delta\cdot\lambda_{\alpha})+O(1)\cdot\sum_{i=\gamma\cdot\lambda_{\alpha}}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}=O\left(\Delta\cdot\lambda_{\alpha}+(m_{\alpha}+\lambda_{\alpha})\cdot\log{\left(\frac{\lambda_{\alpha}}{\gamma\cdot\lambda_{\alpha}}\right)}\right)=O(\Delta\cdot\lambda_{\alpha}+m_{\alpha}),$
-given that $\gamma$ is a constant.
-As such, the time it takes to handle any arbitrary large constant fraction of u-edges will be only $O(m)$ (by summing over choices of $\alpha\in[\Delta+1]$).
-The rest of the proof follows exactly as in Lemma 4.1 by simply removing the small constant fraction of u-edges not handled from the graph for the rest of the proof.
-∎
-Step 2: We remove the extra $O(\log{n})$-term of Lemma 4.2 in a similar way although by focusing on the expected runtime of the algorithm (plus a crucial tail inequality).
-In the last step, we show how to recover the final result even from this weaker guarantee.
-Lemma A.3 (A slight modification of Lemma 4.2).
+All in all, these factors led to the $O(m\log^{3}{n})$ bound of our algorithm in Theorem 6.1.
+We now show how these log factors can be reduced to a single one.
+The two main ideas are: (1) relaxing the requirement of Lemma 6.3 so that its runtime holds in expectation, plus a suitable tail bound.
+Then, instead of maintaining high probability bound on each invocation of this lemma, we only bound the runtime of all invocations of this lemma together with high probability; (2) relaxing the Eulerian partition technique to color most of the graph recursively, instead of the entire graph, and then taking care of the remaining uncolored edges at the end.
+We start by presenting a more fine-grained version of two of our main technical lemmas in the proof of Theorem 6.1 (Lemma 6.3 and Lemma 6.4; recall that Lemma 6.2 already runs in linear time deterministically), which correspond to part (1) above, and then use these to present part (2) of above and conclude the proof.
+10.1 Fine-Grained Variants of  Lemma 6.3 and Lemma 6.4
+We remove the $O(\log{n})$-term of Lemma 6.3 by focusing on the expected runtime of the algorithm (plus a crucial tail inequality).
+In the next part, we show how to recover the final result even from this weaker guarantee.
+Lemma 10.1 (A slight modification of Lemma 6.3).
 There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ and a separable collection of $\lambda$ u-fans, extends $\chi$ to $\Omega(\lambda)$ in $T$ (randomized) time such that for some $T_{0}=O(m)$, we have,
 $\mathbb{E}\left[T\right]\leq T_{0}\quad\text{and for all $\delta>0$}\quad\Pr\left[T\geq(1+\delta)\cdot 2T_{0}\right]\leq\exp\left(-\frac{\delta\cdot\lambda}{100}\right).$
-Proof Sketch.
-The runtime for coloring each single edge in Lemma 4.2 is $O(m/\lambda)$ by Lemma 6.7 assuming the coloring succeeds, which happens with probability at least $1/4$ by Lemma 6.5.
+Proof.
+The amortized runtime for coloring each single edge in Lemma 6.3 is $O(m/\lambda)$ by Lemma 8.7 assuming the coloring succeeds, which happens with probability at least $1/4$ by Lemma 8.5.
 We emphasize that this is in an average sense: the algorithm Color-U-Fans first uses Prime-U-Fans to prime $\Omega(\lambda/\Delta)$ u-fans in expected $O(m/\lambda)$ time per u-fan and thus $O((\lambda/\Delta)\cdot(m/\lambda))=O(m/\Delta)$ expected total time, and then deterministically colors them in $O(m/\Delta)$ time using Activate-U-Fans.
@@
 using a loose upper bound in the last step. Given that the runtime of the algorithm is $t\cdot O(m/\lambda)$, the tail inequality follows. ∎
-We note Lemma A.3 guarantees that as long as $\lambda=\omega(\log{n})$, the $O(m)$ runtime of the algorithm also holds with high probability.
+We note Lemma 10.1 guarantees that as long as $\lambda=\omega(\log{n})$, the $O(m)$ runtime of the algorithm also holds with high probability.
 For smaller values of $\lambda$ (which happens only as a corner case in the algorithm [^16] The only case in our algorithm where this tail bound cannot be replaced with a high probability bound is when $m/\Delta=o(\log{n})$ and at the same time $\Delta\log^{2}\Delta=\omega(m\log{\Delta})$, which means $m=o(n\log{n})$ and yet $\Delta=\omega(n/\log{n})$.), we need the specific tail inequality proven in the lemma instead.
-Step 3: In the next step, we provide a more fine-grained version of Lemma 4.3 that will be used in the last step of the argument to (somewhat) bypass the $O(\log{n})$ factor loss of the original lemma.
-Lemma A.4 (A slight modification of Lemma 4.3).
-There exists an absolute constant $b>0$ such that the following holds.
-There is an algorithm given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges, and an integer $\lambda_{0}\leq\lambda$, extends $\chi$ to all but $\lambda_{0}$ uncolored edges in $T$ (randomized) time such that for some $T_{0}=O(m\cdot\log{(\lambda/\lambda_{0})}+\Delta\lambda)$, we have,
+Furthermore, we provide a similarly fine-grained version of Lemma 6.4 that will be used in the last step of the argument to bypass the $O(\log{n})$ factor loss of the original lemma.
+Lemma 10.2 (A slight modification of Lemma 6.4).
+There is an algorithm that given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges, and an integer $\lambda_{0}\leq\lambda$, extends $\chi$ to all but $\lambda_{0}$ uncolored edges in $T$ (randomized) time such that for some $T_{0}=O(m\cdot\log{(\lambda/\lambda_{0})}+\Delta\lambda)$, we have,
 $\mathbb{E}\left[T\right]\leq T_{0}\quad\text{and for all $\delta>0$}\quad\Pr\left[T\geq(1+\delta)\cdot 2T_{0}\right]\leq\exp\left(-\frac{\delta\cdot\lambda_{0}}{200}\right).$
 Proof.
-The algorithm is verbatim as in Lemma 4.3 by replacing Lemma 4.1 with Lemma A.2 and Lemma 4.2 with Lemma A.3: repeatedly color a constant fraction of remaining edges as long as $\lambda_{0}$ uncolored edges remain.
-Since we start with $\lambda$ uncolored edges and finish with $\lambda_{0}$ many, and reduce uncolored edges by a constant factor each time, we apply Lemmas A.2 and A.3 for a total of $O(\log{(\lambda/\lambda_{0})})$ times.
-Moreover, the $\Delta\lambda$ terms in the runtime of Lemma A.2 form a geometric series and thus their contribution to the runtime is $O(\Delta\lambda)$ in total.
+The algorithm is verbatim as in Lemma 6.4 by replacing Lemma 6.3 with Lemma 10.1: repeatedly color a constant fraction of remaining edges as long as $\lambda_{0}$ uncolored edges remain.
+Since we start with $\lambda$ uncolored edges and finish with $\lambda_{0}$ many, and reduce uncolored edges by a constant factor each time, we apply Lemma 6.2 and Lemma 10.1 for a total of $O(\log{(\lambda/\lambda_{0})})$ times.
+Moreover, the $\Delta\lambda$ terms in the runtime of Lemma 6.2 form a geometric series and thus their contribution to the runtime is $O(\Delta\lambda)$ in total.
 This proves the expected bound on the runtime, i.e., $\mathbb{E}\left[T\right]\leq T_{0}=O(m\cdot\log{(\lambda/\lambda_{0})}+\Delta\lambda)$.
-As for the tail bound, for every application of Lemma A.3 with an intermediate value $\lambda^{\prime}\in[\lambda_{0},\lambda]$, and $T^{\prime}_{0}=O(m)$ being the $T_{0}$-parameter of Lemma A.3, we have,
+As for the tail bound, in each application of Lemma 10.1 with an intermediate value $\lambda^{\prime}\in[\lambda_{0},\lambda]$, and $T^{\prime}_{0}=O(m)$ being the $T_{0}$-parameter of Lemma 10.1, we have,
 […]
 Thus, by union bound,
-Pr[total $\Pr\left[\text{total runtime}\geq k\cdot T_{0}\right]$ $\leq\sum_{\lambda^{\prime}}\exp\left(-\frac{\delta\cdot\lambda^{\prime}}{100}\right)$
+$\Pr\left[\text{total runtime}\geq k\cdot T_{0}\right]$ $\leq\sum_{\lambda^{\prime}}\exp\left(-\frac{\delta\cdot\lambda^{\prime}}{100}\right)$
 $\leq\sum_{i=0}^{\infty}\cdot\exp\left(-\frac{\delta\cdot\lambda_{0}\cdot\gamma^{i}}{100}\right)$ (as number of uncolored edges drops by some constant factor, say, $\gamma=1+\Theta(1)$ each time)
@@
 by (loosely) upper bounding the sum of the geometric series using its first term. ∎
-Step 4: Finally, we show how to put these lemmas together to $(\Delta+1)$ edge color the graph in $O(m\log{\Delta})$ randomized time, and conclude the proof of Theorem A.1.
-Before getting to the main argument, we need the following specialized concentration inequality that we prove here (although, its formulation appears general enough to be of use elsewhere also).
-The proof follows a standard moment generating function argument.
-Proposition A.5.
+Before moving on from this subsection, we mention the following specialized concentration inequality that we need in order to be able to exploit the tail bounds proven in Lemma 10.1 and Lemma 10.2.
+The proof follows a standard moment generating function argument and is postponed to Appendix D.
+Proposition 10.3.
 Let $\{X_{i}\}_{i=1}^{n}$ be $n$ independent non-negative random variables associated with parameters $\{\alpha_{i}\}_{i=1}^{n}$ and $\{\beta_{i}\}_{i=1}^{n}$ such that for each $i\in[n]$, $\alpha_{i},\beta_{i}\geq 1$, and for every $\delta>0$,
@@
 $\Pr\left[\sum_{i=1}^{n}X_{i}\geq\sum_{i=1}^{n}\alpha_{i}+t\right]\leq\exp\left(-\left(\min_{i=1}^{n}\frac{\beta_{i}}{2\alpha_{i}}\right)\cdot\left(t-2\sum_{i=1}^{n}\frac{\alpha_{i}}{\beta_{i}}\right)\right).$
-Proof.
-Let $s:=1/2\cdot\min_{i}\beta_{i}/\alpha_{i}$ and for $i\in[n]$, $Y_{i}:=\max(X_{i}-\alpha_{i},0)$. Letting $Y:=\sum_{i=1}^{n}Y_{i}$, leads
-$\Pr\left[\sum_{i=1}^{n}X_{i}\geq\sum_{i=1}^{n}\alpha_{i}+t\right]\leq\Pr\left[Y\geq t\right],$
-for all $t\geq 0$. We further have,
-$\Pr\left[Y\geq t\right]=\Pr\left[e^{s\cdot Y}\geq e^{s\cdot t}\right]\leq e^{-s\cdot t}\cdot\mathbb{E}\left[e^{s\cdot Y}\right]=e^{-s\cdot t}\cdot\mathbb{E}\left[\prod_{i=1}^{n}e^{s\cdot Y_{i}}\right]=e^{-s\cdot t}\cdot\prod_{i=1}^{n}\mathbb{E}\left[e^{s\cdot Y_{i}}\right]$ (2)
-where the inequality is by Markov inequality and the final equality is by the independence of the variables. Now, for every $i\in[n]$,
-[…]
-[…]
-[…]
-$=\frac{1}{\beta_{i}\cdot(1-s\cdot({\alpha_{i}}/{\beta_{i}}))}$ (since the integral of $e^{-Kx}=-e^{-Kx}/K$ and $s<\beta_{i}/\alpha_{i}$ thus the value of this function in $\infty$ is zero)
-$\leq\frac{1}{1-s\cdot(\alpha_{i}/\beta_{i})}$ (as $\beta_{i}\geq 1$)
-$\leq\exp\left(2s\cdot\alpha_{i}/\beta_{i}\right).$ (as $s\cdot\alpha_{i}/\beta_{i}<1/2$ and $1-x\geq e^{-2x}$ for $x<1/2$)
-Plugging this bounds in Equation 2, implies that
-$\Pr\left[Y\geq t\right]\leq\exp\left(-s\cdot t+\sum_{i=1}^{n}2s\cdot\frac{\alpha_{i}}{\beta_{i}}\right)=\exp\left(-\left(\min_{i=1}^{n}\frac{\beta_{i}}{2\alpha_{i}}\right)\cdot\left(t-2\sum_{i=1}^{n}\frac{\alpha_{i}}{\beta_{i}}\right)\right),$
-by the choice of $s$, concluding the proof. ∎
-We now get to the main parts of the argument. The first part is based on a modification to the Eulerian partition approach used in the proof of Theorem 1.1.
-Lemma A.6.
-There is an algorithm that given a graph $G$, finds a $(\Delta+1)$ edge coloring of all but $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability.
+10.2 Proof of Theorem 1.1
+We are now ready to use the more fine-grained versions of our main technical lemmas to $(\Delta+1)$ edge color the graph in $O(m\log{n})$ randomized time, and conclude the proof of Theorem 1.1.
+The first part is based on a modification to the Eulerian partition approach used in the proof of Theorem 6.1.
+Notice that the runtime of this algorithm is even $O(m\log{\Delta})$ and not $O(m\log{n})$ (the distinction at this point is irrelevant for us in proving Theorem 1.1, but we state the result this way so we can use it later in Appendix C as well).
+Lemma 10.4.
+There is an algorithm that given a graph $G$, finds a $(\Delta+1)$ edge coloring of all but (exactly) $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability.
 Proof.
@@
 Combining $\chi_{1}$ and $\chi_{2}$ and uncoloring the two smallest color classes, gives a $(\Delta+1)$ coloring $\chi$ of all but $\lambda:=O(m/\Delta)$ edges.
-We then run Lemma A.4 to reduce the number of uncolored edges to $\lambda_{0}:=m/\Delta$ edges.
-The correctness of the algorithm follows immediately.
-For the runtime, we have $k:=O(\Delta)$ sub-problems in total, with $2^{d}$ subproblems on $m/2^{d}$ edges and maximum degree $\Delta/2^{d}$ for $d\leq\left\lceil\log{\Delta}\right\rceil$.
+We then run Lemma 10.2 to reduce the number of uncolored edges to $\lambda_{0}:=m/\Delta$ edges.
+The correctness of the algorithm thus follows immediately.
+For the runtime, we have $k:=O(\Delta)$ sub-problems in total, with $2^{d}$ subproblems on $\left\lceil m/2^{d}\right\rceil$ edges and maximum degree $\left\lceil\Delta/2^{d}\right\rceil$ for $d\leq\left\lceil\log{\Delta}\right\rceil$.
 Let $X_{1},\ldots,X_{k}$ denote the runtime of these subproblems and thus $X:=\sum_{i=1}^{k}X_{i}$ is the total runtime.
 We have,
-$\mathbb{E}\left[X\right]=\sum_{i}\mathbb{E}\left[X_{i}\right]=\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot O\left(\frac{m}{2^{d}}\log{(\lambda/\lambda_{0})}+\frac{\Delta}{2^{d}}\cdot\lambda\right)=O(m\log{\Delta})$
-by Lemma A.4, since $\lambda=O(m/\Delta)$, $\lambda_{0}=m/\Delta$, and thus $\log{(\lambda/\lambda_{0})}=O(1)$ and $\Delta\cdot\lambda=O(m)$.
-Moreover, by Lemma A.4, for the sub-problem corresponding to $X_{i}$ at some level $d$ of the recursion, $T^{d}_{0}=O(m/2^{d})$, and every $\delta\geq 0$,
+$\mathbb{E}\left[X\right]=\sum_{i}\mathbb{E}\left[X_{i}\right]=\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot O\left(\frac{m}{2^{d}}\cdot\log{(\lambda/\lambda_{0})}+\frac{\Delta}{2^{d}}\cdot\lambda\right)=O(m\log{\Delta})$
+by Lemma 10.2, since $\lambda=O(m/\Delta)$, $\lambda_{0}=m/\Delta$, and thus $\log{(\lambda/\lambda_{0})}=O(1)$ and $\Delta\cdot\lambda=O(m)$.
+Moreover, by Lemma 10.2, for the sub-problem corresponding to $X_{i}$ at some level $d$ of the recursion, $T^{d}_{0}=O(m/2^{d})$, and every $\delta\geq 0$,
 $\Pr\left[X_{i}\geq(1+\delta)\cdot 2T^{d}_{0}\right]\leq\exp\left(-\delta\cdot\frac{\lambda_{0}}{200}\right)=\exp\left(-\delta\cdot\frac{m}{200\Delta}\right).$
-For this $X_{i}$, define $\alpha_{i}:=2T^{d}_{0}$ and $\beta_{i}:=m/200\Delta$.
-We can now apply Proposition A.5 and for $t:=1000\sum_{i=1}^{k}\alpha_{i}=1000\cdot\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot 2T^{d}_{0}=O(m\log{\Delta})$, have,
+For this $X_{i}$, define $\alpha_{i}:=2T^{d}_{0}$ and $\beta_{i}:=m/200\Delta$. We can now apply Proposition 10.3 and for
+$t:=1000\sum_{i=1}^{k}\alpha_{i}=1000\cdot\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot 2T^{d}_{0}=O(m\log{\Delta}),$
+have,
 $\Pr\left[X\geq t\right]$ $\leq\exp\left(-\left(\min_{i=1}^{n}\frac{\beta_{i}}{2\alpha_{i}}\right)\cdot\left(t-2\sum_{i=1}^{n}\frac{\alpha_{i}}{\beta_{i}}\right)\right)$
@@
 $=\exp\left(-\Theta(1)\cdot\frac{m\log{\Delta}}{\Delta}\right),$
-where we can additionally make the constant in $\Theta(1)$ arbitrarily large without changing the asymptotic runtime of the algorithm.
+where we can make the constant in $\Theta(1)$ arbitrarily large without changing the asymptotic runtime of the algorithm.
 This probability is always at most $1/\operatorname*{poly}(n)$ because either $\Delta<\sqrt{m}$, in which case, the probability is at most $1/2^{\Theta(\sqrt{m})}$, or $\Delta>\sqrt{m}$, and thus $\log{\Delta}=\Theta(\log{n})$ and the probability is $1/\operatorname*{poly}(n)$ at the very least.
@@
 ∎
-In the next part, we design an algorithm that color the remaining $m/\Delta$ uncolored edges left by Lemma A.6.
-This can be easily done in $O(m\log{n})$ time using Lemma A.4, but doing this in $O(m\log{\Delta})$ time instead requires one more ingredient.
+To conclude the proof, we have the following lemma that colors the remaining $m/\Delta$ edges left uncolored by the algorithm of Lemma 10.4.
+The proof is a straightforward application of Lemma 10.2, plus the original Vizing’s Fan and Chain approach stated in Lemma 4.2.
+Lemma 10.5.
+There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ with $\lambda=m/\Delta$ uncolored edges, extends the coloring to all edges in $O(m\log{n})$ time with high probability.
+Proof.
+We first run our Lemma 10.2 with the given parameters $\lambda=m/\Delta$ and $\lambda_{0}=100\log{n}$.
+By Lemma 10.2, the expected runtime will be some $T_{0}=O(m\log{(\lambda/\lambda_{0})}+\Delta\lambda)=O(m\log{n})$ since $\lambda\leq m$ and $\lambda_{0}\geq 1$.
+Moreover, by the same lemma,
+$\Pr\left[\text{runtime of the algorithm}\geq 2000\cdot T_{0}\right]\leq\exp\left(-10\lambda_{0}\right)\leq\exp\left(-1000\log{n}\right)=1/\operatorname*{poly}(n),$
+since $\lambda_{0}\geq 100\log{n}$.
+Thus, after running this part, in $O(m\log{n})$ time, with high probability, we will be left with only $O(\log{n})$ uncolored edges.
+We can then color each remaining uncolored edge in $O(n)$ time using the original Vizing’s Fans and Chains approach of Lemma 4.2, thus obtaining a $(\Delta+1)$ coloring of the entire graph in $O(m\log{n})$ time with high probability.
+∎
+Theorem 1.1 now follows immediately from Lemma 10.4 and Lemma 10.5.
+Acknowledgements
+Part of this work was conducted while Sepehr Assadi and Soheil Behnezhad were visiting the Simons Institute for the Theory of Computing as part of the Sublinear Algorithms program.
+Sepehr Assadi is supported in part by a Sloan Research Fellowship, an NSERC Discovery Grant (RGPIN-2024-04290), and a Faculty of Math Research Chair grant from University of Waterloo.
+Soheil Behnezhad is funded by an NSF CAREER award CCF-2442812 and a Google Faculty Research Award.
+Martín Costa is supported by a Google PhD Fellowship.
+Shay Solomon is funded by the European Union (ERC, DynOpt, 101043159).
+Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council.
+Neither the European Union nor the granting authority can be held responsible for them.
+Shay Solomon is also funded by a grant from the United States-Israel Binational Science Foundation (BSF), Jerusalem, Israel, and the United States National Science Foundation (NSF).
+Tianyi Zhang is supported by funding from the starting grant “A New Paradigm for Flow and Cut Algorithms” (no.
+TMSGI2_218022) of the Swiss National Science Foundation.
+Appendix A Vizing’s Theorem for Multigraphs in Near-Linear Time
+We now show how to extend our arguments to multigraphs.
+A generalization of Vizing’s theorem to multigraphs shows that any multigraph $G$ with maximum degree $\Delta$ and maximum multiplicity $\mu$ can be edge colored with $\Delta+\mu$ colors [Viz64] (and this is worst-case optimal for $\mu\leq\Delta/2$ in the sense that not every multigraph admits an edge coloring with fewer colors).
+We prove the following theorem, which shows how to compute such a coloring in near-linear time.
+Theorem A.1.
+There is a randomized algorithm that, given an undirected multigraph $G=(V,E)$ on $n$ vertices and $m$ edges with maximum degree $\Delta$ and maximum multiplicity $\mu$, finds a $(\Delta+\mu)$-edge coloring of $G$ in $O(m\log{n})$ time with high probability.
+Similar to our results for $(\Delta+1)$ edge coloring, we first focus on only obtaining an $\tilde{O}(m)$ time algorithm, and then use almost exactly the same argument we used to extend Theorem 6.1 to Theorem 1.1, to optimize this runtime to $O(m\log{n})$ time.
+Notation for Multigraphs: We refer to edges in $G$ with the same endpoints as parallel.
+Whenever we refer to an edge $e$ of the multigraph $G$, we are referring to a specific edge and say that this edge is distinct from its parallel edges.
+A.1 Vizing Fans and Separable Collection in Multigraphs
+Let $G=(V,E)$ be an undirected multigraph on $n$ vertices and $m$ edges with maximum degree $\Delta$ and maximum multiplicity $\mu$, and $\chi$ be a partial $(\Delta+\mu)$-edge coloring of $G$.
+We now describe how to generalize the objects used by our algorithm for simple graphs to deal with multigraphs.
+We begin by defining a generalization of Vizing fans for multigraphs [Viz64].
+For convenience, we still refer to them as Vizing fans.
+Definition A.2 (Vizing fan for multigraphs).
+A Vizing fan for a multigraph is a sequence $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},e_{1},C_{1}),\dots,(v_{k},e_{k},C_{k})$ where $u,v_{1},\dots,v_{k}$ are distinct vertices, $C_{1},\dots,C_{k}\subseteq[\Delta+\mu]$ are subsets of colors and $e_{1},\dots,e_{k}$ are edges such that
+  1. 1.
+    $\alpha\in\mathsf{miss}_{\chi}(u)$ and $C_{i}\subseteq\mathsf{miss}_{\chi}(v_{i})$ of size $\mu$ for all $i\in[k]$.
+  2. 2.
+    $v_{1},\dots,v_{k}$ are distinct neighbours of $u$ and $e_{i}$ is an edge with endpoints $u$ and $v_{i}$ for all $i\in[k]$.
+  3. 3.
+    $\chi(e_{1})=\bot$ and, for all $i>1$, there exists $p(i)\in[i-1]$ such that $\chi(e_{i})\in C_{p(i)}$.
+  4. 4.
+    The sets $C_{1},\dots,C_{k-1},\mathsf{miss}_{\chi}(u)$ are mutually disjoint.
+  5. 5.
+    Either $C_{k}\cap\mathsf{miss}_{\chi}(u)\neq\varnothing$ or $C_{k}\cap\left(C_{1}\cup\dots\cup C_{k-1}\right)\neq\varnothing$.
+We say that the Vizing fan $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},e_{1},C_{1}),\dots,(v_{k},e_{k},C_{k})$ is $\alpha$-primed, has center $u$ and leaves $v_{1},\dots,v_{k}$.
+We refer to $C_{i}$ as the colors of $v_{i}$ within $\operatorname*{\mathbf{F}}$.
+A crucial property is that we can rotate colors around the Vizing fan $\operatorname*{\mathbf{F}}$: given any $i\in[k]$, there is a sequence $1=i_{1}<\dots<i_{\ell}=i$ (where $i_{\ell-j}=p^{(j)}(i)$) such that we can set $\chi(e_{i_{1}})\leftarrow\chi(e_{i_{2}}),\dots,\chi(e_{i_{\ell-1}})\leftarrow\chi(e_{i_{\ell}}),\chi(e_{i_{\ell}})\leftarrow\bot$.
+We say that $\operatorname*{\mathbf{F}}$ is a trivial Vizing fan if $C_{k}\cap\mathsf{miss}_{\chi}(u)\neq\varnothing$.
+Note that, if $\operatorname*{\mathbf{F}}$ is trivial, we can immediately extend the coloring $\chi$ to $(e_{1})$ by rotating colors around $\operatorname*{\mathbf{F}}$ to leave $(e_{1})$ uncolored and setting $\chi(e_{k})$ to be any color in $C_{k}\cap\mathsf{miss}_{\chi}(u)$.
+For completeness, we prove the following lemma, which shows how to efficiently construct such a Vizing fan.
+Lemma A.3 ([Viz64]).
+Given an uncolored edge $e=(u,v)$ and a color $\alpha\in\mathsf{miss}_{\chi}(u)$, we can construct an $\alpha$-primed Vizing fan $\operatorname*{\mathbf{F}}$ with center $u$ in $O(\Delta)$ time.
+Proof.
+Let $v_{1}=v$, $e_{1}=e$ and $C_{1}\subseteq\mathsf{miss}_{\chi}(v)$ be a subset of size $\mu$.
+If $C_{1}\cap\mathsf{miss}_{\chi}(u)\neq\varnothing$, then we can return $(u,\alpha),(v_{1},e_{1},C_{1})$ as a Vizing fan.
+Otherwise, we assume inductively that we have constructed a sequence $(u,\alpha),(v_{1},e_{1},C_{1}),\dots,(v_{i},e_{i},C_{i})$ satisfying Conditions 1 to 4 of Definition A.2.
+In this case, we have that $|C_{1}\cup\dots\cup C_{i}|=i\mu$.
+Since there are at most $i\mu-1$ many colored edges that have $u$ as one endpoint and one of $v_{1},\dots,v_{i}$ as the other, it follows that there must be some edge $e_{i+1}=(u,v_{i+1})$ such that $\chi(e_{i+1})\in C_{1}\cup\dots\cup C_{i}$ and $v_{i+1}\notin\{v_{1},\dots,v_{i}\}$.
+Let $C_{i+1}\subseteq\mathsf{miss}_{\chi}(v_{i+1})$ be a subset of size $\mu$.
+If either $C_{i+1}\cap\mathsf{miss}_{\chi}(u)\neq\varnothing$ or $C_{i+1}\cap\left(C_{1}\cup\dots\cup C_{i}\right)\neq\varnothing$, then $(u,\alpha),(v_{1},e_{1},C_{1}),\dots,(v_{i+1},e_{i+1},C_{i+1})$ is a Vizing fan and we are done.
+Otherwise, we continue the induction on this longer sequence.
+Since this process must terminate within at most $\Delta/\mu$ steps, this process always returns a Vizing fan.
+Furthermore, using standard data structures, we can implement this process in $O(\Delta)$ time.
+∎
+Vizing Chains in Multigraphs: Let $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},e_{1},C_{1}),\dots,(v_{k},e_{k},C_{k})$ be a non-trivial $\alpha$-primed Vizing fan, let $c$ be a color in $C_{k}\cap(C_{1},\dots,C_{k-1})$ and let $P$ denote the maximal $\{\alpha,c\}$-alternating path starting at $u$.
+Then, by a similar argument to the case of simple graphs, we can extend the coloring $\chi$ to the edge $e_{1}$ by flipping the path $P$ and rotating colors around $\operatorname*{\mathbf{F}}$.
+Furthermore, this can be done in time $O(\Delta+|P|)$.
+We denote a call to the algorithm that extends the coloring in the manner by $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$.
+Separable Collections in Multigraphs: We define a separable collection of u-components $\operatorname*{\mathcal{U}}$ in a multigraph in the exact same way as simple graphs.
+We emphasise that, given a separable collection $\operatorname*{\mathcal{U}}$, any distinct u-components $\operatorname*{\boldsymbol{g}}_{1},\operatorname*{\boldsymbol{g}}_{2}\in\operatorname*{\mathcal{U}}$ are edge-disjoint but may contain parallel edges.
+In other words, it is possible for two distinct parallel uncolored edges $e_{1}$ and $e_{2}$ to be contained in u-components within the separable collection $\operatorname*{\mathcal{U}}$.
+We define $\operatorname*{\mathcal{U}}$-avoiding Vizing fans for multigraphs in the following analogous way.
+Definition A.4.
+Let $\operatorname*{\mathcal{U}}$ be a separable collection and $\operatorname*{\mathbf{F}}=(u,\alpha),(v_{1},e_{1},C_{1}),\dots,(v_{k},e_{k},C_{k})$ be a Vizing fan.
+We say that the Vizing fan $\operatorname*{\mathbf{F}}$ is $\operatorname*{\mathcal{U}}$-avoiding if $C_{i}\subseteq\mathsf{miss}_{\chi}(v_{i})\setminus C_{\operatorname*{\mathcal{U}}}(v_{i})$ for each $v_{i}\in\operatorname*{\mathbf{F}}$.
+Using these new definitions, we can directly extend Lemmas 5.8 and 5.9 to multigraphs by slightly changing the proofs, giving us the following lemmas.
+Lemma A.5.
+Given a u-edge $\operatorname*{\boldsymbol{e}}\in\operatorname*{\mathcal{U}}$, there exists a $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}$ of $\operatorname*{\boldsymbol{e}}$.
+Furthermore, we can compute such a Vizing fan in $O(\Delta)$ time.
+[^17] Recall that we say that $\operatorname*{\mathbf{F}}$ is a Vizing fan of $\operatorname*{\boldsymbol{e}}=(u,v,\alpha)$ if $\operatorname*{\mathbf{F}}$ is $\alpha$-primed, has center $u$, and its first leaf is $v$.
+Lemma A.6.
+Let $\chi$ be a $(\Delta+\mu)$-edge coloring of a graph $G$ and $\operatorname*{\mathcal{U}}$ be a separable collection.
+For any u-edge $\operatorname*{\boldsymbol{e}}=(u,v,\alpha)\in\operatorname*{\mathcal{U}}$ with a $\operatorname*{\mathcal{U}}$-avoiding Vizing fan $\operatorname*{\mathbf{F}}$, we have the following:
+  1. 1.
+    Rotating colors around $\operatorname*{\mathbf{F}}$ does not damage any u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}\}$.
+  2. 2.
+    Calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ damages at most one u-component in $\operatorname*{\mathcal{U}}\setminus\{\operatorname*{\boldsymbol{e}}\}$.
+Furthermore, we can identify the damaged u-component in $O(1)$ time.
+A.2 Proof of Theorem A.1
+Our main algorithm for multigraphs is completely analogous to our main algorithm for simple graphs, which we describe in Section 6.
+In particular, the algorithm also consists of two main components that are combined in the same way, which we summarize in the following lemmas.
+Lemma A.7 (A modification of Lemma 6.2 for multigraphs).
+There is an algorithm that, given a multigraph $G$, a partial $(\Delta+\mu)$-edge coloring $\chi$ of $G$ and a set of $\lambda$ uncolored edges $U$, does one of the following in $O(m+\Delta\lambda)$ time:
+  1. 1.
+    Extends the coloring to $\Omega(\lambda)$ uncolored edges.
+  2. 2.
+    Modifies $\chi$ to obtain a separable collection of $\Omega(\lambda)$ u-fans $\mathcal{U}$.
+The proof of this lemma is similar to the proof of Lemma 6.2 but with some minor changes. In Section A.3, we sketch how to modify the proof of Lemma 6.2 to extend it to Lemma A.7.
+Lemma A.8 (A modification of Lemma 6.3 for multigraphs).
+There is an algorithm that, given a multigraph $G$, a partial $(\Delta+\mu)$-edge coloring $\chi$ of $G$ and a separable collection of $\lambda$ u-fans $\mathcal{U}$, extends $\chi$ to $\Omega(\lambda)$ edges in $O(m\log n)$ time w.h.p.
+The proof of this lemma is almost identical to the proof of Lemma 6.3, thus we omit the details.
+Using Lemmas A.7 and A.8, we get the following lemma which allows us to extend an edge coloring $\chi$ to a set of uncolored edges.
+Lemma A.9 (A modification of Lemma 6.4 for multigraphs).
+There is an algorithm that, given a multigraph $G$ and a partial $(\Delta+\mu)$-edge coloring $\chi$ of $G$ with $\lambda$ uncolored edges $U$, extends $\chi$ to the remaining uncolored edges in time $O((m+\Delta\lambda)\log^{2}n)$ w.h.p.
+The proof of this lemma is verbatim the same as the proof of Lemma 6.4, and hence is omitted.
+Finally, in the same way as in the proof of Theorem 1.1, we can apply standard Euler partitioning using Lemma A.9 to merge the colorings.
+We only note that in multigraphs, splitting an $m$-edge multigraph $G$ via the Euler partition leads to two multigraphs, each with $\left\lceil m/2\right\rceil$ edges, maximum degree $\left\lceil\Delta/2\right\rceil$, and maximum multiplicity $\left\lceil\mu/2\right\rceil$.
+Thus, the number of colors used to color the graph recursively will be $2\cdot(\left\lceil\Delta/2\right\rceil+\left\lceil\mu/2\right\rceil)=\Delta+\mu+O(1)$ colors.
+Hence, we again only need to uncolor $O(m/\Delta)$ edges to obtain a $(\Delta+\mu)$ coloring and then extend this coloring to the remaining $O(m/\Delta)$ edges using Lemma A.9.
+This leads to an $O(m\log^{3}{n})$ time algorithm for finding a $(\Delta+\mu)$ edge coloring with high probability.
+We can also optimize this algorithm to run in $O(m\log{n})$ time with high probability exactly as in our $(\Delta+1)$ edge coloring algorithm.
+Specifically, using exactly the same argument as in Section 10, we have the following analogues of Lemma 10.4 and Lemma 10.5 for multigraphs.
+Lemma A.10 (A modification of Lemma 10.4 for multigraphs).
+There is an algorithm that given a multigraph $G$, finds a $(\Delta+\mu)$ edge coloring of all but (exactly) $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability.
+Lemma A.11 (A modification of Lemma 10.5 for multigraphs).
+There is an algorithm that, given a multigraph $G$, a partial $(\Delta+\mu)$-edge coloring $\chi$ with $\lambda=m/\Delta$ uncolored edges, extends the coloring to all edges in $O(m\log{n})$ time with high probability.
+Theorem A.1 now follows immediately from Lemma A.10 and Lemma A.11.
+A.3 Proof Sketch of Lemma A.7
+The algorithm for Lemma A.7 is almost identical to the algorithm for Lemma 6.2, except that we replace the relevant definitions with their generalizations for multigraphs, leading to a few extra cases.
+The overall structure of the algorithm is still the same as Algorithm 4, but using modifications of Prune-Vizing-Fans and Reduce-U-Edges for multigraphs.
+Initializing the Seperable Collection $\operatorname*{\mathcal{U}}$: Given a set of $\lambda$ uncolored edges $U$, we can construct a separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$ in the exact same way as simple graphs.
+However, there may exist distinct u-edges $\operatorname*{\boldsymbol{e}},\operatorname*{\boldsymbol{e}}^{\prime}\in\operatorname*{\mathcal{U}}$ that correspond to parallel (but distinct) uncolored edges.
+Modifying Prune-Vizing-Fans for Multigraphs: Suppose that we want to run Prune-Vizing-Fans on a separable collection $\operatorname*{\mathcal{U}}$ with color $\alpha$.
+In this call to the subroutine, we consider all of the $\alpha$-primed u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})\subseteq\operatorname*{\mathcal{U}}$.
+We implement this subroutine for multigraphs in the same way as for simple graphs, but with the following additional preprocessing step to ensure that none of the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ are parallel.
+Suppose that we have distinct $\operatorname*{\boldsymbol{e}},\operatorname*{\boldsymbol{e}}^{\prime}\in\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ that correspond to parallel uncolored edges with endpoints $u$ and $v$.
+Then, by the definition of a separable collection, one of these u-edges must have $u$ as a center and the other must have $v$ as a center.
+Thus, $\alpha\in\mathsf{miss}_{\chi}(u)\cap\mathsf{miss}_{\chi}(v)$, so we can extend the coloring to one of these u-edges by coloring it with $\alpha$ and then remove them both from $\operatorname*{\mathcal{U}}$.
+We can scan through all of the $\alpha$-primed u-edges and perform this operation whenever we encounter parallel u-edges.
+This maintains the invariant that $\operatorname*{\mathcal{U}}$ is a separable collection and ensures that none of the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ are parallel.
+After performing this preprocessing step, we can implement the rest of Prune-Vizing-Fans in the exact same way as described in Section 7.1, except that we must use the generalization of Vizing fans for multigraphs.
+In particular, all of the lemmas in Section 7.1 describing the properties of Prune-Vizing-Fans still hold.
+Modifying Reduce-U-Edges for Multigraphs: Suppose that we want to run Reduce-U-Edges on a separable collection $\operatorname*{\mathcal{U}}$ with color $\alpha$ and have obtained a collection of vertex-disjoint $\operatorname*{\mathcal{U}}$-avoiding Vizing fans $\mathcal{F}$ corresponding to the u-edges in $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ from calling Prune-Vizing-Fans.
+We implement the subroutine Reduce-U-Edges in the same way as described in Section 7.2 but with the following modification to Algorithm 5 to account for an additional case.
+Suppose that we call $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ and take one more step along the path $\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})$ and observe the edge $e=(x,y)$.
+If the edge $e$ has already been seen (i.e.
+is contained in $S$) then we proceed as normal.
+[^18] That is, this exact copy of the edge is in $S$, not just an edge parallel to $e$.
+If neither the edge $e$ nor any edge parallel to $e$ is contained in $S$, then we again proceed as normal.
+However, if we observe that $e$ is not already contained in $S$ but some edge $e^{\prime}$ that is parallel to $e$ is contained in $S$, then we do the following: Let $\operatorname*{\mathbf{F}}^{\prime}\in\mathcal{F}$ be the Vizing fan such that $e^{\prime}\in P_{\operatorname*{\mathbf{F}}^{\prime}}$, set $\chi(e)\leftarrow\bot$ and $\chi(e^{\prime})\leftarrow\bot$, call $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ and $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}}^{\prime})$, set $\chi(e)\leftarrow\alpha$, remove $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ from $\mathcal{F}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ from $\operatorname*{\mathcal{U}}$, and set $S\leftarrow S\setminus(P_{\operatorname*{\mathbf{F}}}\cup P_{\operatorname*{\mathbf{F}}^{\prime}})$; see Figure 10 for an illustration.
+We can verify that, if this does happen, the edges $e$ and $e^{\prime}$ must appear in different orientations in the paths $P_{\operatorname*{\mathbf{F}}}$ and $P_{\operatorname*{\mathbf{F}}^{\prime}}$.
+Consequently, this operation removes 2 u-edges from $\operatorname*{\mathcal{U}}$ and extends the coloring to one more edge.
+[…]
+Figure 10: If $P_{\operatorname*{\mathbf{F}}}$ and $P_{\operatorname*{\mathbf{F}}^{\prime}}$ meet in the middle at a pair of parallel edges $e,e^{\prime}$, then we can shift the uncolored edge from $\operatorname*{\mathbf{F}}^{\prime}$ to $e^{\prime}$ and apply $\textnormal{{Vizing}}(F)$ which assigns $\chi(e)\leftarrow\alpha$; here $\alpha$ is blue.
+After modifying Algorithm 5 with this additional case, we strengthen 7.5 so that no two edges contained in distinct prefix paths in $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$ are parallel.
+This is sufficient to ensure that the analysis of Reduce-U-Edges in Section 7.2 extends to multigraphs.
+In particular, all of the lemmas in Section 7.2 describing the properties of Reduce-U-Edges still hold.
+Appendix B Shannon’s Theorem for Multigraphs in Near-Linear Time
+A classical theorem by Shannon shows that any multigraph $G$ with maximum degree $\Delta$ can be edge colored with $\left\lfloor 3\Delta/2\right\rfloor$ colors [Sha49].
+Unlike Vizing’s theorem for multigraphs, this bound does not depend on the maximum multiplicity of the graph.
+We show that we can also extend our approach to a near-linear time randomized algorithm for computing such a coloring.
+Previously, only the original $O(mn)$ time algorithm of [Sha49] and a recent $O(n\cdot\Delta^{18})$ time algorithm of [Dha24a] were known for this problem.
+Theorem B.1.
+There is a randomized algorithm that, given an undirected multigraph graph $G=(V,E)$ on $n$ vertices and $m$ edges with maximum degree $\Delta$, finds a $\left\lfloor 3\Delta/2\right\rfloor$-edge coloring of $G$ in $O(m\log{n})$ time with high probability.
+Yet again, we first focus on obtaining a near-linear time algorithm, and postpone optimizing the runtime to $O(m\log{n})$ to the end of the proof.
+B.1 Shannon Fans
+Let $G=(V,E)$ be an undirected multigraph on $n$ vertices and $m$ edges with maximum degree $\Delta$ and $\chi$ be a partial $\left\lfloor 3\Delta/2\right\rfloor$-edge coloring of $G$.
+We begin by defining the notion of a Shannon fan, which is conceptually very similar to a u-fan.
+Definition B.2 (Shannon fan).
+A Shannon fan is a tuple $\operatorname*{\boldsymbol{S}}=(u,v,w,\alpha,\beta,\gamma)$ where $u$, $v$ and $w$ are distinct vertices and $\alpha$, $\beta$ and $\gamma$ are distinct colors such that:
+  1. 1.
+    $(u,v)$ is an uncolored edge and $(u,w)$ is an edge with color $\gamma$.
+  2. 2.
+    $\alpha\in\mathsf{miss}_{\chi}(u)$, $\beta\in\mathsf{miss}_{\chi}(v)\cap\mathsf{miss}_{\chi}(w)$ and $\gamma\in\mathsf{miss}_{\chi}(v)$.
+We say that $u$ is the center of $\operatorname*{\boldsymbol{S}}$ and that $v$ and $w$ are the leaves of $\operatorname*{\boldsymbol{S}}$.
+We also say that the Shannon fan $\operatorname*{\boldsymbol{S}}$ is $\{\alpha,\beta\}$-primed, and that $\gamma$ is the auxiliary color.
+We say that $\operatorname*{\boldsymbol{S}}$ is the Shannon fan of the uncolored edge $(u,v)$.
+Activating Shannon Fans: Let $\operatorname*{\boldsymbol{S}}$ be an $\{\alpha,\beta\}$-primed Shannon fan with center $u$ and leaves $v$ and $w$.
+The key property of Shannon fans is that at most one of the $\{\alpha,\beta\}$-alternating paths starting at $v$ or $w$ ends at $u$.
+Suppose that the $\{\alpha,\beta\}$-alternating path starting at $v$ does not end at $u$.
+Then, after flipping this $\{\alpha,\beta\}$-alternating path, both $u,v$ are missing color $\alpha$.
+Thus, we can extend the coloring $\chi$ by assigning $\chi(u,v)\leftarrow\alpha$.
+If the path does not end a $w$ instead, we can set $\chi(u,w)\leftarrow\bot$ and $\chi(u,v)\leftarrow\gamma$ and carry out an analogous argument.
+We refer to this as activating the Shannon fan $\operatorname*{\boldsymbol{S}}$.
+Constructing Shannon Fans: The following lemma shows how to construct a Shannon fan.
+Lemma B.3.
+Given an uncolored edge $e$, we can either construct a Shannon fan $\operatorname*{\boldsymbol{S}}$ of the edge $e$ or extend $\chi$ to $e$ in $O(\Delta)$ time.
+Proof.
+Let $u$ and $v$ be the endpoints of $e$ and let $\gamma$ be any color in $\mathsf{miss}_{\chi}(v)$.
+If $\gamma\in\mathsf{miss}_{\chi}(u)$, then we can set $\chi(u,v)\leftarrow\gamma$ and we are done.
+Otherwise, let $(u,w)$ be the edge incident on $u$ with color $\gamma$ (note that $w\neq v$ since $\gamma\in\mathsf{miss}_{\chi}(v)$).
+Now, suppose that $\mathsf{miss}_{\chi}(u)$ is disjoint from both $\mathsf{miss}_{\chi}(v)$ and $\mathsf{miss}_{\chi}(w)$.
+Otherwise, we can extend $\chi$ to $e$ by shifting the color $\gamma$ between $(u,v)$ and $(u,w)$ and directly coloring the uncolored edge.
+The following claim shows that we can find a color in $\mathsf{miss}_{\chi}(v)\cap\mathsf{miss}_{\chi}(w)$.
+Claim B.4.
+We have that $\mathsf{miss}_{\chi}(v)\cap\mathsf{miss}_{\chi}(w)\neq\varnothing$.
+Proof.
+Suppose that $\mathsf{miss}_{\chi}(v)$ and $\mathsf{miss}_{\chi}(v)$ are disjoint. Then we have that
+$2\cdot\left(\left\lfloor\frac{3\Delta}{2}\right\rfloor-(\Delta-1)\right)+\left(\left\lfloor\frac{3\Delta}{2}\right\rfloor-\Delta\right)\geq|\mathsf{miss}_{\chi}(u)|+|\mathsf{miss}_{\chi}(v)|+|\mathsf{miss}_{\chi}(w)|\geq\left\lfloor\frac{3\Delta}{2}\right\rfloor,$
+which implies that $\left\lfloor 3\Delta/2\right\rfloor\geq(3\Delta/2)-1$, giving a contradiction. ∎
+Now, we can find colors $\alpha\in\mathsf{miss}_{\chi}(u)$ and $\beta\in\mathsf{miss}_{\chi}(v)\cap\mathsf{miss}_{\chi}(w)$ in $O(\Delta)$ time.
+It follows that $\operatorname*{\boldsymbol{S}}=(u,v,w,\alpha,\beta,\gamma)$ is a Shannon fan of the edge $e$.
+∎
+Pre-Shannon Fans: For notational convenience, we introduce the notion of a pre-Shannon fan.
+Definition B.5 (pre-Shannon fan).
+A pre-Shannon fan is a tuple $\operatorname*{\boldsymbol{S}}^{\star}=(u,v,w,\gamma)$ where $u$, $v$ and $w$ are distinct vertices and $\gamma$ is the auxiliary color such that $(u,v)$ is an uncolored edge, $(u,w)$ is an edge with color $\gamma$ and $\gamma\in\mathsf{miss}_{\chi}(v)$.
+By slightly modifying the proof of Lemma B.3, we get the following lemma which shows that we can either convert a pre-Shannon fan $\operatorname*{\boldsymbol{S}}^{\star}$ into a Shannon fan $\operatorname*{\boldsymbol{S}}$ or extend the coloring to the uncolored edge in $\operatorname*{\boldsymbol{S}}^{\star}$.
+Lemma B.6.
+Given a pre-Shannon fan $\operatorname*{\boldsymbol{S}}^{\star}=(u,v,w,\gamma)$, we can do one of the following in $O(\Delta)$ time:
+  1. 1.
+    Find colors $\alpha$ and $\beta$ such that $\operatorname*{\boldsymbol{S}}=(u,v,w,\alpha,\beta,\gamma)$ is a Shannon fan.
+  2. 2.
+    Extend the coloring to $(u,v)$ while only changing the color of $(u,w)$.
+B.2 Proof of Theorem B.1
+Our final algorithm for Shannon’s theorem is completely analogous to our final algorithm for simple graphs, which we describe in Section 6.
+In particular, the algorithm also consists of two main components that are combined in the same way, which we summarize in the following lemmas.
+Lemma B.7 (A modification of Lemma 6.2 for Shannon fans).
+There is an algorithm that, given a multigraph $G$, a partial $\left\lfloor 3\Delta/2\right\rfloor$-edge coloring $\chi$ and $\lambda=O(m/\Delta)$ uncolored edges that form a matching, does one of the following in $O(m)$ time deterministically:
+  1. 1.
+    Extends the coloring to $\Omega(\lambda)$ uncolored edges.
+  2. 2.
+    Returns a set of $\Omega(\lambda)$ vertex-disjoint Shannon fans of the uncolored edges that all share the same auxiliary color.
+We defer the proof of this lemma to Section B.3.
+Lemma B.8 (A modification of Lemma 6.3 for Shannon fans).
+There is an algorithm that, given a multigraph $G$, a partial $\left\lfloor 3\Delta/2\right\rfloor$-edge coloring $\chi$ and set of $\lambda$ vertex-disjoint Shannon fans $\mathcal{U}$ sharing the same auxiliary color, extends $\chi$ to $\Omega(\lambda)$ edges in $O(m\log n)$ time with high probability.
+The proof of this lemma is almost identical to the proof of Lemma 6.3, thus we omit the details.
+In particular, the set of vertex-disjoint Shannon fans $\operatorname*{\mathcal{U}}$ in this lemma has the same key properties as the separable collection of u-fans in Lemma 6.3, so the proof extends directly by simply replacing the u-fans with Shannon fans.
+We remark that the additional condition that the Shannon fans must share the same auxiliary color $\gamma$ is crucial.
+In particular, it ensures that we can prime these Shannon fans by flipping alternating paths which will not contain the color $\gamma$, and thus will not interfere with the auxiliary colors of the other Shannon fans in $\operatorname*{\mathcal{U}}$.
+Using Lemmas B.7 and B.8, we get the following lemma which allows us to extend an edge coloring $\chi$ to a matching of uncolored edges.
+Lemma B.9 (A modification of Lemma 6.4 for Shannon fans).
+There is an algorithm that, given a multigraph $G$, a partial $\left\lfloor 3\Delta/2\right\rfloor$-edge coloring $\chi$ and a set of $\lambda=O(m/\Delta)$ uncolored edges that form a matching, extends $\chi$ to these uncolored edges in $O(m\log^{2}n)$ time with high probability.
+The proof of this lemma is almost identical to the proof of Lemma 6.4, with the observation that repeatedly applying Lemmas B.7 and B.8 does not change the fact that the remaining uncolored edges form a matching.
+[^19] Note that this is not true when we repeatedly apply Lemmas 6.2 and 6.3 in our algorithm for simple graphs, since Lemma 6.2 changes the locations of the uncolored edges.
+Finally, in the same way as in the proof of Theorem A.1, we can apply standard Euler partitioning using Lemma B.9 to merge the colorings, applying it separately for each color class that we uncolor.
+This allows for obtaining an $O(m\log^{3}{n})$ time algorithm.
+The extension to an $O(m\log{n})$ time algorithm is verbatim as in Theorem A.1 and Theorem 1.1 and we only mention the relevant lemmas below without proving them.
+Lemma B.10 (A modification of Lemma 10.4 for Shannon’s Theorem).
+There is an algorithm that, given a multigraph $G$, finds a $\left\lfloor 3\Delta/2\right\rfloor$-edge coloring of all but (exactly) $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability.
+Lemma B.11 (A modification of Lemma 10.5 for Shannon’s Theorem).
+There is an algorithm that, given a multigraph $G$, a partial $\left\lfloor 3\Delta/2\right\rfloor$-edge coloring $\chi$ with $\lambda=m/\Delta$ uncolored edges, extends the coloring to all edges in $O(m\log{n})$ time with high probability.
+This concludes the proof of Theorem B.1.
+B.3 Proof of Lemma B.7
+Let $U$ be the set of $\lambda=O(m/\Delta)$ uncolored edges forming a matching.
+We first scan through the edges $(u,v)\in U$ and check if $\mathsf{miss}_{\chi}(u)\cap\mathsf{miss}_{\chi}(v)\neq\varnothing$ in $O(\Delta)\cdot\lambda=O(m)$ time.
+If this holds for at least $\lambda/2$ of these edges, we can extend the coloring to these edges in $O(m)$ time and we are done.
+Otherwise, let $U^{\prime}\subseteq U$ be the subset of uncolored edges where this does not hold.
+Now, for each color $c$, we compute
+$\texttt{freq}(c):=\sum_{(u,v)\in U^{\prime}}\mathmybb{1}[c\in\mathsf{miss}_{\chi}(v)]$
+by scanning over each $(u,v)\in U^{\prime}$, taking $O(\lambda\Delta)$ time in total. We now define an auxiliary color
+$\gamma:=\arg\max_{c}\texttt{freq}(c)$
+and proceed to construct a vertex-disjoint collection of pre-Shannon fans $\operatorname*{\mathcal{U}}^{\prime}$ with auxiliary color $\gamma$ as described in Algorithm 9.
+[…]
+2 if $(u,v)\in U^{\prime}$ then
+3 if $\gamma\in\mathsf{miss}_{\chi}(v)$ then
+4      Let $(u,w)$ be the edge with $\chi(u,w)=\gamma$
+5    if $\{u,v,w\}\cap S=\varnothing$ then
+6       $\operatorname*{\mathcal{U}}^{\prime}\leftarrow\operatorname*{\mathcal{U}}^{\prime}\cup\{(u,v,w,\gamma)\}$
+7       $S\leftarrow S\cup\{u,v,w\}$
+8
+return $\operatorname*{\mathcal{U}}^{\prime}$
+Algorithm 9 $\textsf{Construct-Pre-Shannon-Fans}(U^{\prime})$
+We can observe that the set $\operatorname*{\mathcal{U}}^{\prime}$ produced by Algorithm 9 consists of vertex-disjoint pre-Shannon fans with auxiliary color $\gamma$, and that this can be computed in $O(\lambda)$ time.
+The following lemma shows that $\operatorname*{\mathcal{U}}^{\prime}$ contains $\Omega(\lambda)$ Shannon fans.
+Lemma B.12.
+We have that $|{\operatorname*{\mathcal{U}}^{\prime}}|\geq\lambda/24$.
+Proof.
+We begin by proving the following claim, which shows that $\gamma$ is available at the endpoints of $\Omega(\lambda)$ many edges in $U^{\prime}$.
+Claim B.13.
+We have that $\textnormal{{freq}}(\gamma)\geq\lambda/12$.
+Proof.
+We can first observe that
+$\sum_{c}\texttt{freq}(c)=\sum_{c}\sum_{(u,v)\in U^{\prime}}\mathmybb{1}[c\in\mathsf{miss}_{\chi}(v)]=\sum_{(u,v)\in U^{\prime}}\sum_{c}\mathmybb{1}[c\in\mathsf{miss}_{\chi}(v)]$
+$\geq\sum_{(u,v)\in U^{\prime}}\left(\left\lfloor\frac{3\Delta}{2}\right\rfloor-\Delta\right)\geq\left(\frac{\Delta-1}{2}\right)\cdot|U^{\prime}|\geq\frac{\Delta}{4}\cdot|U^{\prime}|\geq\frac{\lambda\Delta}{8},$
+where we are using the facts that each vertex has at least $\left\lfloor 3\Delta/2\right\rfloor-\Delta$ missing colors, that $\Delta>1$ (otherwise, the problem is trivial), and that $|U^{\prime}|\geq\lambda/2$.
+Thus, we have that
+$\texttt{freq}(\gamma)\geq\frac{1}{\left\lfloor 3\Delta/2\right\rfloor}\cdot\sum_{c}\texttt{freq}(c)\geq\frac{2}{3\Delta}\cdot\frac{\lambda\Delta}{8}=\frac{\lambda}{12}.\qed$
+It follows from this claim that we can create pre-Shannon fans with $\gamma$ as an auxiliary color with at least $\lambda/12$ of the uncolored edges in $U^{\prime}$.
+Since the uncolored edges in $U^{\prime}$ are vertex-disjoint, two pre-Shannon fans with auxiliary color $\gamma$ on these edges, $(u_{i},v_{i},w_{i},\gamma)$ and $(u_{j},v_{j},w_{j},\gamma)$, intersect if and only if $u_{i}=w_{j}$ and $w_{i}=u_{j}$.
+Thus, any such fan can intersect at most one other such fan.
+It follows that the algorithm produces a collection $\operatorname*{\mathcal{U}}^{\prime}$ of size at least $\lambda/24$.
+∎
+Finally, we can take the set $\operatorname*{\mathcal{U}}^{\prime}$ of vertex-disjoint pre-Shannon fans and apply Lemma B.6 to each $\operatorname*{\boldsymbol{S}}^{\star}\in\operatorname*{\mathcal{U}}^{\prime}$ to compute a set $\operatorname*{\mathcal{U}}$ of at least $|\mathcal{U}^{\prime}|/2=\Omega(\lambda)$ vertex-disjoint Shannon fans or extend $\chi$ to at least $|\mathcal{U}^{\prime}|/2=\Omega(\lambda)$ edges in $|\mathcal{U}^{\prime}|\cdot O(\Delta)\leq\lambda\cdot O(\Delta)\leq O(m)$ time.
+Appendix C Edge Coloring Algorithms for Small Values of $\Delta$
+The main contribution and the primary message of our work is presenting the first near-linear time algorithm for $(\Delta+1)$-edge coloring and polynomial time factor improvements over prior work.
+However, it turns out that a simple corollary of our main results, plus the prior work of [BD23], is also enough to obtain quite efficient algorithms in the regime when $\Delta=n^{o(1)}$, namely, an algorithm with only $O(m\log{\Delta})$ runtime with high probability (basically, replacing $\log{n}$-term in our main results with a $\log{\Delta}$-term instead).
+In this regime, previously, the best known bounds where $O(m\Delta^{4}\log\Delta)$ time randomized algorithm of [BD24] for $(\Delta+1)$-edge coloring and $O(m\log{\Delta})$ time randomized algorithm of [Ass25] for $(\Delta+O(\log{n}))$-edge coloring.
+Corollary C.1.
+There are randomized algorithms that, given any undirected multigraph $G$ on $n$ vertices and $m$ edges with maximum degree $\Delta$ and maximum edge multiplicity $\mu$, in $O(m\log{\Delta})$ time, with high probability, output:
+  1. 1.
+    a $(\Delta+1)$-edge coloring of any simple graph $G$ (Vizing’s theorem for simple graphs);
+  2. 2.
+    a $(\Delta+\mu)$-edge coloring of any multigraph $G$ (Vizing’s theorem for multigraphs);
+  3. 3.
+    a $\left\lfloor 3\Delta/2\right\rfloor$-edge coloring of any multigraph $G$ (Shannon’s theorem for multigraphs).
+Corollary C.1 improves the runtime dependence of Theorem 1.1Theorem A.1, and Theorem B.1 from $O(m\log{n})$ time to $O(m\log{\Delta})$ time instead.
+In the following, we prove the first part of this result, namely, for $(\Delta+1)$-edge coloring of simple graphs, and then show how this argument extends to the other two parts as well.
+C.1 Proof of Corollary C.1 for Simple Graphs
+Recall that by Lemma 10.4, we already have an algorithm that, with high probability, in $O(m\log{\Delta})$ time, $(\Delta+1)$ colors all but $m/\Delta$ edges of the graph.
+Thus, our task is now to extend the coloring to these edges as well.
+Previously, in Lemma 10.5, we did this by coloring a constant fraction of the edges repeatedly (using Lemma 10.2); since, we need $\Theta(\log{(m/\Delta)})$ attempts to color all edges and each one requires $O(m)$ time, this lead to an $O(m\log{n})$ time algorithm.
+To prove Corollary C.1, we split this final task: we will again run the algorithm of Lemma 10.2, but this time to reduce the number of uncolored edges from $m/\Delta$ to $m/\operatorname*{poly}(\Delta)$ for some arbitrarily large polynomial in $\Delta$; thus, we need $\Theta(\log{((m/\Delta)/(m/\operatorname*{poly}(\Delta)))})=\Theta(\log{\Delta})$ steps of coloring, and hence $O(m\log{\Delta})$ time in total.
+Then, to handle the very last $m/\operatorname*{poly}(\Delta)$ remaining edges, we will run the algorithm of [BD24], which extends the coloring to each uncolored edge in essentially (but not formally) $\operatorname*{poly}(\Delta)$ time per uncolored edge for some polynomial in $\Delta$.
+Thus, the second part of the algorithm only takes $O(m)$ time now, leaving us with an $O(m\log{\Delta})$ time algorithm all in all.
+We formalize this approach in the following (which specifically requires extra care for obtaining a high probability guarantee).
+Lemma C.2.
+Let $c\geq 1$ be any arbitrary fixed constant and $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$.
+There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ with $\lambda=m/\Delta$ uncolored edges, extends the coloring to all but $\Theta(\lambda_{0})$ uncolored edges in $O(m\log{\Delta})$ time with high probability.
+Proof.
+We simply run our Lemma 10.2 with the given parameters $\lambda$ and $\lambda_{0}$.
+By Lemma 10.2, the expected runtime will be some $T_{0}=O(m\log{(\lambda/\lambda_{0})}+\Delta\lambda)=O(m\log{\Delta})$ since $\lambda_{0}\geq m/\Delta^{2c}$ for some absolute constant $c\geq 1$.
+Moreover, by the same lemma,
+$\Pr\left[\text{runtime of the algorithm}\geq 2000\cdot T_{0}\right]\leq\exp\left(-10\lambda_{0}\right)\leq\exp\left(-\Theta(m\log{n})/n\right)=1/\operatorname*{poly}(n),$
+since $\lambda_{0}\geq m\log{n}/n$. Thus, the algorithm finishes in $O(m\log{\Delta})$ time with high probability. ∎
+We now present the final part for coloring the last remaining edges.
 For this step, we use the following result from [BD23] (itself based on [Ber22, Chr23]) that allows for coloring a random uncolored edge in essentially (but not exactly) $\operatorname*{poly}(\Delta)$ time.
 We note that a recent work of [BD24] presents a further improvement to these bounds (in terms of the exponent of $\Delta$) but as our dependence on $\Delta$ is only logarithmic, this improvement is inconsequential and thus we stick with the simpler work of [BD23].
-Proposition A.7 ([BD23]).
+Proposition C.3 ([BD23]).
 There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$, and $\lambda$ uncolored edges, extends the coloring to a randomly chosen uncolored edge in expected $\operatorname*{poly}(\Delta)\cdot\log{(n/\lambda)}$ time.
 We use this result to obtain the following lemma for our purpose.
-Lemma A.8.
-Let $c$ be the $\operatorname*{poly}(\Delta)$-exponent in Proposition A.7, and $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$.
+Lemma C.4.
+Let $c$ be the $\operatorname*{poly}(\Delta)$-exponent in Proposition C.3, and $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$.
 There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ and $\lambda=\Theta(\lambda_{0})$ uncolored edges, extends the coloring to all uncolored edges in $O(m\log\Delta)$ time with high probability.
@@
 Now consider the case when $\lambda_{0}=m/\Delta^{2c}$.
-We run the algorithm of Proposition A.7 with the following modification: when coloring a random edge, if the runtime becomes two times more than the expected time guarantee of Proposition A.7, we terminate the coloring and repeat from the beginning.
+We run the algorithm of Proposition C.3 with the following modification: when coloring a random edge, if the runtime becomes two times more than the expected time guarantee of Proposition C.3, we terminate the coloring and repeat from the beginning.
 Thus, each step takes $O(\Delta^{c}\cdot\log{(n/i)})$ time deterministically and succeeds in extending the coloring by one edge with probability at least half.
@@
 where the first inequality is because the probability that $\ell_{i}$ increases by one each time is at most $1/2$ by Markov’s inequality.
-While $X_{i}$’s are technically not independent (the choice of some edge being colored may change the structure of the graph for coloring next uncolored edge), they are stochastically dominated by independent random variables with above tail bounds (given the guarantee of Proposition A.7 for each uncolored edge).
-Thus, we can apply Proposition A.5 with parameters $\{\alpha_{i}\}_{i=1}^{\lambda}$ as above and $\beta_{i}=1$ for all $i\in[\lambda]$ to obtain that for $t:=m/\Delta^{c}$,
+While $X_{i}$’s are technically not independent (the choice of some edge being colored may change the structure of the graph for coloring next uncolored edge), they are stochastically dominated by independent random variables with above tail bounds (given the guarantee of Proposition C.3 for each uncolored edge).
+Thus, we can apply Proposition 10.3 with parameters $\{\alpha_{i}\}_{i=1}^{\lambda}$ as above and $\beta_{i}=1$ for all $i\in[\lambda]$ to obtain that for $t:=m/\Delta^{c}$,
 $\Pr\!\left[\sum_{i=1}^{\lambda}X_{i}\geq\frac{2m}{\Delta^{c}}\right]\leq\Pr\left[\sum_{i=1}^{\lambda}X_{i}\geq\sum_{i=1}^{\lambda}\alpha_{i}+t\right]$ $\leq\exp\left(-\frac{1}{4\log{n}}\cdot\frac{m}{2\Delta^{c}}\right)\ll 1/\operatorname*{poly}(m),$
@@
 ∎
-Finally, we present an algorithm for taking care of the gap between the $m/\Delta$ uncolored edges left by Lemma A.6 and the $\Theta(1)\cdot\max\{m/\Delta^{2c},m\log{n}/n\}$ edges that Lemma A.8 can handle.
-Lemma A.9.
-Let $c=\Theta(1)$ and $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$ be as in Lemma A.8.
-There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ with $\lambda=m/\Delta$ uncolored edges, extends the coloring to all but $\Theta(\lambda_{0})$ uncolored edges in $O(m\log{\Delta})$ time with high probability.
+The proof of Corollary C.1 for $(\Delta+1)$ edge coloring is now as follows: run Lemma 10.4 to color all but $m/\Delta$ edges.
+Then, let $c$ be the constant in the exponent of $\operatorname*{poly}(\Delta)$ in Proposition C.3 and run Lemma C.2 to extend the coloring to all but $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$ edges.
+Finally, run Lemma C.4 to extend the coloring to all these remaining $\lambda_{0}$ edges.
+Since each of these algorithms take $O(m\mathrm{lo}{\Delta})$ time and succeed with high probability, we obtain the final algorithm.
+C.2 Proof of Corollary C.1 for Multigraphs
+To extend the proof to multigraphs, recall that by Lemma A.10 and Lemma B.10, we can, respectively, $(\Delta+\mu)$ color or $\left\lfloor 3\Delta/2\right\rfloor$ color all but $m/\Delta$ edges of a multigraph in $O(m\log{\Delta})$ time.
+Thus, we can follow the same exact strategy as in $(\Delta+1)$ coloring approach outlined in the previous subsection, and we only need an analogue of the result of [BD23] to conclude the proof.
+Such results are also already established by [Dha24a] and we do not need to reinvent the wheel here.
+Proposition C.5 (A subroutine in [Dha24a, Theorem 1.6; part 1]).
+There is an algorithm that, given a multigraph $G$, a partial $(\Delta+\mu)$-edge coloring $\chi$, and $\lambda$ uncolored edges, extends the coloring to a randomly chosen uncolored edge in expected $\operatorname*{poly}(\Delta)\cdot\log{(n/\lambda)}$ time.
+Proposition C.6 (A subroutine in [Dha24a, Theorem 1.5; part 1]).
+There is an algorithm that, given a multigraph $G$, a partial $\left\lfloor 3\Delta/2\right\rfloor$-edge coloring $\chi$, and $\lambda$ uncolored edges, extends the coloring to a randomly chosen uncolored edge in expected $\operatorname*{poly}(\Delta)\cdot\log{(n/\lambda)}$ time.
+The rest of the proof is verbatim as in simple graphs in Corollary C.1 and we omit it here.
+Appendix D A Specialized Concentration Inequality
+Throughout the paper, for optimizing log factors in the runtime of our algorithms, we needed the following specialized concentration inequality.
+For completeness, we now prove this inequality using a standard argument based on moment generating functions.
+Proposition (Restatement of Proposition 10.3).
+Let $\{X_{i}\}_{i=1}^{n}$ be $n$ independent non-negative random variables associated with parameters $\{\alpha_{i}\}_{i=1}^{n}$ and $\{\beta_{i}\}_{i=1}^{n}$ such that for each $i\in[n]$, $\alpha_{i},\beta_{i}\geq 1$, and for every $\delta>0$,
+$\Pr\left[X_{i}\geq(1+\delta)\cdot\alpha_{i}\right]\leq\exp\left(-\delta\cdot\beta_{i}\right);$
+then, for every $t\geq 0$,
+$\Pr\left[\sum_{i=1}^{n}X_{i}\geq\sum_{i=1}^{n}\alpha_{i}+t\right]\leq\exp\left(-\left(\min_{i=1}^{n}\frac{\beta_{i}}{2\alpha_{i}}\right)\cdot\left(t-2\sum_{i=1}^{n}\frac{\alpha_{i}}{\beta_{i}}\right)\right).$
 Proof.
-We simply run our Lemma A.4 with the given parameters $\lambda$ and $\lambda_{0}$.
-By Lemma A.4, the expected runtime will be some $T_{0}=O(m\log{(\lambda/\lambda_{0})}+\Delta\lambda)=O(m\log{\Delta})$ since $\lambda_{0}\geq m/\Delta^{2c}$ for some absolute constant $c\geq 1$.
-Moreover, by the same lemma,
-Pr[runtime of the $\Pr\left[\text{runtime of the algorithm}\geq 2000\cdot T_{0}\right]\leq\exp\left(-10\lambda_{0}\right)\leq\exp\left(-\Theta(m\log{n})/n\right)=1/\operatorname*{poly}(n),$
-since $\lambda_{0}\geq m\log{n}/n$. Thus, the algorithm finishes in $O(m\log{\Delta})$ time with high probability. ∎
-Putting our above lemmas together concludes the proof of Theorem A.1 as follows:
-Proof of Theorem A.1.
-The proof follows immediately from the lemmas in Step 4:
-  • By Lemma A.6, we can color all but $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability;
-  • Then, by Lemma A.9, we can extend the coloring to all but $\Theta(\lambda_{0})$ edges of the graph in $O(m\log{\Delta})$ time with high probability, where $\lambda_{0}$ is as in Lemma A.8;
-  • Finally, by Lemma A.8, we can color all the remaining $\Theta(\lambda_{0})$ uncolored edges in $O(m\log{\Delta})$ time with high probability.
-This concludes the proof. ∎
+Let $s:=1/2\cdot\min_{i}\beta_{i}/\alpha_{i}$ and for $i\in[n]$, $Y_{i}:=\max(X_{i}-\alpha_{i},0)$. Letting $Y:=\sum_{i=1}^{n}Y_{i}$, leads
+$\Pr\left[\sum_{i=1}^{n}X_{i}\geq\sum_{i=1}^{n}\alpha_{i}+t\right]\leq\Pr\left[Y\geq t\right],$
+for all $t\geq 0$. We further have,
+$\Pr\left[Y\geq t\right]=\Pr\left[e^{s\cdot Y}\geq e^{s\cdot t}\right]\leq e^{-s\cdot t}\cdot\mathbb{E}\left[e^{s\cdot Y}\right]=e^{-s\cdot t}\cdot\mathbb{E}\left[\prod_{i=1}^{n}e^{s\cdot Y_{i}}\right]=e^{-s\cdot t}\cdot\prod_{i=1}^{n}\mathbb{E}\left[e^{s\cdot Y_{i}}\right]$ (4)
+where the inequality is by Markov inequality and the final equality is by the independence of the variables. Now, for every $i\in[n]$,
+$\mathbb{E}\left[e^{s\cdot Y_{i}}\right]$ $=\int_{y=0}^{\infty}\Pr\left[Y_{i}=y\cdot\alpha_{i}\right]\cdot\exp\left(s\cdot y\cdot\alpha_{i}\right)~dy$
+$=\int_{y=0}^{\infty}\Pr\left[X_{i}=(1+y)\cdot\alpha_{i}\right]\cdot\exp\left(s\cdot y\cdot\alpha_{i}\right)~dy$ (by the connection between $X_{i}$ and $Y_{i}$ when $y\geq 0$)
+$\leq\int_{y=0}^{\infty}\exp\left(-y\cdot\beta_{i}\right)\cdot\exp\left(s\cdot y\cdot\alpha_{i}\right)~dy$ (by the tail inequality of $X_{i}$ in the proposition statement)
+$=\frac{1}{\beta_{i}\cdot(1-s\cdot({\alpha_{i}}/{\beta_{i}}))}$ (since the integral of $e^{-Kx}=-e^{-Kx}/K$ and $s<\beta_{i}/\alpha_{i}$ thus the value of this function in $\infty$ is zero)
+$\leq\frac{1}{1-s\cdot(\alpha_{i}/\beta_{i})}$ (as $\beta_{i}\geq 1$)
+$\leq\exp\left(2s\cdot\alpha_{i}/\beta_{i}\right).$ (as $s\cdot\alpha_{i}/\beta_{i}<1/2$ and $1-x\geq e^{-2x}$ for $x<1/2$)
+Plugging this bounds in Equation 4, implies that
+$\Pr\left[Y\geq t\right]\leq\exp\left(-s\cdot t+\sum_{i=1}^{n}2s\cdot\frac{\alpha_{i}}{\beta_{i}}\right)=\exp\left(-\left(\min_{i=1}^{n}\frac{\beta_{i}}{2\alpha_{i}}\right)\cdot\left(t-2\sum_{i=1}^{n}\frac{\alpha_{i}}{\beta_{i}}\right)\right),$
+by the choice of $s$, concluding the proof. ∎